### v1

# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

Abstract
Migration has been a core topic in German political debate, from millions of expellees post World War II over labor migration to refugee movements in the recent past. Studying political speech regarding such wide-ranging phenomena in depth traditionally required extensive manual annotations, limiting the scope of analysis to small subsets of the data. Large language models (LLMs) have the potential to partially automate even complex annotation tasks. We provide an extensive evaluation of a multiple LLMs in annotating (anti-)solidarity subtypes in German parliamentary debates compared to a large set of thousands of human reference annotations (gathered over a year). We evaluate the influence of model size, prompting differences, fine-tuning, historical versus contemporary data; and we investigate systematic errors. Beyond methodological evaluation, we also interpret the resulting annotations from a social science lense, gaining deeper insight into (anti-)solidarity trends towards migrants in the German post-World War II period and recent past. Our data reveals a high degree of migrant-directed solidarity in the postwar period, as well as a strong trend towards anti-solidarity in the German parliament since 2015, motivating further research. These findings highlight the promise of LLMs for political text analysis and the importance of migration debates in Germany, where demographic decline and labor shortages coexist with rising polarization.
keywords:
solidarity, large language models, annotation, political discourse
$\cup@datesfalse$
Aida Kostikova]aida.kostikova@uni-bielefeld.de
1 Introduction
Migration is a defining feature of the 21st century, explicitly recognized in the 2030 Agenda for Sustainable Development wilkinson2022climate (105, 1). As the worldbank2023 (106) stresses, global economic imbalances, demographic divergence, and climate change make migration increasingly unavoidable, especially from the Global South to the Global North almulhim2024climate (3, 55, 69). However, if governed responsibly, migration can also drive prosperity and contribute to achieving the Sustainable Development Goals worldbank2023 (106, 38, 1).
Within Europe, Germany occupies a particularly important position as a major migration destination eurostat2024migration (32). As the EU’s largest economy, it offers labor market opportunities, institutional support, and stability bertoli2016crisis (10). At the same time, demographic decline and skilled-labor shortages make sustained immigration economically beneficial smith2025demographic (27, 29, 4). Migration has also profoundly shaped Germany’s political development, from the resettlement of millions of expellees after WWII, to the labor migration of the 1950s–1970s, to refugee movements in the past decade kaya2017inclusion (52, 36). This raises important questions: how does German politics approach migration, and to what extent does it express solidarity – or anti-solidarity – toward migrants?
Answering such questions requires systematic analysis of political discourse across long time spans, traditionally relying on manual textual analysis and qualitative interpretation, which remain powerful tools randour2020twenty (78). However, such methods are labor-intensive bhattacherjee2012social (11) and thus often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized babalola2021current (7). Computational social science (CSS) emerged to address the scale and complexity of modern data, yet it often struggles with theoretical alignment and interpretive depth lazer2024future (62). Large language models (LLMs) promise to bridge this gap by enabling automated, in-depth analysis of large-scale text. While LLMs have shown strong performance on tasks like ideological stance detection ziems2024can (109), it remains an open question whether they can handle more complex, theory-driven sociopolitical annotation tasks. We provide a partial answer: If accompanied by substantial validation and careful interpretation, LLMs can approximate the performance of trained human annotators and generate annotations that help quantify large-scale trends of complex constructs. In particular, we apply LLMs to annotate solidarity and anti-solidarity towards women and migrants in German parliamentary speeches, compare LLM annotations to thousands of human reference annotations, and interpret the (anti-)solidarity trends between 1949-1969 and 2009-2025 in light of German political history.
Figure 1: Overview of the annotation scheme adapted from thijssen2012mechanical (97). Statements are first classified at a high level as expressing solidarity, anti-solidarity, mixed, and none. Solidarity and anti-solidarity instances are then further categorized into subtypes based on the underlying rationale: group-based, exchange-based, compassionate, and empathic. The example sentence in the figure illustrates exchange-based anti-solidarity.
We focus on solidarity as a particularly demanding test case for such an approach. While frequently invoked in political discourse, solidarity has often been treated as a symbolic notion rather than an operational concept kozlowska2024solidarity (58), which has limited large-scale empirical research. To address this gap, we approach solidarity as a socially meaningful category that can be systematically studied in political text. For this purpose, we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97), which provides a structured and annotatable framework. It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure 1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification). Germany offers a particularly suitable context for such a study: solidarity is deeply rooted in its political vocabulary fiegle2003solidarite (34, 94, 84), and its multi-party system enables us to trace how meanings of (anti-)solidarity vary across parties and over time.
In a prior conference paper, we already performed a first analysis of LLM annotations of (anti-)solidarity towards women and migrants on this data set kostikova-fine (57). In this paper, we 1) evaluate a much wider range of LLMs, 2) systematically evaluate the effect of prompts, model size, and historic (versus contemporary) data, 3) include a much larger data set of human annotations, 4) validate robustness by comparing the time series of model-predicted (anti-)solidarity shares from the selected model (Llama-3.3-70B) and an independent model (GPT-4), and 5) most importantly, we interpret the (anti-)solidarity trends towards migrants over time to derive novel insights for political science. In particular, we find a) that different groups of migrants are associated with different types of (anti-)solidarity in German parliamentary debates during the post-World War II period; and b) that anti-solidarity towards migrants has dramatically increased in the German parliament over the past 10 years, indicating a shift to the right (especially in the liberal and conservative parties) on this issue that motivates further research.
1.1 Related Work
Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; and (ii) the use of LLMs as annotation tools in political and social science contexts. While prior studies have engaged with each of these dimensions individually, few have combined them to enable large-scale, longitudinal analysis of solidarity in political speech. We briefly review these literatures below.
Solidarity in Political Discourse and NLP
Sociological theories of solidarity often begin with Durkheim’s classical distinction between mechanical and organic forms, which describe different bases for social cohesion. While influential, this binary model has been criticized for its structuralism and linear view of social change. thijssen2012mechanical (97) expands on Durkheim by integrating recognition-based theories such as those developed by honneth1996struggle (45, 46) and proposing a fourfold typology with group-based, compassionate, exchange-based, and empathic subtypes, which can be applied to both historical and contemporary political discourse. thijssen2022s (98) apply this typology to Belgian party manifestos, showing how different parties emphasize different solidarity frames in line with broader ideological divides. We build on this by applying the typology to parliamentary debates, extending the temporal scope and scaling the analysis with LLMs. While some studies have examined solidarity longitudinally using interviews or survey data schonweitz2024solidarity (89, 92), these approaches are not designed for large-scale text analysis or fine-grained typological classification.
At the same time, solidarity has received relatively little attention in computational text analysis, where more application-driven concerns such as hate speech and misinformation detection have dominated the field. Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements weber2024social (103, 28), the COVID-19 pandemic kalaloi2021mediated (51), and natural disasters santhanam2019stand (88). These works typically use supervised models to detect solidarity or anti-solidarity in large-scale Twitter data and track how such expressions evolve over short timeframes. While they demonstrate the feasibility of automated solidarity detection, they focus on specific, time-bound crises and treat solidarity as a binary sentiment-like category rather than a differentiated moral frame. In contrast, our work applies a theory-driven typology to political speech, enabling fine-grained analysis of solidarity in parliamentary discourse.
LLMs as Annotation Tools in Political and Social Science.
LLMs are increasingly used in political and social science research. Most applications focus on standard classification tasks such as political ideology detection (heseltine2024large, 43, e.g.); sentiment analysis lashitew2025corporate (61, 37); stance detection (cao2024can, 19, e.g.); and misinformation detection (huang2025unmasking, 47, e.g.). For broader surveys of LLM applications, benefits, and challenges in social and political science, see thapa2025large (95, 66, 109). However, conclusions about LLM performance in these fields vary. Some studies report that LLMs show near-human performance rytting2023towards (87) or even outperform humans, as well as traditional models, on short texts and tasks like ideology classification le2025positioning (63, 43). In contrast, several studies show that smaller fine-tuned models like BERT still outperform LLMs, especially when labeled data is available wang2024selecting (102, 18) or when tasks require expert-defined taxonomies and contextual reasoning ziems2024can (109). Our work provides a systemic evaluation of LLMs for such a complex, theory-informed task, namely (anti-)solidarity detection.
Beyond standard classification tasks, a subset of research has applied language models to more complex, theory-driven categories. For example, erhard2025popbert (30) develop PopBERT, a fine-tuned transformer trained to detect populist rhetoric and ideological dimensions in German Bundestag speeches using a custom annotation scheme. card2022computational (20) analyze 140 years of U.S. political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, identify rhetorical frames, and detect dehumanizing metaphors, demonstrating the utility of fine-tuned LMs for large-scale, theory-aligned analysis. Related work has applied fine-tuned models to pragmatic and rhetorical categories in German parliamentary debates, such as speaker attribution, speech acts, and opinion role labelling reinig2024politics (83, 81, 80). In contrast to fine-tuned approaches, ranjit2024oath (79) use general-purpose LLMs (e.g., GPT-4, Flan-T5) to apply a custom typology of attitudes toward homelessness, developed from framing theory and grounded theory. Their evaluation focuses on aggregate trends and correlations with external structural indicators, using LLMs to annotate a large corpus of tweets at scale and over time.
Our work differs in several important ways. Unlike erhard2025popbert (30) and card2022computational (20), we use state-of-the-art LLMs and avoid the need for supervised data sets and fine-tuning. Unlike ranjit2024oath (79), who also use prompted LLMs, we adopt a typology grounded in established sociological theory rather than one developed inductively for a specific domain. We apply this typology to more formal political speech rather than short, informal social media texts, and evaluate model performance through systematic comparison with human annotations. Finally, our analysis spans over 150 years of parliamentary debate, encompassing two case studies, historical and contemporary, and enabling fine-grained, longitudinal insight into political discourse.
2 Methodology
2.1 Data
We build on the DeuParl dataset introduced in DeuParl (101), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022. We extend this corpus to include protocols up to June 2025. We follow the preprocessing steps from DeuParl (101) and include metadata such as date, session number, and legislative period.
We reuse the keyword sets (see a full list of keywords in A) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in kostikova-fine (57), consisting of 32 and 18 terms respectively. Based on these keywords, we extract 63k migrant-related and 138k women-related sentences (instances) from DeuParl, additionally including three sentences before and after each instance for context. We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) provided by poradaautomatische (77). See B for details. Fig. 2 shows the number of instances over time, as well as their share relative to all sentences in the DeuParl dataset. Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015, coinciding with refugee crisis debates. Notably, migrant mentions also surpass women-related discourse in the immediate post-war years.
Figure 2: Absolute number of instances (left) and their share relative to all DeuParl sentences (right) for the Woman and Migrant categories.
2.2 Data Annotation
To validate our models and enable further qualitative analysis, we build on the human-annotated dataset introduced in kostikova-fine (57), which comprises 2,864 instances (1,437 for migrants, 1,427 for women) from the DeuParl corpus. [^1] Annotation was carried out by five trained student assistants through an iterative process, achieving Cohen’s $\kappa$ of 0.62 (high level) and 0.42 (fine-grained). Final labels in the dataset are assigned by majority vote. Each instance is labeled according to a fine-grained solidarity framing scheme that distinguishes four high-level stance categories – solidarity (willingness to share resources, following Lahusen.2018 (60) and TwitterDataset (50)), anti-solidarity (restriction or exclusion of groups), mixed (both supportive and opposing elements), and none (neutral or unrelated content). Instances labeled as solidarity or anti-solidarity are further assigned one of four frame types adapted from thijssen2022s (98): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance in contributions), and empathic (respect for or rejection of social diversity).
We extend the evaluation to a newly annotated set of 756 instances for Migrant with more historical coverage (before 1950). For this set, annotations were conducted by 4 annotators following the same procedure as in kostikova-fine (57) Statistics on the human-annotated dataset are presented in Fig. 3. Overall, the label distribution differs notably by target group: woman-related instances are predominantly labeled as solidarity (74.1%), while migrant-related instances show a more diverse pattern, with lower solidarity (51.6%) and a higher share of anti-solidarity (21.3%).
Figure 3: Distribution of instances in the human annotated dataset across time and target groups. Total indicates number of all annotated samples for the given target group, including the None label. See Table 2 for the actual numbers of instances in the human annotated dataset.
2.3 Models and Experiments
We evaluate a diverse suite of models, ranging from smaller supervised models (BERT, fine-tuned SBERT) to proprietary LLMs (GPT-4) and recent open-weight instruction-tuned models. The final list includes:
  1. (i)
    smaller pre-trained LMs like BERT and SBERT;
  2. (ii)
    closed-source LLMs like gpt-4-1106-preview;
  3. (iii)
    open-weight LLMs such as Llama-3-8B-It, Llama-3.3-70B-It (and fine-tuned Llama-3-70B-It), Gemma-2-9b-It, Gemma-2-27b-It, and Mistral-2-Large.
Data and Metrics
We adopt the 70/15/15 train/dev/test split strategy as in kostikova-fine (57), generating three random splits of approximately 2,000 training, 400 development, and 430 test instances for each target group. These are used to train and evaluate supervised models. For inference-only experiments, we use only the test sets: Test 1 (original) and Test 2 (newly added, 756 instances), the latter offering broader historical coverage (1860s–1950s). We report Macro F1 to account for class imbalance, along with per-class F1 scores for both high-level and fine-grained classification tasks. [^2] High-level classification refers to the four stance categories: solidarity, anti-solidarity, mixed, and none. Fine-grained classification includes ten labels: the four subtypes of solidarity, the four subtypes of anti-solidarity, plus mixed and none.
Models
We use manually designed prompts for each target group. Each prompt consists of a classification instruction concatenated with the parliamentary instance, i.e., the keyword sentence plus its three-sentence left/right context as defined in 2.1. Note that, although speaker and party metadata are available in the corpus, this information is removed from the model input. The classification instruction includes (i) chain-of-thought reasoning in zero- and few-shot settings (via think step by step); (ii) label definitions; and (iii) a two-step classification format: high-level first, followed by subtype for solidarity or anti-solidarity prediction. [^3] For GPT-4, both steps are combined in a single prompt (see Figure 21 for Migrant and Figure 22 in the Appendix for Woman). For open-weight models, we instead use two prompts: one for high-level classification and, conditional on that output, a second for subtype selection (see Figure 24 for Migrant and Figure 27 in the Appendix for Woman). In few-shot mode, prompts include one representative example per label.
We apply these prompts to closed-source models like gpt-4-1106-preview, and to open-weight LLMs including Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it (full precision) in inference-only setting. Larger models (Llama-3.3-70B-Instruct and Mistral-Large-Instruct-2407) are run with $Q4_{0}$ quantization. Additionally, we fine-tune Llama-3-8B-It and Llama-3-70B-It using a three-model pipeline (high-level and two subtype classifiers), applying instruction-style fine-tuning with true labels paired with GPT-4-generated explanations as training targets.
We use BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details for all models used in our experiments are provided in C.
3 Results
In this section, we evaluate models against a human-annotated dataset to determine reliability for social scientific analysis of solidarity frames in parliamentary texts. Our aim is twofold: (i) to identify a model whose predictions approximate human-level annotation quality across both high-level and fine-grained stance categories, and (ii) to ensure the selected model is practically feasible to deploy at scale. Specifically, we ask whether recent open-weight models can match the performance of GPT-4, which was previously found to be highly reliable (though costly) kostikova-fine (57), through techniques such as few-shot prompting or lightweight fine-tuning.
3.1 Model Performance Overview
Table 1 reports macro F1 scores for each model across both inference-based and fine-tuned approaches. We benchmark these results against a human upper bound, estimated using a leave-one-out (LOO) procedure, [^4] For each annotator, we exclude their labels, compute a consensus label from the remaining annotators, and calculate the macro F1 score between the excluded annotator and the consensus. We then average these scores across all annotators. and report performance for both high-level and fine-grained classification tasks, with the latter shown in parentheses. Our observations are structured around two dimensions: (i) overall model performance, (ii) task and label-related challenges.
Model Method Woman Migrant (T1) Migrant (T2)
GPT-4 Zero-shot 0.37 (0.60) 0.42 (0.73) 0.29 (0.57)
Few-shot 0.37 (0.54) 0.43 (0.63) 0.29 (0.51)
Mistral-2-Large Zero-shot 0.20 (0.50) 0.27 (0.56) 0.24 (0.45)
Few-shot 0.26 (0.51) 0.38 (0.57) 0.26 (0.44)
Llama-3.3-70B Zero-shot 0.23 (0.50) 0.32 (0.54) 0.25 (0.47)
Few-shot 0.30 (0.50) 0.37 (0.65) 0.28 (0.53)
Llama-3-70B Fine-tuned 0.26 (0.49) 0.27 (0.65) 0.21 (0.47)
Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.17 (0.40)
Few-shot 0.14 (0.35) 0.19 (0.34) 0.15 (0.40)
Fine-tuned 0.16 (0.38) 0.20 (0.40) 0.13 (0.27)
Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.25 (0.45)
Few-shot 0.25 (0.55) 0.33 (0.48) 0.26 (0.46)
Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.22 (0.46)
Few-shot 0.23 (0.44) 0.31 (0.54) 0.24 (0.46)
BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.23 (0.45)
SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.18 (0.36)
Human upper bound 0.34 (0.62) 0.48 (0.75) 0.35 (0.57)
Table 1: Comparative performance (macro F1) of models on high-level (in parentheses) and fine-grained tasks. ?Migrant (T1)? uses the original benchmark; ?Migrant (T2)? is newly introduced in this paper. Green cells indicate the best-performing score for each target group, based on a model’s best setting (e.g., zero-shot or few-shot); yellow cells indicate the best-performing score from a different model (i.e., the second-best model overall).
  1. (i)
    Overall model performance.
    • GPT-4 achieves the highest scores, but Llama-3.3-70B is a strong alternative. Across all target groups (Woman and both Migrant test sets), GPT-4 (zero-shot) consistently shows the top performance, coming closest to the human upper bound in both high-level and fine-grained classification. Llama-3.3-70B is the best-performing open model, showing meaningful gains from few-shot prompting (in particular in the high-level classification).
    • Model size correlates with performance, but not perfectly. Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure 4 for the relationship between model size and performance). Mid-sized models follow, and smaller models perform worst. However, this relationship is not strictly linear: for example, Llama3-8B-it underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B-it.
      Figure 4: Model size vs. performance on high-level and fine-grained classification tasks in Migrant (T1). Each model is plotted with its best-performing configuration (zero- or few-shot). The x-axis is logarithmic, so spatial distances do not represent linear differences in parameter count.
    • Examples help mid-sized models more than small or very large ones. Few-shot prompting improves performance primarily for mid-sized models such as Gemma-9B-it, Gemma-27B-it, Mistral-2-Large, and Llama-3.3-70B-it. In contrast, smaller models like Llama-3-8B-it show no benefit, and GPT-4 sees minimal gains—likely because its zero-shot performance is already near-optimal. Fine-tuned variants (e.g., Llama-3.3-70B-
      it
      , Llama-3-8B-it) often show no improvement over inference-only models.
  2. (ii)
    Task and label-related challenges.
    • All models perform worse on the fine-grained task, underscoring its complexity: unlike the more intuitive high-level categories (e.g., solidarity, anti-solidarity), it requires alignment with annotation guidelines grounded in social theory, where even human annotators show limited agreement (reaching macro F1 score 0.34 for Woman and 0.48 for Migrant).
    • At the high level, solidarity is the easiest to detect (see Figure 5; full scores in Table 3 in the Appendix), especially in the Woman category, where even smaller models like BERT (0.77) and Gemma-9B (0.78) perform strongly. Anti-solidarity and mixed stance are the most challenging categories, with GPT-4 achieving the highest high-level anti-solidarity scores (0.64 for Migrant, 0.65 for Woman), but mid-size models (e.g., Mistral, Llama-3.3-70B, Gemma-27B, Gemma-9B) show moderate performance (with scores ranging from 0.52 to 0.54). Only GPT-4 and Llama-3.3-70B reliably handle mixed stances (e.g., 0.38–0.43 for Migrant, 0.30–0.37 for Woman).
    • On the fine-grained level, GPT-4, Llama-3.3-70B, and Gemma models offer the most balanced coverage across subtypes. GPT-4 stands out on exchange-based solidarity (0.56), group- and exchange-based anti-solidarity for Migrant (0.47 and 0.46), and on empathic and compassionate anti-solidarity for Woman (0.67 and 0.50). Among open models, Llama-3.3-70B performs particularly well on compassionate solidarity (0.51), and shows strong results on group- and exchange-based anti-solidarity (0.42 and 0.35). Interestingly, the Gemma models are able to distinguish the rare compassionate anti-solidarity subtype for Migrant with some success. Most other models struggle with fine-grained anti-solidarity, with scores close to zero.
      Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For the Migrant group, the two test sets (Test 1 and Test 2) are combined.
    • Performance drops on historical data. All models perform worse on the historical Migrant (T2) set, including the strongest ones. This reflects a broader pattern: model performance tends to decline on older data, as shown by macro F1 scores over time (see Figure 16)—so does agreement among human annotators (Cohen’s Kappa). However, while human trends are smooth and monotonic, model trajectories are less stable. Stronger models such as GPT-4, Mistral-2-Large, Llama-3.3-70B, and BERT show a general upward trend but exhibit clear dips in the 1930s and 1960s. In contrast, smaller models (e.g., Llama-3-8B, SBERT, Gemma-9B) show little to no temporal coherence.
In summary, GPT-4 remains the strongest model overall. While performance generally scales with model size, mid-sized models can approximate GPT-4 when given examples. Among open-weight alternatives, Llama-3.3-70B emerges as the most viable option, particularly for the Migrant category, delivering competitive results in both high-level and fine-grained classification, while offering a more cost-effective solution. We therefore adopt it for the subsequent social scientific analyses. However, since it still lags behind GPT-4 in capturing certain fine-grained subtypes, we further validate its reliability by comparing the temporal patterns produced by both models in 4.1.
3.2 Error Analysis
In this subsection, we examine the errors made by the two strongest models in our experiments, GPT-4 and Llama-3.3-70B, as determined by macro F1 scores. To identify where these models struggle, we compare their confusion matrices on the full Migrant test set (Figure 6 for high-level; see Figure 17 in the Appendix for confusion matrices for all models included in the study and Figure 18 in the Appendix for fine-grained label confusion matrices) with a human reference, constructed by averaging confusion matrices from four leave-one-out (LOO) comparisons among annotators.
Figure 6: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four LOO comparisons, where each annotator is compared to the consensus of the others.
Both models exhibit notable differences from human annotators. In particular, both models frequently annotate true Solidarity instances as None (108 instances for GPT-4, 128 for Llama) and true Anti-solidarity instances as None (59 for GPT-4, 60 for Llama), especially for common subtypes such as group-based and compassionate. These errors are far less frequent in human annotations (32 and 12, respectively), suggesting that models may more often underestimate both solidarity and anti-solidarity. Llama additionally shows a tendency to misclassify true Solidarity as Anti-solidarity (30 cases vs. 11 for GPT-4 and 13 for human annotators) and, to a lesser extent, true Anti-solidarity as Solidarity (18 vs. 4 for GPT-4 and 10 for humans). Some of these misclassifications arise when a speaker cites an anti-solidarity statement by another party, which Llama occasionally misinterprets as anti-solidarity (see Examples 9 and 11 in Table 14 in the Appendix; these are discussed in more detail in Section 4.3. However, manual inspection indicates that such cases are rare.
4 Temporal Trends in Solidarity Towards Migrants
In this section, we apply the best-performing open-weight model identified in 3, Llama-3.3-70B, to the full Migrant dataset of 63k instances to analyze trends across time. We focus on the Migrant dataset in particular because migration has long been one of the most contested issues in German politics, consistently framed through ?crisis narratives? that have shaped national identity and migration governance across decades frohlich2025migration (36). This makes it a compelling case for evaluating whether model-based annotations align with historical and political science literature. It is also reflected in our human-annotated dataset, where migrant-related instances display a more diverse label distribution, including substantially more anti-solidarity, compared to the predominantly solidarity-oriented discourse on women. More specifically, 4.1 compares the long-term (anti-)solidarity trends across human annotators and LLMs and thereby further validates the results of our models. In doing so, we also note two periods with notable shifts in solidarity, namely the post-war period and the recent past, which we cover in more detail in Section 4.2 and Section 4.3.
(a) GPT-4
(b) Llama-3.3-70B
Figure 7: Solidarity trends for Migrant predicted by GPT-4 and Llama-3.3-70B. Values are normalized within each decade; percentages may not total 100% because instances classified as ?None? are not shown (though included in the calculation). The period 1933–1949 (grey shaded area) is excluded due to limited data availability during and immediately after the Nazi dictatorship.
4.1 Trends with GPT-4 vs with Llama-3.3-70B
(a) GPT-4
(b) Llama-3.3-70B
Figure 8: (Anti-)solidarity frames trends for Migrant predicted by GPT-4 and Llama-3.3-70B. Percentages are normalized within each decade across all eight subtypes; thus, the curves in both panels together sum to 100% per decade. The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
Figures 7 and 8 show the development of (anti-)solidarity for GPT-4 and Llama-3.3-70B on a high and fine-grained level across 155 years. The GPT-4 reference data is taken from kostikova-fine (57), based on a subset of 18.3k migrant-related instances (from an original 58k-instance dataset; see 2.1) spanning 1867–2022. While Llama-3.3-70B is applied to the full dataset (expanded to 63k instances through 2025) for the main analysis, we restrict it to the same 18.3k-instance subset for this comparison.
To validate the annotations of the models, we first compare them to each other. Comparing the trends between GPT-4 and Llama-3.3-70B, we observe a high correlation of 0.96 on the high level and between 0.77 and 0.95 across the more frequent fine-grained subtypes (see Table 4 in the Appendix). This indicates that the models show a high degree of agreement in temporal trends. [^5] Additionally, Llama-3.3-70B shows near-identical trends between the 18.3k subset and the full 63k dataset (e.g., $\rho$ = 0.996 for solidarity and anti-solidarity, and $\sim$0.97–0.98 for subtypes).
For reference, Figure 3 shows the high-level trends across time for human annotations ($N=2,084$). We observe similar relative distributions of (anti-)solidarity on a high level. Direct trend correlations between Llama-3.3-70B and human annotations are also substantial ($\rho$ = 0.84 for anti-solidarity, $\rho$ = 0.76 for solidarity, $\rho$ = 0.47 for mixed). In terms of Cohen’s $\kappa$ (computed on the 2,084 human annotated instances), GPT-4 slightly outperforms Llama-3.3-70B ($\kappa$ = 0.57 vs. 0.50 at the high level; 0.38 vs. 0.36 for subtypes), but differences are modest. Overall, GPT-4 and Llama-3.3-70B appear to produce similar results that agree to a moderate level with human annotations. Given that it is an open weight model that we can run on our own computing hardware at low further cost, we utilize Llama-3.3-70B for the subsequent analysis.
4.1.1 Trend Stability Across Keywords.
Before we move on to interpreting the observed trends derived from Llama-3.3-70B predictions, we verify that they are not artifacts of our keyword-based data collection (see 2.1). As Figures 19 and 20 in the Appendix show, individual keywords vary considerably in their temporal distribution and in the balance of solidarity vs. anti-solidarity. To rule out the possibility that our long-term trends merely reflect such keyword effects, we perform two robustness tests.
First, we conduct a stability test by randomly sampling 200 subsets of at least 5 keywords each and comparing the resulting high-level trends using Pearson’s correlation coefficient. Across all 19,900 pairwise comparisons, the trends are highly consistent (average correlation of 0.92 for solidarity, 0.81 for anti-solidarity, and 0.94 for mixed; with 25% quantiles of 0.90, 0.70, and 0.92, respectively). [^6] For more detail on keywords and the stability test, see A. Second, we restrict the analysis to the two most frequent and temporally widespread keywords, Ausländer (en.: foreigners) and Flüchtlinge (en.: refugees), which occur in every decade (19.7k and 12k mentions, respectively). Together, the resulting plots (Figure 23 in the Appendix) and the stability test results confirm that overall trends in solidarity and anti-solidarity are preserved both under this restricted vocabulary and across different keyword selections.
4.1.2 Notable Trends in 1867-2022.
Inspecting the trends in Figure 7, we observe four periods with notable shifts in solidarity: Between 1890-1920, in the 1930s, in the post-war period, and in the recent past after 2010.
Regarding the 1890-1920 period, we observe some disagreement between the models: GPT-4 shows an increase in solidarity between 1900 and 1920, while Llama-3.3-70B does not. We also note that the agreement between human annotators and between models and annotators is lower for historic data, such that we do not consider this trend to be sufficiently reliable for further analysis.
The 1930s and 1940s data points are outliers, as the parliament was swiftly sidelined by the Nazi regime and the only 109 available samples from 1933 to 1949 mostly concern inter-state relations (mentioning, e.g., ?foreign creditors?) that fall into the ?None? category. Hence, we exclude these data points from further analysis, as well.
For the two remaining periods, 1949–1957 and 2009–2025, we take a more detailed look. In the postwar years (Section 4.2), parliamentary discourse reveals high levels of solidarity and little anti-solidarity toward migrants, despite the severe economic and infrastructural strain of the time braun2014immigration (14). This raises our first guiding question: Why is there so much solidarity (and so little anti-solidarity), and what was it based on? In contrast, the recent period (Section 4.3) shows a clear rise in anti-solidarity and a decline in solidarity, beginning in the early 2010s. These shifts likely reflect the political impact of recent refugee movements and the growing influence of anti-immigration views, such as those promoted by the far-right party Alternative für Deutschland (AfD) (en.: Alternative for Germany), which has gained substantial electoral support in Germany since its founding in 2013. This leads to our second question: What are the drivers of anti-solidarity in the contemporary period?
4.2 Germany after the Second World War, 1949-1957
The Second World War set off the largest wave of forced migrations in European history, with tens of millions of people, both non-Germans and ethnic Germans, displaced or resettled across Central and Eastern Europe schulze2017commemoration (90). The populations displaced by the war were highly heterogeneous, including refugees, displaced persons (DPs), forced laborers, and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans. For those who settled in postwar Germany, integration posed significant political and societal challenges, shaped by regional disparities, economic hardship, and tensions with the local population chevalier2024forced (21).
This subsection focuses on the early postwar period (1949–1957), when the Federal Republic of Germany institutionalized solidarity as part of nation-building. This period is analytically important because solidarity was no longer merely a moral gesture, but a tool of governance, embedded in legal frameworks like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy kittel2020stiefkinder (54). Crucially, such measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups from the benefits provided by these laws. However, since solidarity in this period was neither static nor uniform, shaped by diverse migrant groups, competing political ideologies, and contested notions of deservingness ther1996integration (96), we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97) to guide our analysis. By distinguishing between identity-, need-, and contribution-based forms, this framework allows us to examine how solidarity was constructed and sustained amid institutional strain.
4.2.1 Overall solidarity distribution
Throughout the 1949–1957 period, solidarity dominated parliamentary discourse on postwar population movements. As shown in Figure 9, [^7] We use raw counts as they provide a more interpretable view of stance distribution during this period, given the overwhelming dominance of solidarity and the low frequency of categories like anti-solidarity. mentions of solidarity consistently outpaced both anti-solidarity and mixed statements, at times reaching ten times their frequency.
Figure 9: Distribution of (Anti-)solidarity mentions in Bundestag debates, 1949–1957 (raw counts). LAG refers to the Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG to the Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG to the Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
Mentions of solidarity closely track key legislative moments, particularly in the debates leading up to the passage of major laws. A major spike in March 1950 aligns with early discussions on the integration of expellees and Article 131, which addressed the rehabilitation of former public servants of the Nazi regime, many of them displaced. A second cluster from January to May 1952 reflects debates on the LAG, preceding its adoption in May (see an excerpt from such a debate in Example 1 in Table 10 in the Appendix). Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified rights and integration measures, see Example 2 for such a debate. Later peaks in September 1954 and June 1955 reflect continued debate over LAG implementation and budget issues. Smaller but sustained increases in 1956 relate to new support measures for Sowjetzonenflüchtlinge (en.: Soviet-zone refugees) and Aussiedler (en.: resettlers).
Figure 10: Distribution of (Anti-)solidarity frame subtypes from 1949 to 1957 (raw counts). An inset plot shows the evolving composition of solidarity subtypes, normalized within each 6-month period. Values represent the proportional share of each subtype among all solidarity mentions during that period.
The prominence of solidarity in this period is unsurprising. Parliamentary debates used it not only to justify compensation, but also to navigate postwar fragmentation and competing claims to victimhood and belonging hughes1999shouldering (49). This multifaceted nature of solidarity is evident in the distribution of all four subtypes across key legislative moments, as shown in Figure 10. Group-based and compassionate frames were especially prominent, reflecting an emphasis on both shared national belonging and moral obligation toward displaced populations (see, e.g., Example 1 in Table 10, where solidarity is tied to the right to the homeland and framed as a legal entitlement). More striking is the shift in the balance between these two subtypes over time. As shown in the inset of Figure 10, group-based solidarity dominated in the early 1950s but declined after mid-1953, while compassionate steadily increased. This suggests a rhetorical transition: once legal belonging was formalized, most notably with the LAG in 1952 and BVFG in 1953, solidarity no longer needed to be framed primarily in terms of shared identity. The shift may also reflect demographic change within the displaced populations, as newer arrivals like Sowjetzonenflüchtlinge from the GDR and Aussiedler (en.: resettlers) from the Eastern Bloc countries were less easily framed as co-nationals boke1996fluchtlinge (13). This trend aligns with Thijssen’s reading of Durkheim: as group-based solidarity extends to more diverse groups, it gives way to need-based appeals thijssen2012mechanical (97) (see Example 3 in Table 10 in the Appendix).
These patterns show that solidarity was used in parliamentary debates to manage the challenges of postwar reconstruction, as reflected in key legislative reforms. While the claimant groups of the broadest measure LAG were highly diverse (bombed out, evacuated, expelled, financially damaged), their eligibility for compensation under the law was based on Volkszugehörigkeit (en.: German ethnicity), excluding many from formal recognition kittel2020stiefkinder (54, 73). The concept of Volksdeutsche was a fluid and constructed Nazi category whose ambiguous definition nonetheless had concrete and deadly implications bergen2005tenuousness (9). We therefore examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in parliamentary discourse in that period.
4.2.2 Solidarity Frames Across Keywords in 1949-1957
This analysis includes only migrant-related keywords that appear at least 200 times in combination with a solidarity, anti-solidarity subtype or mixed stance. Keywords dominated by none or with low total counts are excluded (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants), etc.); see Table 5 for full distributions.
Figure 11 shows the distribution of (anti-)solidarity frames across migrant-related keywords in 1949–1957. Across all groups, solidarity was far more common than anti-solidarity, with group-based and compassionate frames dominating. However, the distribution of subtypes varies between groups.
Vertriebene (en.: expellees) are most often associated with group-based solidarity, accounting for around half of all expressions with a stance, while also receiving a substantial share of compassionate solidarity (nearly 30%). This reflects their dual position: as key representatives of German victimhood (alongside prisoners of war) during and after the World War II moeller2001war (71), they were both integrated into West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice. Still, around 5–6% of stance expressions toward Heimatvertriebene fall into the anti-solidarity category, particularly group-based (solidarity limited to ?us?, the in-group). These expressed concerns that expellees strained local infrastructures and were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment kossert2008kalte (56). Perceptions of injustice were heightened by the uneven distribution of expellees across West German states braun2014employment (15), particularly in regions where their prominent presence in public institutions and politics sparked fears of Überfremdung (en.: overforeignization) and contributed to a sense of unequal burden-sharing levy2005memories (64) see Example 4 in Table 10.
Figure 11: Relative distribution of (Anti-)Solidarity subtypes across selected migrant-related keywords, 1949–1957. Bars represent the proportion of each subtype within each keyword group, thus all subtypes for a given group sum to 100%.
Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are primarily framed through compassionate solidarity (45–52%), followed by group-based solidarity (25–30%). The label Flüchtlinge initially functioned as a broad catch-all category in the late 1940s nachum2018semantics (74). With the 1953 BVFG, it was more clearly differentiated into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler. Yet their Germanness was not valued equally: while Vertriebene were framed as victims of expulsion and central to West German nation-building, the label Flüchtlinge retained post-war connotations of marginality and foreignness nachum2018semantics (74). This helps explain a higher share of compassionate solidarity, as well as the notable anti-solidarity (around 10%). Sowjetzonenflüchtlinge, in turn, were primarily framed as Cold War refugees fleeing communist oppression and thus not integrated into the nation building narrative like the Vertriebene, prompting more compassionate than group-based solidarity nachum2018semantics (74, 13) (see Example 5).
Ausländer (en.: foreigners) were the main targets of anti-solidarity. Not a single coherent group, they were a discursively constructed ?other?, including DPs, Eastern European border crossers, wartime laborers, and early foreign workers smith1994germany (93). Often contrasted with German groups like Vertriebene or Einheimische, they were framed as non-belonging, economically burdensome, and sometimes criminalized holian1945missing (44) (see Example 6). However, despite this exclusionary framing, solidarity with Ausländer, especially group-based, exceeded anti-solidarity. This might be partly due to the category’s fluidity, which included ethnic Germans not yet formally recognized as such ruhkopf2023institutionalisierte (86) in the early postwar years. Politicians like Lukaschek also linked DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community ruhkopf2023institutionalisierte (86) (Example 7). The second-highest solidarity subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy within the Western alliance ruhkopf2023institutionalisierte (86) (Example 8). The 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure, granted DPs civic and social rights largely to meet conditions for sovereignty and improve the country’s international standing holian1945missing (44, 86). At the same time, instances of genuine solidarity toward non-ethnic German foreigners, particularly in CDU and SPD, statements, show that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 7 and 9).
In sum, solidarity was expressed differently depending on the migrant group and was shaped by perceived ?deservingness?. Vertriebene were largely included as part of the national community, while others, especially Ausländer, were treated as outsiders and received comparatively less support. These differences reflected not only group characteristics but also the ideological strategies of political parties. The next subsection examines how different parties in the early Bundestag framed (anti-)solidarity in their parliamentary speech.
4.2.3 Distribution across parties
Figure 12 shows the distribution of solidarity and anti-solidarity subtypes across parties for the period 1949–1957. While solidarity was the dominant stance overall, parties varied in how they framed it, driven by ideological and strategic choices. Some parties used solidarity to promote national integration, while others employed it to defend regional concerns. Broadly, we divide parties into three categories: mainstream governing parties, expellee advocacy parties, and smaller parties with ideological or regional agendas.
The main parties – governing CDU/CSU and FDP, and opposition SPD – supported the integration of expellees and refugees through legal and institutional measures, primarily expressing group-based or compassionate solidarity (see Figure 12). The CDU/CSU played a leading role in proposing the LAG, which framed solidarity as a national obligation toward fellow Germans, and even included expellee representatives in the party ranks hughes1999shouldering (49). The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers connor2018refugees (22), resulting in occasional anti-solidarity statements. The FDP, while formally aligned with CDU/CSU and SPD on key refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation connor2018refugees (22) (see Example 10 in Table 10, where FDP speakers argued that expellees should build their own homes through self-help and cooperative effort, with state support limited to loans and low-interest credit).
Expellee advocacy parties (GB/BHE and DP) [^8] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights) and Deutsche Partei (DP) (en.: German Party) explicitly positioned themselves as political representatives of displaced Germans. Following multiple successes in state elections, GB/BHE entered the federal government coalition in 1953, thereby exerting significant pressure on other parties to address the demands and concerns of expellees connor2018refugees (22). They framed solidarity primarily in terms of historical injustice and economic entitlement, calling for wealth redistribution, fair pensions, and affordable housing. Their rhetoric relied heavily on group-based solidarity, portraying expellees as full members of the national community who deserved compensation and full integration (see Example 11). Eventually, both major parties of the Federal Republic CDU and SPD managed to absorb the demands and concerns of the expellees, which made GB/BHE and DP less successful connor2018refugees (22, 54).
Regionalist parties such as the WAV and BP [^9] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized group-based solidarity, but restricted it to ethnic Germans. The WAV, a short-lived nationalist party, appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories kranenpohl1998zwischen (59) (Example 12). The BP, responding to Bavaria’s disproportionate refugee burden, called for collective responsibility in crisis, fairer redistribution and showed preference for groups like the Sudeten Germans braun2020101330 (16) (see Example 13,where the BP argued that Sudeten and Southeast Germans should be concentrated in Bavaria because they were ?closest to the Bavarian cultural sphere? and thus easier to integrate). Both parties also expressed the highest levels of anti-solidarity, either by excluding non-German migrants (WAV) connor2018refugees (22) or opposing specific groups such as Silesians (BP).
Figure 12: (Anti-)solidarity frames distribution across major parties in the Bundestag from 1949 to 1957, by election period. Percentages are normalized within each party, such that all subtype bars (solidarity and anti-solidarity) together sum to 100% per party. We include only those parties with a sufficient number of stanced instances $(N>100)$. The full label distribution across all parties is provided in Figure 6 in the Appendix.
The KPD [^10] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) framed displaced populations through a lens of socialist integration, emphasizing group-based solidarity to support their incorporation into the new social order schwartz2008vertriebene (91). This solidarity, however, was assimilationist rather than inclusive: the party avoided the term Vertriebene, in line with Soviet repression of expellee identity heidemeyer2007deutsche (42), and this was reflected in its minimal use of empathic solidarity. The KPD referred almost exclusively to ?refugees?, advocating land reform, compensation, and integration as part of a broader effort to portray them as victims of fascism and capitalism and to gain political support connor2018refugees (22) (see Example 14, where the KPD argued that refugee employment could not be solved under capitalism).
Some parties active in the early postwar years, such as the KPD, WAV, and BP, did not stay in parliament for long, but they played an important role in shaping political debates of the time. Rather than merely reflecting existing social tensions, as suggested by lipset1967cleavage (67), these parties actively contributed to defining what solidarity meant and who was considered deserving of support. As suggested by thijssen2022s (98), such framings were shaped by the political and social challenges of the period and reflected broader efforts to maintain or renegotiate social cohesion.
4.3 Recent German history 2009-2025
Between 2009-2015, migration into Germany increased from just under 800,000 people per year to over 2 mio. people per year gessler2023politicization (41) in the context of the war in Syria. The peak in migration in 2015 has since been associated with the label ?migration crisis? gessler2023politicization (41, 36). Since then, the right-wing AfD has made electoral gains and entered the German federal parliament in 2017. Another spike in migration occurred in 2022 due to the Russian invasion of Ukraine bamf2025 (33). Alongside these historical events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
4.3.1 Overall solidarity distribution
Fig. 13 shows the (anti-)solidarity trends according to Llama-3.3-70b annotations between 2009-2025. We observe a decrease in solidarity between 2014-2018, a somewhat stable period between 2018-2022, and another decrease in solidarity 2022-2025. The decrease in solidarity is associated with an almost corresponding increase in anti-solidarity In terms of (anti-)solidarity subtypes, we observe a peak in compassionate solidarity in 2014, followed by a sharp decrease. The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity. In terms of anti-solidarity, we observe a strong increase in group-based anti-solidarity, accounting for over 50% of all migrant-related statements that either expressed solidarity or anti-solidarity. Exchange-based anti-solidarity exhibited a small increase between 2015-2020 but dropped again with the beginning of the last parliamentary period in 2021.
Overall, the 2020s mark the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements. This key finding is not only visible in the data provided by Llama-3.3-70B but also in human annotations (Figure 2). To provide more insight into this finding, we contextualize it with recent Germany history and a deeper look into party distribution.
(a) Solidarity and anti-solidarity distribution
(b) Solidarity and anti-solidarity frames distribution
Figure 13: Fig. 13(a) shows the fraction of solidarity, anti-solidarity, and mixed stance towards migrants from 2009 to 2025 (instances classified as ?None? are not shown, though included in the calculation). Fig. 13(b) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes according to Llama-3.3-70B, where percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year).
The spike in compassionate solidarity 2014 coincides with a time of heightened public solidarity with Syrian refugees: 15.4% of German survey respondents indicated solidarity as the most important personal value in 2014, up from 12.3% in 2013 (DeLeur2023, 23, p. 38). Syrian refugees were commonly framed as needing help, motivating calls for a ?German welcome culture? Gebauer2023 (39) (see Example 15 in Table 14 in the Appendix). Over the course of 2015, a counter-narrative emerged, framing the influx of refugees as overburdening German capacities and as a ?refugee crisis? Gebauer2023 (39) (see Example 16). During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests and moved notably right, followed by substantial electoral gains in both state and national elections and the AfD entering the German parliament in 2017 Frey2020 (35, 39). Regarding the broader European context, DeLeur2023 (23) diagnose a ?solidarity fatigue? across Europe after 2015, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level. The counter-narratives mentioned by Gebauer2023 (39, 35, 23) match several notions of anti-solidarity according to thijssen2022s (98): refugees as an outgroup that stands in competition to the German population (group-based anti-solidarity), refugees as taking away resources and not contributing sufficiently (exchange-based anti-solidarity), and refugees as not deserving of compassion (compassionate anti-solidarity); see Example 17 where refugees are described as not contributing, straining resources, and threatening the welfare of the ?real? German population. We also note a general shift towards right-wing sentiments in the general German population Zick2023 (108), which may be associated with heightened anti-solidarity.
More recently, the Russian invasion in Ukraine and the subsequent migration of Ukrainian refugees into Germany triggered another spike in public solidarity, albeit much less prominent Weber2024 (104). In our data, solidarity increases slightly in 2022, followed by a continued, rapid decline. Bojarskich2025 (12) identified several factors promoting anti-solidarity towards Ukraine, such as conspiracy mentality and populist voting, but also fears about economic and societal consequences, indicating exchange-based (insufficient contribution) or group-based (competition between groups) anti-solidarity (Example 18). Additionally, anti-Ukrainian framing in parliamentary debates may have been reinforced by the AfD’s long-standing pro-Russian alignment arzheimer2023russia (5), including opposition to sanctions and calls to end German military aid to Ukraine.
In contrast to the postwar period, (anti-)solidarity shifts in the German parliament between 2009 and 2025 appear less attached to specific legislative initiatives. Because of Germany’s membership in the European Union, migration policy has become more complex and is negotiated between European institutions and member states geddes2020migration (40). Additionally, the AfD integrated anti-migrant talking points across a wide range of legislative topics. As such, we do not look deeper into specific laws passed during this time frame but rather focus on the trends for specific parties.
Figure 14: Raw yearly counts of parliamentary statements by selected political parties, 2009–2025.
4.3.2 Distribution across parties
(a) AfD
(b) CDU/CSU
(c) SPD
(d) FDP
(e) Die Linke
(f) Die Grünen
Figure 15: Distribution of solidarity and anti-solidarity subtypes over time (2009–2025) for selected German parties. For each party, percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year). Note that AfD was first represented in the Bundestag in 2017, and the FDP was not represented from 2013 to 2017. Therefore, data points are absent for these parties during those respective periods.
Fig. 14 shows the raw numbers of migration-related statements per party and year, Fig. 15 shows the fine-grained solidarity trends.
The right-wing AfD
utters almost no expression of solidarity and the anti-solidarity subtypes remain relatively stable over time, with group-based anti-solidarity being consistently dominant (ca. 60% in 2018-2019 and ca. 75% in 2021-2025), although a notable number of samples with exchange-based solidarity (ca. 25%) occurs in the parliamentary period 2017-2021. Although the AfD is not the largest faction in parliament, they account for the largest share of anti-solidarity samples overall (72.5%). The sharp rise of anti-solidarity after 2017 (Figure 13(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure 14). This can be explained by the fact that many AfD speeches, across topics, contain references to migrants illegally crossing the border, criminality of migrants, or migrants imposing additional burden on the German society leaving less resources to address the actual topic of legislative debate (see Example 19 in Table 14 in the Appendix where an AfD member argues that refugees displace elderly people from care homes).
The center-right CDU/CSU and the liberal FDP
changed the most in our data. In the period 2009-2015, we observe a clear majority of solidarity samples, with a spike of compassionate solidarity for the CDU/CSU in 2014 (the FDP was not in parliament at that time). The main difference between both parties is more emphasis on exchange-based solidarity in the FDP, e.g. the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: skilled labor migration law) in 2023 (see Example 20, where the FDP highlights the law as enabling foreign workers to contribute to the German economy). However, for both parties, we observe a rise in group-based anti-solidarity, accounting for 80% of all samples in 2025. When inspecting the samples, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, which is classified as group-based anti-solidarity by Llama-3.3-70B (see Example 21 where a CDU/CSU speaker frames migration as a threat to political stability and an excessive burden on local infrastructure, with calls for deportations, faster procedures, and rejections at the border; and Example 22 where an FDP speaker portrays asylum as being misused for state-subsidized residence).
The drastic shift toward group-based anti-solidarity over the last three years is insufficiently explained by recent literature. Partially, it may be that politicians from CDU/CSU and FDP take over populist rhetoric from AfD politicians who entered parliament since then Esguerra2023Accommodation (31). However,Lewandowsky2022 (65) and kesting2018discourse (53) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality. We also note that the first shift toward group-based anti-solidarity already occurred before the AfD enters parliament in 2017. Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population Zick2023 (108) by appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD. Overall, we cannot establish a full explanation but encourage further research into this striking and rapid trend.
The center-left SPD
also exhibits a spike in compassionate solidarity in 2014 and a slow decline in solidarity towards more group-based anti-solidarity (ca. 20% in 2024). Part of this is explained by mis-classifications of Llama-3.3-70B because SPD speakers often cite anti-solidarity stances of other parties (especially AfD) and these citations are mis-classified as anti-solidarity by the speaker (see Example 23 where an SPD speaker cites AfD proposals putting foreigners under general suspicion in order to criticize them). However, there are also samples where SPD speakers refer positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 24).
The left Green and Linke parties
only display very few instances classified as anti-solidarity and almost all of them are misclassifications where Green or Linke speakers cite anti-solidarity statements or parliamentary motions by other parties (see Example 25 ). These samples are nonetheless interesting because they express a polarizing influence of the AfD entering parliament, leading left-wing parties to positioning themselves in contrast Atzpodien2022 (6). Overall, Die Grünen initially emphasized compassionate solidarity, but later shifted to group-based appeals stressing integration and rights. In the 2024/25 campaign, their rhetoric focused strongly on exchange-based arguments about migrants’ economic and societal contributions Gruene2025 (17) (see Example 26). Similarly, Die Linke began with predominantly compassionate appeals, framing migration as a humanitarian duty, but later emphasized migration as integral to society and advanced a detailed pro-immigration agenda, thereby increasing group-based solidarity. In the 2024/25 campaign, they moved further toward empathic solidarity, criticizing stereotyping and positioning themselves against exclusionary discourses DieLinke2021 (25, 26) (see Example 27). A special case is the BSW, which split off from the Linke party in 2024, partially due to explicit anti-migration positions. However, our data set contains only very few samples from this party (38 statements), so we cannot provide a reliable quantitative analysis.
5 Discussion and Conclusion
In this paper, we have evaluated LLMs on the complex task of annotating (anti-)solidarity in German parliamentary speeches and interpreted the annotated data to derive new insights into (anti-)solidarity towards migrants in the post-war period as well as recent past. In more detail, we provide both methodological and sociological insights:
Methodological perspective.
We find that instruction-tuned LLMs like Llama-3.3-70B and GPT-4 are effective tools for annotating solidarity frames in parliamentary debates, with open models offering a more cost-effective alternative to closed ones. In our experiments, Llama-3.3-70B performed competitively with GPT-4, especially in few-shot settings. Still, no model consistently outperformed GPT-4, and even GPT-4 did not fully attain the human upper bound. Existing research offers mixed evidence on the reliability of LLMs for annotation tasks: some highlight their limitations in handling subjective or expert-defined taxonomies ziems2024can (109), while others show they can substitute or outperform human coders heseltine2024large (43, 99), or suggest that smaller, fine-tuned models are more reliable in political science applications wang2024selecting (102). In our case, model performance generally improved with model size, but also varied by target group, historical period (with older data proving more difficult), and inference method. These findings highlight the absence of a one-size-fits-all solution: rather than relying on a single model, both model and method selection must be adjusted to the requirements of the specific task.
Moreover, we emphasize the need for validation at every stage of the analysis pipeline. Even keyword search as a first filtering step can introduce biases in the data or can aggregate data that concerns very different sub-populations (such as different groups of migrants in the German post-war period). Regarding the LLM annotation itself, we validated model predictions against a large data set of human annotations (aggregating a year of labor involving multiple annotators), test multiple inference configurations (including zero- and few-shot prompting), and validate long-term trends produced by our two best-performing models (Llama-3.3-70B and GPT-4) against human annotations. We also conducted error analyses to identify where and why models fail to make sure that outputs can be reliably interpreted. This aligns with recent work advocating for human-in-the-loop workflows and partial human validation when using LLMs for annotation tasks, to ensure output quality and interpretability pangakis2024keeping (75, 100).
Despite these efforts, it remains essential to acknowledge model-specific weaknesses when interpreting results. For example, we observed that Llama-3.3-70B occasionally misclassifies anti-solidarity, particularly in instances where criticism of anti-migrant rhetoric is sometimes mistaken for anti-solidarity framing itself. In such cases, models may overlook logical direction or target relevance and assign misleading labels despite otherwise coherent reasoning zhang2025mpvstance (107). One possible reason for this behavior could be the influence of complex political language or long prompt context, which may distract the model from the intended target huang2025breaking (48). Importantly, we could only reveal model errors because we checked the primary data before drawing conclusions.
Social science perspective.
Beyond the methodological contributions, our study also demonstrates how an LLM-based pipeline can support a theoretically informed quantitative analysis through two illustrative case studies: one historical (1949–1957) and one contemporary (2009–2025).
For the early post-war period (1949–1957), we demonstrated how solidarity was shaped and institutionalized in the context of national reconstruction. While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insight for historical and social science research. First, it enables us to move beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity, and shifting rhetorical strategies across different migrant groups. Second, our approach helps integrate migration histories that are often treated separately, such as expellees, DPs, and refugees, into a shared analytical framework, allowing for direct comparison of how solidarity was articulated across groups. This also sheds light on the emerging distinction between early DPs and the broader category of Ausländer, including those later identified as Gastarbeiter, showing that solidarity was not uniformly absent but selectively expressed.
For the recent past (2009-2025), we see a striking shift from solidarity to anti-solidarity following the 2015 migration crisis. This may be partially explained by solidarity fatigue DeLeur2023 (23) or may be part of a larger right-wing shift Zick2023 (108) on migration issues. While stricter migration stances across Europe have been analyzed before, our study is the first to quantify the shift from solidarity to anti-solidarity in German parliament. Our data also revealed that the shift is most striking in the liberal and conservative parties, motivating further research. This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives aktas2024rise (2), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists bayerlein2021chasing (8). Comparative survey evidence further shows that anti-immigrant attitudes remain a central driver of right-wing populist support across the continent messing2021anti (70).
This type of fine-grained, longitudinal analysis of solidarity would hardly be feasible at such scale using manual annotation alone which, in our project, cost over €29,000 for only 3,500 instances. By contrast, with access to computing resources (e.g., in our case, in-house hardware with NVIDIA A40 GPUs), open-weight models such as Llama-3.3-70B can annotate tens of thousands of instances at minimal additional cost. Even closed models like GPT-4 are far cheaper, with one third of our dataset annotated for about €500. While machine annotation is therefore not ?free?, it is vastly more scalable than human annotation and enables large-scale analyses that would otherwise be financially prohibitive. In this sense, our method bridges ?distant reading? approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded systematic interpretation of parliamentary speech while retaining full corpus coverage.
6 Limitations and Future Work
While our analysis was performed carefully to avoid methodological pitfalls, we nonetheless acknowledge remaining limitations of our study.
Human Label Variation:
Even after months of training, discussion and improvements of the annotation manual, we still observe substantial disagreement between human annotators, especially on the fine-grained level. This indicates that (anti-)solidarity is intrinsically a challenging concept to annotate, especially on parliamentary speeches, and that reasonable disagreement may persist (e.g. whether anti-solidarity should be regarded as group-based or exchange-based when competition between groups around limited resources is highlighted). For this investigation, we used the majority label, but future work should consider methods that explicitly account for human label variation and embrace it plank-2022-problem (76).
Political Bias:
It may be that the LLMs compared in this paper may be subject to political bias as investigated by several prior works motoki2024more (72, 85). In particular, it may be that a left-wing political bias leads to more anti-solidarity being annotated. However, we do not observe a systematic bias toward anti-solidarity in the error analysis (Fig. 6); instead, models often mislabel solidarity and anti-solidarity samples as ?None?. It may also be that the human annotators are biased towards certain labels, but this would need to affect all annotators equally because we do not observe a systematic disagreement towards one label, either. Finally, even if bias were present, this would not explain differences and trends in (anti-)solidarity (which are the main subject of our analysis) but rather the overall level of (anti-)solidarity.
Analysis of (Anti-)solidarity towards women:
In the same way we performed an analysis of (anti-)solidarity trends towards migrants, the (anti-)solidarity towards women could be analyzed. For example, recent work by Diabate2023 (24) suggests that there may be a shift to the left in Germany regarding gender, which would be an opposite trend compared to what we found regarding migration. Due to space and time constraints, such an analysis is not included in this work but appears highly promising for future work.
Further model variations:
We considered a range of models as well as different prompts, and fine-tuning (for Llama models and BERT models) but a vast range of possible model variations remains to be explored, such as ensembles of models, less quantized models, newer models, etc. It may well be that human annotation quality can be fully matched with more refined tools.
Temporal Reliability:
Both model performance (macro F1) and human agreement vary over time, with notable dips in some decades (e.g., the 1960s for Llama-3.3-70B; Fig. 16). This reflects the difficulty of annotating historical language, but overall trend correlations between GPT-4, Llama-3.3-70B, and human annotations remain high. Since our analysis emphasizes broad temporal trends rather than isolated data points, these fluctuations do not undermine the validity of our conclusions.
Appendix
Appendix A List of Keywords
For Frau (en.: woman) we use Frauen, Frau, Mutter, Mädchen, Mütter, Ehefrau, Müttern, Hausfrauen, Hausfrau, Ehefrauen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Großmutter, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
Since Frau is the German word for woman but also for Mrs./Ms., we only include occurrences of the word Frau that are not followed by a capitalized word (which are probably surnames).
For Migrant (en.: refugee) we use Flüchtlinge, Ausländer, Flüchtlingen, Zuwanderung, Vertriebenen, Ausländern, Asylbewerber, Migranten, Migration, Heimatvertriebenen, Aussiedler, Einwanderung, Ansiedler, Vertriebene, Zuwanderer, Asylbewerbern, Flüchtling, Heimatvertriebene, Sowjetzonenflüchtlinge, Aussiedlern, Einwanderer, Asylsuchenden, Asylsuchende, Bürgerkriegsflüchtlinge, Zuwanderern, Ansiedlern, Migrantinnen, Vertriebener, Emigranten, Kriegsflüchtlinge, Ausländerinnen, and Immigranten.
When running stability tests (see Subsubsection 4.1.1), we ensure that each sample includes at least 5 keywords, covers at least 10% of the dataset, and spans at least 75% of the timeline, such that enough data is present to create the plots. For the analysis of frequency of keywords over time, we calculate the percentages normalized for each keyword, i.e., a value of pp% in year $y$ implies that in year $y$ pp% of all sentences with this keyword occurred. The trends are shown in the Fig. 19 in the Appendix.
Appendix B Parties
For speaker and party extraction, we parse the XML protocols of the Bundestag and Reichstag. For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by poradaautomatische (77).
Each new intervention is marked in the XML by a <redner> element containing both the speaker’s name and their party. During preprocessing (XML parsing, transcript cleaning, and sentence segmentation), we convert these into standardized <SPEAKER> tags and extract the associated party information. During extraction of migrant- or women-related sentences, the most recent <SPEAKER> marker (with party) is assigned to each sentence until a new one appears.
List of the parties included in the dataset, along with the variations of their names or abbreviations as they have been recorded: AfD (Alternative for Germany); Die Linke (The Left) with variations such as PDS, Gruppe der PDS; Bündnis 90/Die Grünen (Alliance 90/The Greens); CDU/CSU (Christian Democratic Union/Christian Social Union); SPD (Social Democratic Party of Germany); FDP (Free Democratic Party); DP (German Party) with variations such as DP/DPB, DP/FVP, FVP; GB/BHE (All-German Bloc/League of Expellees and Deprived of Rights); KPD (Communist Party of Germany); BP (Bavarian Party); WAV (Economic Reconstruction Union).
Appendix C Models Inference and Training Details
Hardware and Inference Settings.
All inference runs were conducted on an NVIDIA A40 (80GB) with temperature set to 0.6 and top-p to 0.9. For Mistral-Large-Instruct-2407, we used an NVIDIA A100 to accommodate memory requirements.
Prompt Format.
Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure 24 for Migrant and in Figure 27 in the Appendix for Woman).
In few-shot settings, one representative example is included for each class label. GPT-4 receives all instructions in a single prompt (see the prompt text in Figure 21 for Migrant and Figure 22 in the Appendix for Woman). Open-weight models receive separate prompts: one for high-level classification, and one for subtype prediction if the label is solidarity or anti-solidarity.
Baseline fine-tuning details.
For our baseline, we use a BERT-based pipeline with 110M parameters, comprising a high-level category classifier and two subtype models for solidarity and anti-solidarity. All models share a similar architecture, processing inputs with a target (Frau or Migrant) and full text comprising the focus sentence and left and right context. We add a fully connected layer with softmax activation atop the pooled output of the BERT-based models, with 4 output units for each model. To counter class imbalance, minority classes are oversampled to parity with the majority class. We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer AdamW (68) and categorical crossentropy loss.
Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 reimers2019sentence (82). We modify the network by expanding the embeddings from 512 to 1024 dimensions, then reducing them back to 512 dimensions. The model incorporates ReLU activation and dropout (0.1), and maps the embeddings to four and ten target classes for high-level and fine-grained tasks, respectively. To manage class imbalance, we use oversampling. The finetuning is conducted over 20 epochs with a batch size of 8, starting with a learning rate of 5e-4, a warmup ratio of 0.03, and using AdamW for optimization with learning rate decay and categorical crossentropy.
Fine-Tuned Llama Models
We fine-tune Llama-3-8B-It on a single A40 GPU. Llama-3-70B-It is trained across two A40s using PyTorch FSDP and QLoRA [^11] https://www.philschmid.de/fsdp-qlora-llama3. Training spans 3 epochs with a constant learning rate of 2e-4, AdamW optimizer, and max gradient norm of 0.3.
Label Woman Migrant Total per label
Group-based Solidarity 112 (7.8%) 270 (12.4%) 382 (10.6%)
Exchange-based Solidarity 54 (3.8%) 79 (3.6%) 133 (3.7%)
Empathic Solidarity 125 (8.7%) 32 (1.5%) 157 (4.3%)
Compassionate Solidarity 732 (51.0%) 692 (31.7%) 1424 (39.3%)
Solidarity (no subtype) 41 (2.9%) 55 (2.5%) 96 (2.7%)
Total Solidarity 1064 (74.1%) 1128 (51.6%) 2192 (60.5%)
Group-based Anti-solidarity 10 (0.7%) 288 (13.2%) 298 (8.2%)
Exchange-based Anti-solidarity 0 (0.0%) 55 (2.5%) 55 (1.5%)
Empathic Anti-solidarity 17 (1.2%) 6 (0.3%) 23 (0.6%)
Compassionate Anti-solidarity 7 (0.5%) 97 (4.4%) 104 (2.9%)
Anti-solidarity (no subtype) 5 (0.3%) 19 (0.9%) 24 (0.7%)
Total Anti-solidarity 39 (2.7%) 465 (21.3%) 504 (13.9%)
Mixed 60 (4.2%) 149 (6.8%) 209 (5.8%)
None 273 (19.0%) 442 (20.2%) 715 (19.8%)
Total 1436 (100.0%) 2184 (100.0%) 3620 (100.0%)
Table 2: Distribution of labels in the human-annotated dataset for Woman and Migrant instances. Percentages are relative to each column total.
Figure 16: Macro F1 scores over time for each model (best configuration), and average pairwise Cohen’s Kappa for human annotators. Evaluated on the full test set, including Frau and both test sets for Migrant, grouped by decade.
Figure 17: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
Figure 18: Fine-grained confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
Table 3: F1 scores per model for fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman. The top section contains fine-grained categories, and the bold rows show performance on high-level classification. Scores are chosen for the best performing configuration for the model in terms of Macro F1.
Label
GPT-4
Mistral
2-Large
Llama
3.3-70B
Llama
3-8B
Gemma
2-27B
Gemma
2-9B
BERT SBERT
Group-based Solidarity 0.33 0.30 0.29 0.25 0.25 0.34 0.32 0.16
Exchange-based Solidarity 0.56 0.39 0.36 0.25 0.41 0.20 0.21 0.19
Empathic Solidarity 0.04 0.27 0.10 0.08 0.15 0.14 0.00 0.03
Compassionate Solidarity 0.48 0.42 0.51 0.22 0.45 0.43 0.51 0.41
Solidarity (high-level) 0.78 0.52 0.69 0.60 0.58 0.63 0.66 0.51
Group-based Anti-solidarity 0.47 0.25 0.42 0.32 0.42 0.41 0.28 0.20
Exchange-based Anti-solidarity 0.46 0.29 0.35 0.12 0.28 0.29 0.09 0.22
Empathic Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Compassionate Anti-solidarity 0.00 0.18 0.07 0.00 0.19 0.11 0.17 0.12
Anti-solidarity (high-level) 0.64 0.53 0.52 0.44 0.56 0.54 0.35 0.35
Mixed 0.38 0.26 0.43 0.10 0.17 0.19 0.24 0.15
None 0.67 0.60 0.63 0.56 0.57 0.58 0.59 0.51
Label
GPT-4
Mistral
2-Large
Llama
3.3-70B
Llama
3-8B
Gemma
2-27B
Gemma
2-9B
BERT SBERT
Group-based Solidarity 0.28 0.24 0.25 0.00 0.20 0.20 0.24 0.08
Exchange-based Solidarity 0.31 0.00 0.00 0.00 0.22 0.30 0.35 0.15
Empathic Solidarity 0.27 0.36 0.41 0.24 0.33 0.41 0.22 0.14
Compassionate Solidarity 0.43 0.32 0.51 0.00 0.37 0.21 0.62 0.52
Solidarity (high-level) 0.85 0.67 0.79 0.51 0.75 0.70 0.77 0.78
Group-based Anti-solidarity 0.31 0.36 0.22 0.00 0.27 0.26 0.00 0.00
Exchange-based Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Empathic Anti-solidarity 0.67 0.00 0.25 0.07 0.00 0.17 0.00 0.00
Compassionate Anti-solidarity 0.50 0.33 0.22 0.36 0.15 0.00 0.00 0.00
Anti-solidarity (high-level) 0.65 0.44 0.39 0.31 0.48 0.35 0.12 0.10
Mixed 0.30 0.40 0.37 0.16 0.32 0.30 0.15 0.26
None 0.62 0.54 0.47 0.41 0.66 0.41 0.21 0.40
Table 4: Agreement between GPT-4, Llama-3.3-70B, and human annotations on the Migrant dataset. Subtable LABEL:tab:model-human-agreement shows instance-level agreement; Subtable LABEL:tab:trend-agreement reports trend-level correlation across decades.
Comparison Cohen’s $\kappa$ / Macro F1 N
GPT-4 vs Llama-3.3-70B 0.47 (0.58) / — 17,696
GPT-4 vs Human 0.38 (0.57) / 0.34 (0.63) 2,084
Llama-3.3-70B vs Human 0.36 (0.50) / 0.35 (0.61) 2,084
Label Pearson (p)
Group-based Solidarity 0.560 ($p=2.41\times 10^{-2}$)
Exchange-based Solidarity 0.777 ($p=4.00\times 10^{-4}$)
Empathic Solidarity 0.108 ($p=0.691$)
Compassionate Solidarity 0.915 ($p=7.10\times 10^{-7}$)
Solidarity (high-level) 0.966 ($\boldsymbol{p=1.25\times 10^{-9}}$)
Group-based Anti-solidarity 0.945 ($p=3.65\times 10^{-8}$)
Exchange-based Anti-solidarity 0.898 ($p=2.29\times 10^{-6}$)
Empathic Anti-solidarity 0.160-0.160 ($p=0.553$)
Compassionate Anti-solidarity 0.841 ($p=4.44\times 10^{-5}$)
Anti-solidarity (high-level) 0.976 ($\boldsymbol{p=1.15\times 10^{-10}}$)
Mixed 0.685 ($p=0.0034$)
Figure 19: Distribution of all Migrant keywords over the years, normalized per keyword. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
Figure 20: Percentage of sentences showing solidarity/anti-solidarity per decade for all Migrant keywords from 1867 to 2025. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
Figure 21: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Migrant dataset.
Figure 22: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Frau dataset.
Figure 23: (Anti-)solidarity trends in 1867-2022, restricted to the keywords Flüchtlinge and Ausländer. Figure 23(a) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes predicted by Llama-3.3-70B. Percentages are normalized within each decade across all eight subtypes, so the curves in both panels together sum to 100% per decade. Figure 23(b) shows the aggregated proportions of solidarity, anti-solidarity, and mixed labels per decade; totals may not reach 100% because instances classified as ?None? are omitted (though included in the calculation). The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
(a) High-level prompt (Step 1).
Figure 24: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B. The model is first given the high-level classification prompt in Figure 27(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
(a) Subtype prompt for Solidarity (Step 2a).
Figure 25: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
(a) Subtype prompt for Anti-solidarity (Step 2b).
Figure 26: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
(a) High-level prompt (Step 1).
Figure 27: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset. The model is first given the high-level classification prompt in Figure 27(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
(a) Subtype prompt for Solidarity (Step 2a).
Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
(a) Subtype prompt for Anti-solidarity (Step 2b).
Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
Keyword Solidarity Anti-Solidarity Mixed None Total
Asylsuchende 10 0 0 5 15
Ausländer 168 52 9 525 754
Aussiedler 99 1 5 61 166
Einwanderer 23 1 0 21 45
Emigranten 66 5 0 18 89
Flüchtlinge 1828 23 84 1179 3114
Heimatvertriebene 2471 20 68 1814 4373
Sowjetzonenflüchtlinge 264 5 10 178 457
Zuwanderer 32 0 2 46 80
Zuwanderung 7 0 1 29 37
Table 5: Distribution of Solidarity, Anti-Solidarity, Mixed, and None labels per keyword included in the post-war trends analysis (Section 4.2).
BP CDU/CSU DP DRP DZP FDP GB/BHE KPD SPD WAV Z
Group-based S. 95 707 56 13 23 221 118 50 667 93 5
Exchange-based S. 5 163 5 4 1 46 14 8 73 13 0
Compassionate S. 64 670 38 9 8 135 92 69 537 72 7
Empathic S. 6 91 8 1 0 39 29 4 89 8 1
Solidarity with no subtype 1 18 0 0 0 6 2 1 6 2 0
Group-based AS 28 100 9 3 1 22 9 30 68 23 0
Exchange-based AS 4 14 3 0 0 1 3 3 11 2 0
Compassionate AS 1 3 1 0 0 0 0 3 3 0 0
Empathic AS 0 1 0 0 0 0 0 0 0 0 0
Mixed 15 63 5 1 2 23 8 4 33 3 0
None 81 1564 78 10 11 389 94 54 754 47 4
Total per party 300 3394 203 41 46 882 369 226 2241 263 17
Table 6: Distribution of (anti-)solidarity subtypes by party included in the post-war trends analysis (Section 4.2). AS stands for Anti-Solidarity; S for Solidarity.
Original Text Translation
(1) Predicted label: group-based solidarity towards expellees
(May 14, 1952)
?[…] das Lastenausgleichsgesetz und die Eingliederung der Vertriebenen [steht] in keinem Widerspruch zu unserm Anspruch und unserm Recht auf die Heimat. Im Gegenteil, […] daß die Rücksiedlung […] nur durchgeführt werden kann, wenn die körperliche, seelische und auch die wirtschaftliche Substanz der Vertriebenen erhalten geblieben ist, und das kann nur geschehen, wenn sie in die westdeutsche Wirtschaft eingegliedert werden. […] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten; […] aber der Rechtsanspruch darf nicht so behandelt werden, als sei er nur etwas Zweitrangiges, angewiesen auf das, was übrigbleibt.? ?[…] the Equalization of Burdens Act and the integration of the expellees does not contradict our right to the homeland. On the contrary, resettlement […] can only be achieved if the physical, mental, and economic substance of the expellees has been preserved, and that can only happen if they are integrated into the West German economy. […] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations; […] but the legal entitlement should not be treated as secondary, dependent only on what is left over.?
(2) Predicted label: exchange-based solidarity towards expellees
(Feb 25, 1953)
?Ich darf […] auf einige Gesichtspunkte hinweisen […] zum Bundesvertriebenengesetz. Der Ihnen vorliegende Entwurf geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen. Die Eingliederungsbestimmungen sollen der Erfüllung berechtigter Ausgleichsansprüche hinsichtlich der verlorenen Existenz dienen […] mit dem Ziel, daß sich das gesamte Arbeitspotential, das der Einheimischen und der Vertriebenen, so betätigen kann, daß eine größtmögliche Wirkung erzielt wird.? ?I would like to point out some aspects […] on the Federal Expellees Act. The draft before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy. The integration provisions are intended to meet legitimate compensation claims with regard to lost livelihoods […] with the goal that the entire labor potential, that of natives and of expellees alike, can be deployed in such a way that the greatest possible effect is achieved.?
(3) Predicted label: compassionate solidarity towards Soviet-zone refugees
(June 15, 1955)
?[…] es ist notwendig, zu den noch immer vielseitigen Problemen der Sowjetzonenflüchtlinge Stellung zu nehmen […]. Wir erkennen selbstverständlich an, daß durch die Neuregelung des Härtefonds im Lastenausgleich eine Verbesserung für die Sowjetzonenflüchtlinge erfolgte, weil die Ermessensfragen, unter denen dieser Personenkreis bislang besonders zu leiden hatte, dadurch eingeschränkt werden. […] Wenn wir aber die Sicherheit haben wollen, daß sich diese Neuregelung auch wirklich positiv für den Flüchtling auswirkt, so ist es notwendig, daß die vorgesehene Erhöhung des Härtefonds […] wirklich realisiert wird.? ?[…] it is necessary to address the still manifold problems of the Soviet-zone refugees […]. We naturally recognize that the reform of the hardship fund in the Equalization of Burdens has brought improvements […] because the discretionary decisions that this group had particularly suffered from have been curtailed. But if we want to be sure that this reform really has a positive effect for the refugee, then […] the planned increase of the hardship fund […] must actually be implemented.?
(4) Predicted label: group-based anti-solidarity towards expellees
(Dec 13, 1950)
?In unseren Landkreisverwaltungen sind heute 42,2% Heimatvertriebene. […] Auch die einheimische Bevölkerung beklagt sich […] man höre auf den Landratsämtern überwiegend ostpreußisch, und das Plattdeutsche sei zur Seite gedrängt. Unter den Lehrern […] 51% Heimatvertriebene, unter den vollbeschäftigten Richtern […] mehr als 50%, unter der Polizei sogar 61%; zweifellos eine Überfremdung, die Besorgnis erregt! Aber all diese ehrlichen Bemühungen um Eingliederung der Heimatvertriebenen sind angesichts der ungeheuren Zahl vergeblich. Arbeitslosenzahl […] beträgt unter den Heimatvertriebenen fast 60%. Die Lasten sind unerschwinglich.? ?In our district administrations today, 42.2% are expellees. The native population […] complains […] that in the district offices one mostly hears East Prussian, and that Low German has been pushed aside. Among the teachers 51% are expellees, among the full-time judges also more than 50%, among the police even 61% – undoubtedly an over-foreignization that arouses concern! But all these honest efforts to integrate the expellees are in vain in view of their immense numbers. Unemployment […] almost 60% among the expellees […]. The burdens are unbearable.?
Table 7: Examples for postwar period discussed in Section 4.2. Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(5) Predicted label: group-based solidarity towards Soviet-zone refugees
(Sept 23, 1954)
?[…] den Ländern [wurden] für die wohnungsmäßige Unterbringung der […] Sowjetzonenflüchtlinge […] 350 Millionen DM […] zur Verfügung gestellt […]. Der Bundestag ist […] davon ausgegangen, daß die Sowjetzonenflüchtlinge im Lastenausgleich nicht berücksichtigt werden konnten und daß die Verweisung […] auf den Härtefonds nur als Interimslösung gedacht war. […] Die Bundesregierung prüft […] noch […] Maßnahmen, um den bei den Sowjetzonenflüchtlingen auftretenden Notständen abzuhelfen.? ?[…] the federal government provided 350 million DM for the housing of […] Soviet-zone refugees […]. Parliament assumed that Soviet-zone refugees could not be included in the Equalization of Burdens and that referral […] to the hardship fund was intended only as an interim solution. […] The federal government is still examining measures to address the emergencies arising among the Soviet-zone refugees.?
(6) Predicted label: group-based anti-solidarity towards foreigners
(June 21, 1956)
?[…] Wir müssen das Bundesgebiet nicht nur gegen politisch subversive Elemente, wir müssen uns auch gegen kriminelle Elemente abschirmen. […] Für die Bundesrepublik kommt […] hinzu, daß sie […] Anlaß hat, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden. […] alle internationalen Verpflichtungen, die wir […] erfüllen, [gehören] selbstverständlich nicht unter das, was ich hier ausgeführt habe.? ?[…] We must shield the Federal territory not only against politically subversive elements, we but also against criminal elements. […] For the Federal Republic there is the added circumstance that it has reason to shield itself against tens of thousands of stateless foreigners who […] come with the intention of seeking economic support or political asylum here. […] all international obligations which we […] fulfill, of course, do not fall under what I have set out here.?
(7) Predicted label: group-based solidarity towards foreigners
(Oct 18, 1950)
?[…] Auch in der öffentlichen Fürsorge sollen sie Leistungen in gleicher Höhe erhalten wie deutsche Staatsangehörige. Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen wie Deutschen gewährt werden. […] die Bundesrepublik [gewährt] ihnen einen Rechtsschutz, dessen sie bedürfen, weil sie nicht mehr den Schutz ihres Heimatstaates genießen […]. Der Gesetzentwurf geht […] über den Konventionsentwurf, dessen Grundsätze Mindestforderungen für die Behandlung heimatloser Ausländer darstellen, hinaus. […]? ?[…] In public welfare too, they are to receive benefits equal to those of German citizens. They are to be granted access to schools of all kinds under the same conditions as Germans. […] the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state […]. The draft law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners. […]?
(8) Predicted label: group-based solidarity towards foreigners
(April 19, 1951)
?[…] Wir begrüßen die Verabschiedung des Gesetzes für heimatlose Ausländer, weil uns damit die Möglichkeit geboten war, dem Ausland zu beweisen, wie wir Angehörige anderer Nationen in unserem eigenen Staate zu behandeln gewillt sind, und weil wir uns […] erhoffen, daß sie […] dazu beitragen wird, die Frage der Internationalisierung unseres Flüchtlingsproblems und des Zuflusses von Mitteln aus dem Ausland zu fördern. […]? ?[…] We welcome the adoption of the law for stateless foreigners, because it gave us the opportunity to demonstrate to foreign countries how we intend to treat nationals of other nations within our own state, and because we […] hope that it will make a contribution to promoting the internationalization of our refugee problem and the inflow of funds from abroad. […]?
(9) Predicted label: empathic solidarity towards foreigners
(June 15, 1955)
?[…] Es ist kein Zweifel, daß die heimatlosen Ausländer natürlich vom Arbeitsministerium und vom Innenministerium […] eine rein materielle Unterstützung bekommen. Aber diese materielle Unterstützung gibt keine Möglichkeit, ihnen kulturell oder sonstwie unter die Arme zu greifen. […] [Es ist sehr schwer] diese Dinge heute den Ländern nahezulegen.? ?[…] There is no doubt that the stateless foreigners, of course, receive […] purely material support from the Ministry of Labor and the Ministry of the Interior. But this material support provides no opportunity to assist them culturally or in other ways. […] [It is very difficult] today to persuade the Länder on these matters. […]?
Table 8: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(10) Predicted label: exchange-based solidarity towards expellees,
Party: FDP
(Oct 20, 1949)
?Die werden […] von früh morgens bis spät abends arbeiten, wenn man ihnen das nötige Gelände, [Baumaterialien oder das Geld] zur Verfügung stellt. So werden sie in Gemeinschaftsarbeit wesentlich mehr leisten, als wir das sonst heute bei diesen Bauten sehen. In diesem Falle wäre das Geld, das wir dafür aufnehmen, praktisch nur als Kredit anzusehen. Wir hätten die Möglichkeit, daß diese Heimatvertriebenen sich ihr Heim selbst bauen, daß es allmählich ihr Eigentum wird. Wenn wir nur normale Mietsätze verrechnen und das Kapital ohne Zinsen oder zu geringen Zinssätzen verzinsen lassen, so wäre in wenigen, spätestens 15 Jahren, dieses Haus […] Eigentum des Betreffenden. ? ?They would […] work from early morning until late at night if they were provided with the necessary land, [building materials or the money]. In cooperative work they would achieve far more than we otherwise see with such construction today. In this case the money we raise would practically only be regarded as a loan. We would have the possibility that these expellees build their homes themselves, gradually becoming their property. If we only charged normal rents and let the capital bear no or low interest, then within at most fifteen years this house […] would become the property of the person concerned.?
(11) Predicted label: group-based solidarity towards expellees,
Party: GB/BHE
(Oct 29, 1953)
?[…] Wir dürfen die Augen auch nicht vor der Tatsache verschließen, daß Millionen von Vertriebenen und Kriegsgeschädigten noch nicht den Anschluß an unsere soziale Entwicklung gefunden haben. Auch hier mußten bisher Zehntausende arbeitswilliger tüchtiger Mitbürger abseitsstehen. Bei unserem Verlangen nach beschleunigter Eingliederung dieser Bevölkerungskreise geht es nicht nur um die Ablösung einer unproduktiven Fürsorge durch aufbauende Mitarbeit. Es geht auch nicht nur um die Erfüllung von ?Rechtsansprüchen?, und ich darf diesen Ausdruck Rechtsansprüche hier noch einmal betonen. ? ?[…] We must not close our eyes to the fact that millions of expellees and war victims have still not been integrated into our social development. Here too, tens of thousands of willing and capable fellow citizens have so far had to stand aside. In our demand for the accelerated integration of these groups, it is not only a matter of replacing unproductive welfare with constructive participation. It is also not only a matter of fulfilling legal claims — and I would like to emphasize this expression ?legal claims? once again.?
(12) Predicted label: compassionate solidarity towards refugees,
Party: WAV
(March 23, 1950)
?Die Aufnahme dieser Flüchtlinge wird in manchen Ländern bewußt erschwert, sabotiert und unmöglich gemacht. […] Seitdem die Spannung zwischen Osten und Westen zunimmt, hat ein neuer Strom von Flüchtlingen eingesetzt. […] Polen und die Tschechei [durchführen] die neue Ausweisungsaktion auf Befehl Moskaus, […] um das Flüchtlingsheer in Westdeutschland zu vermehren und noch mehr Arbeitslose zu schaffen, den Staat zu belasten und seine Wirtschaft zu zerstören. […]? ?The admission of these refugees is being deliberately obstructed, sabotaged, and made impossible in some states. […] Since the tension between East and West has increased, a new stream of refugees has begun. […] Poland and Czechoslovakia are carrying out the new expulsions on orders from Moscow, […] in order to enlarge the refugee army in West Germany, to create even more unemployed, to burden the state, and to destroy its economy. […]?
(13) Predicted label: group-based anti-solidarity towards expellees,
Party: BP
(Oct 20, 1949)
?Die Sudetendeutschen […] und die Südostdeutschen [stehen] dem süddeutschen, also dem bayerischen Kulturkreis am nächsten. […] Die aus dem Sudetenland bzw. aus Südosteuropa kommenden Deutschen [sollen] zunächst in Süddeutschland zusammen[gefasst werden], […] damit die Eingliederung […] schneller und reibungsloser vor sich gehen kann. [Wir] wollen […] nicht Schwierigkeiten schaffen, wo Schwierigkeiten nicht überwunden werden müssen, und nicht Gräben aufreißen, wo solche nicht vorhanden sind. Das hat mit Rassentheorie […] gar nichts zu tun. […] ? ?The Sudeten Germans […] and the Southeast Germans are closest to the South German, that is, the Bavarian cultural sphere. […] The Germans coming from the Sudetenland and from Southeastern Europe [should] initially be concentrated in southern Germany, […] so that integration […] can proceed more quickly and smoothly. […] we do not want to create difficulties where difficulties need not arise, and we do not want to open up rifts where none exist. This has nothing to do with racial theory […]. ?
Table 9: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(14) Predicted label: group-based solidarity towards refugees,
Party: KPD
(Dec 13, 1950)
?Wir glauben, daß das Problem der Beschaffung von Arbeit für [Umsiedler] unter dem kapitalistischen System nicht in der Form gelöst werden kann […]. Wir sind also der Auffassung, daß die Umsiedlung eine Notwendigkeit ist […]. Wenn hier der Sprecher der CDU das schöne, stolze und richtige Wort geprägt hat ?Unser Haus ist deutsch?, wenn er am Schluß gesagt hat ?Das ganze Deutschland soll es sein?, wie kann man dann […] gleichzeitig die Tatsache, daß sich die Vertriebenen, Flüchtlinge und Umsiedler, wie Sie sie bezeichnen, politisch zu Parteien zusammengeschlossen haben, als die Gefahr einer Überfremdung bezeichnen? ? ?We believe that the problem of providing work for [resettlers] cannot be solved under the capitalist system […]. We therefore take the view that resettlement is a necessity […]. If the CDU speaker here has coined the fine, proud, and correct phrase ?Our house is German?, and if he concluded by saying ?It shall be the whole of Germany?, how then can one at the same time describe the fact that the expellees, refugees, and resettlers, as you call them, have politically organized themselves into parties as a danger of being overrun by foreigners??
Table 10: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(15) Predicted label: compassionate solidarity towards refugees
Party: Die Grünen
(Nov 6, 2014)
?Ich bin froh, dass es heute bei uns in Deutschland eine andere Willkommenskultur gibt und dass sich viele Menschen mit hohem Engagement für das Wohl der Flüchtlinge, die im Augenblick zu uns kommen, einsetzen. […] Ich will ganz klar sagen: Das Boot ist nicht voll; das Boot ist niemals voll. Heute geht es um Grundrechte von Menschen, denen die Grundrechte in ihren Heimatländern versagt wurden […]. ? ?I am glad that today in Germany we have a different welcoming culture, and that many people are working with great commitment for the well-being of the refugees who are currently coming to us. […] I want to state very clearly: the boat is not full; the boat is never full. Today it is about the fundamental rights of people who have been denied those rights in their home countries […].?
(16) Predicted label: group-based anti-solidarity towards refugees
(Oct 1, 2015)
?[…] Wenn die Zahlen weiterhin auf diesem hohen Niveau bleiben, dann wird Deutschland über kurz oder lang überfordert sein. Was die Registrierung, Antragstellung und Antragsbearbeitung anbelangt, ist dies eine riesige Herausforderung, und die Unterbringung sowie die Integration in unsere Gesellschaft und in den Arbeitsmarkt sind […] auf die Dauer nicht zu leisten […]. allein im September sind mehrere zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen. Sie wurden nicht registriert. Sie wurden nicht kontrolliert. […] Damit besteht auch ein großes Sicherheitsrisiko. ? ?[…] If the numbers remain at this high level, Germany will sooner or later be overburdened. Registration, application processing and decision-making are a huge challenge, and accommodation as well as integration into our society and the labor market will […] not be sustainable in the long run […]. In September alone, several tens of thousands of refugees and asylum seekers came to Germany. They were not registered. They were not controlled. […] This poses a major security risk.?
(17) Predicted label: exchange-based anti-solidarity towards refugees
Party: AfD
(May 16, 2018)
?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug nach derzeitigem Stand um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind. In Deutschland, meine Damen und Herren, lebt bereits eine große soziale Unterschicht […]. Diese Zahl wächst seit 2015 durch die Zuwanderung von Gering- und Unqualifizierten rasant. Meine Damen und Herren, Deutschland ist kein reiches Land, sondern ein leistungsfähiges. ? ?[…] each year about 140,000 highly qualified people leave this country, while in return, according to the current situation, about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate. In Germany, ladies and gentlemen, there already lives a large social underclass […]. Since 2015, this number has been growing rapidly through the immigration of low- and unqualified people. Ladies and gentlemen, Germany is not a rich country, but a capable one.?
Table 11: Examples for the 2009-2025 period discussed in Section 4.3. Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(18) Predicted label: group-based anti-solidarity towards refugees
Party: AfD
(Oct 19, 2022)
?[…] Echte Kriegsflüchtlinge aus der Ukraine brauchen Schutz – keine Frage. Wir sagen aber Nein zu Fluchttourismus durch halb Europa, und wir sagen Nein zu bevorzugter Behandlung auf Kosten der deutschen Steuerzahler. Wir fordern daher die sofortige Abschaltung dieses Sozialstaatsmagneten, […] Wir fordern außerdem die Umstellung von Geldleistungen auf Sachleistungen für Asylbewerber, damit Migrationsanreize gesenkt werden und der Missbrauch effektiv unterbunden wird. […] Und wer sich wie Frau Baerbock in erster Linie der ukrainischen Bevölkerung verpflichtet fühlt, der soll sich in der Ukraine um ein Amt bewerben, aber nicht in Deutschland. ? ?[…] Genuine war refugees from Ukraine need protection — no question. But we say no to flight tourism across half of Europe, and we say no to preferential treatment at the expense of German taxpayers. We therefore demand the immediate shutdown of this welfare magnet, […] We also demand the switch from cash benefits to benefits in kind for asylum seekers, so that incentives for migration are reduced and abuse is effectively curbed. […] And whoever, like Ms. Baerbock, feels primarily obligated to the Ukrainian population should run for office in Ukraine, but not in Germany.?
(19) Predicted label: group-based anti-solidarity towards refugees
Party: AfD
(Dec 5, 2023)
?[…] alte Leute sollen lieber aus ihren großen Wohnungen ausziehen, weil grün-linke Politik Millionen Gäste aus Afrika und dem Nahen Osten einlädt und gleichzeitig den Wohnungsbau abwürgt. […] Dafür werden bereits Zimmer in angeblich christlichen Altenheimen gekündigt, weil angebliche Flüchtlinge lukrativer sind, Messerzuwanderung inklusive. […] Bei diesem Recht auf Wohnungstausch der Linken handelt es sich um eine Mischung aus Neid zulasten alter Leute und Zwang zulasten von Vermietern. Es ist die Fortsetzung einer inländerfeindlichen… ? ?[…] elderly people are supposed to move out of their large apartments because green-left politics invites millions of guests from Africa and the Middle East while simultaneously stifling housing construction. […] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration [^12] Messereinwanderung: a propagandistic term coined by the AfD implying migrants bring knife crime. included. […] This ?right to apartment swaps? proposed by the Left is a mixture of envy at the expense of elderly people and coercion at the expense of landlords. It is the continuation of an anti-native policy…?
(20) Predicted label: exchange-based solidarity towards immigrants
Party: FDP
(May 25, 2023)
?[…] Menschen, die bei uns arbeiten wollen, sollten wir eine echte Chance geben; denn viele, die an unserer Grenze ankommen, wollen ja gerade arbeiten. Wir arbeiten in der Ampel an der Reform des Fachkräfteeinwanderungsgesetzes. Da geht es um mehr Regulierung, gezielte Einwanderung in unseren Arbeitsmarkt. Menschen aus dem Ausland sollten eine realistische Chance bekommen, zu uns zu kommen und arbeiten zu können. Das Fachkräfteeinwanderungsgesetz wird das jetzt endlich besser machen und mehr Menschen die Chance geben, die unsere Wirtschaft auch dringend braucht. […] ? ?[…] People who want to work here should be given a real chance; because many who arrive at our border actually want to work. Within the coalition we are working on reforming the Skilled Immigration Act. This is about more regulation and targeted immigration into our labor market. People from abroad should have a realistic chance to come to us and be able to work. The Skilled Immigration Act will now finally improve this and give more people the opportunity that our economy urgently needs. […]?
(21) Predicted label: group-based anti-solidarity towards migrants
Party: CDU/CSU
(May 14, 2025)
?[…] Die gute Nachricht ist: Der Politikwechsel hat begonnen. […] Die illegale Migration der letzten zehn Jahre gefährdet die politische Stabilität Deutschlands und Europas. […] Unsere Städte und Gemeinden, die Landkreise: Alle sind über dem Limit. In den Kitas, in den Schulen, auf dem Wohnungsmarkt, bei der Sicherheit an den Bahnhöfen und auf den Marktplätzen wird diese Überforderung sehr konkret. Deswegen ist es gut, dass wir gemeinsam im Koalitionsvertrag die Migrationswende vereinbart haben: mehr Rückführungen, schnellere Verfahren und, ja, ebenso die Zurückweisung an den deutschen Grenzen auch bei Asylgesuchen. ? ?[…] The good news is: the policy shift has begun. […] The illegal migration of the last ten years endangers the political stability of Germany and Europe. […] Our towns and municipalities, the districts—everyone is over the limit. In day-care centers, in schools, in the housing market, in security at train stations and in marketplaces, this overburdening becomes very concrete. That is why it is good that we have agreed together in the coalition agreement on a migration turnaround: more deportations, faster procedures, and, yes, also rejections at the German borders, even in the case of asylum applications.?
Table 12: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(22) Predicted label: group-based anti-solidarity towards immigrants
Party: FDP
(Jan 30, 2025)
?Das Grundrecht auf Asyl darf nicht dazu führen, dass allein ein Antrag auf Asyl zu mehreren Jahren freiem und staatlich subventioniertem Aufenthalt in Deutschland führt, wenn er völlig ohne Grundlage erfolgt. Solche Missstände müssen wir beseitigen, damit wir uns auf jene konzentrieren können, die unsere Hilfe wirklich brauchen […]. SPD und Grüne […] wollen die illegale Einwanderung nach Deutschland kaum begrenzen und fordern lieber weiteren Familiennachzug. Diesen falschen Weg ist Deutschland leider viel zu lange gegangen. […] Eine immer linkere Politik wird […] dazu führen, dass die AfD ihr teuflisches Werk immer weiter fortsetzen kann. […] ? ?The fundamental right to asylum must not result in a situation where merely submitting an asylum application leads to several years of free and state-subsidized residence in Germany when it is entirely unfounded. Such abuses must be eliminated so that we can focus on those who truly need our help […]. The SPD and the Greens […] hardly want to limit illegal immigration to Germany and instead prefer to demand further family reunification. Unfortunately, Germany has followed this wrong path for far too long. […] Ever more left-leaning policies […] will only enable the AfD to continue its diabolical work […].?
(23) Predicted label: group-based anti-solidarity towards foreigners
Misclassified; true: group-based solidarity
Party: SPD
(Apr 10, 2024)
?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht. Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird. Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […]. Sie ?übersehen? in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben. Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen. Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist. […] ? ?Let us turn to the AfD’s draft and to what makes it simply atrocious. You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved. Not only does this place all foreigners under general suspicion […], but it also violates EU law […]. You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency. But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is. […]?
(24) Predicted label: group-based anti-solidarity towards migrants
Party: SPD
(Dec 9, 2024)
?Das Sicherheitspaket, das heute hier zur Debatte steht, ist der erste Schritt, um Konsequenzen aus der furchtbaren Katastrophe in Solingen zu ziehen. Waffenverbotszonen […] helfen [aber] Waffenverbotszonen nicht. Deshalb brauchen wir mehr Befugnisse für das Bundeskriminalamt und das Bundesamt für Migration. Sie müssen in die Lage versetzt werden, mit den öffentlich zugänglichen Informationen Straftaten zu verhindern. Das ist dringend erforderlich, und deshalb werden wir das jetzt umsetzen. Wir werden weitere Schritte veranlassen, damit ausreisepflichtige Personen unser Land verlassen. ? ?The security package being debated here today is the first step toward drawing consequences from the terrible catastrophe in Solingen. Weapons prohibition zones […] do not, however, help against terrorists. That is why we need more powers for the Federal Criminal Police Office and the Federal Office for Migration. They must be put in a position to use publicly available information to prevent crimes. That is urgently necessary, and that is why we will now implement it. We will take further steps to ensure that persons obliged to leave the country actually leave our country.?
(25) Predicted: group-based anti-solidarity towards migrants
Misclassified; true: empathic solidarity
Party: Die Linke
(June 5, 2025)
?[…] Migration wird pauschal als sicherheitspolitisches Problem dargestellt. Die AfD führt die sogenannte gescheiterte Migrationspolitik in einer Liste von Krisen auf – neben dem Krieg in der Ukraine, der Coronapandemie und der hybriden Bedrohung. Damit betreibt sie eine gefährliche und spaltende Gleichsetzung. Sie erklärt das bloße Vorhandensein von Migration zur Bedrohung der inneren und äußeren Sicherheit und nennt sie in einem Atemzug mit Terrorismus, Krieg und Katastrophen. ? ?Migration is presented across the board as a security policy problem. The AfD lists the so-called failed migration policy alongside the war in Ukraine, the coronavirus pandemic, and hybrid threats. In doing so, it engages in a dangerous and divisive equation. It declares the mere existence of migration to be a threat to internal and external security, mentioning it in the same breath as terrorism, war, and disasters.?
Table 13: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
Original Text Translation
(26) Predicted label: exchange-based solidarity towards immigrants
Party: Die Grünen
(Jan 19, 2024)
?Wissenschaftliche Institute sprechen von 400 000 Menschen jedes Jahr. Darauf müssen wir die Gesellschaft und die Politik vorbereiten. Wir müssen also ohnehin die entsprechende Infrastruktur ausbauen – Wohnungen, Schulen, Kinderbetreuung. Wir brauchen gezielte Zuwanderung, und dafür haben wir das Fachkräfteeinwanderungsgesetz gemacht. Aber wir wären doch mit dem Klammerbeutel gepudert, wenn wir nicht auch das Potenzial der Menschen, die aus humanitären Gründen zu uns kommen, nutzten und dafür sorgten, dass sie aus- und weitergebildet werden und Teil unserer Gesellschaft werden. Das wäre die Hauptaufgabe. Und weil wir Zuwanderung brauchen, brauchen wir eine Willkommenskultur. ? ?Scientific institutes speak of 400,000 people every year. Society and politics must be prepared for that. We therefore have to expand the relevant infrastructure anyway — housing, schools, childcare. We need targeted immigration, and for that we created the Skilled Workers Immigration Act. But we would have to be out of our minds not to also make use of the potential of people who come to us for humanitarian reasons, and ensure that they receive education and training and become part of our society. That would be the main task. And because we need immigration, we need a culture of welcome.?
(27) Predicted label: empathic solidarity towards migrants
Party: Die Linke
(Jan 30, 2025)
?Sie sagen, Sie wollen Frauen und Kinder schützen. Sie wollen Kinder und Frauen in Kriegsgebiete abschieben. So ist es. Wir führen eine Debatte über Sicherheit und Migration. Und leider erklären alle Parteien außer der Linken, dass sie mit Abschottung dieses Land sicherer machen wollen, und reden von ?Abschiebungen im großen Stil?, als seien Migranten grundsätzlich eine Gefahr. So werden Menschen unter Generalverdacht gestellt. Weniger Migration mache das Land sicherer, das ist doch Unsinn. ? ?You say you want to protect women and children. You want to deport children and women to war zones. That is the reality. We are having a debate about security and migration. And unfortunately, all parties except the Left declare that they want to make this country safer through isolation, and talk about ?deportations on a large scale,? as if migrants were inherently a danger. In this way, people are placed under general suspicion. Less migration makes the country safer — that is nonsense.?
Table 14: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.

### v2

@@
-LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade
+LLM Analysis of 150+ Years of German Parliamentary Debates on Migration Reveals a Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade
 Abstract
-Migration has been a core topic in German political debate, from millions of expellees post World War II over labor migration to refugee movements in the recent past.
-Studying political speech regarding such wide-ranging phenomena in depth traditionally required extensive manual annotations, limiting the scope of analysis to small subsets of the data.
-Large language models (LLMs) have the potential to partially automate even complex annotation tasks.
-We provide an extensive evaluation of a multiple LLMs in annotating (anti-)solidarity subtypes in German parliamentary debates compared to a large set of thousands of human reference annotations (gathered over a year).
-We evaluate the influence of model size, prompting differences, fine-tuning, historical versus contemporary data; and we investigate systematic errors.
-Beyond methodological evaluation, we also interpret the resulting annotations from a social science lense, gaining deeper insight into (anti-)solidarity trends towards migrants in the German post-World War II period and recent past.
-Our data reveals a high degree of migrant-directed solidarity in the postwar period, as well as a strong trend towards anti-solidarity in the German parliament since 2015, motivating further research.
-These findings highlight the promise of LLMs for political text analysis and the importance of migration debates in Germany, where demographic decline and labor shortages coexist with rising polarization.
-keywords:
-solidarity, large language models, annotation, political discourse
-$\cup@datesfalse$
-Aida Kostikova]aida.kostikova@uni-bielefeld.de
+Migration has been a core topic in German political debate, from the postwar displacement of millions of expellees to labor migration and recent refugee movements.
+Studying political speech across such wide-ranging phenomena in depth has traditionally required extensive manual annotation, limiting analysis to small subsets of the data.
+Large language models (LLMs) offer a potential way to overcome this constraint.
+Using a theory-driven annotation scheme, we examine how well LLMs annotate subtypes of solidarity and anti-solidarity in German parliamentary debates and whether the resulting labels support valid downstream inference.
+We first provide a comprehensive evaluation of multiple LLMs, analyzing the effects of model size, prompting strategies, fine-tuning, historical versus contemporary data, and systematic error patterns.
+We find that the strongest models, especially GPT-5 and gpt-oss-120B, achieve human-level agreement on this task, although their errors remain systematic and bias downstream results.
+To address this issue, we combine soft-label model outputs with Design-based Supervised Learning (DSL) to reduce bias in long-term trend estimates.
+Beyond the methodological evaluation, we interpret the resulting annotations from a social-scientific perspective to trace trends in solidarity and anti-solidarity toward migrants in postwar and contemporary Germany.
+Our approach shows relatively high levels of solidarity in the postwar period, especially in group-based and compassionate forms, and a marked rise in anti-solidarity since 2015, framed through exclusion, undeservingness, and resource burden.
+We argue that LLMs can support large-scale social-scientific text analysis, but only when their outputs are rigorously validated and statistically corrected.
 1 Introduction
-Migration is a defining feature of the 21st century, explicitly recognized in the 2030 Agenda for Sustainable Development wilkinson2022climate (105, 1).
-As the worldbank2023 (106) stresses, global economic imbalances, demographic divergence, and climate change make migration increasingly unavoidable, especially from the Global South to the Global North almulhim2024climate (3, 55, 69).
-However, if governed responsibly, migration can also drive prosperity and contribute to achieving the Sustainable Development Goals worldbank2023 (106, 38, 1).
-Within Europe, Germany occupies a particularly important position as a major migration destination eurostat2024migration (32).
-As the EU’s largest economy, it offers labor market opportunities, institutional support, and stability bertoli2016crisis (10).
-At the same time, demographic decline and skilled-labor shortages make sustained immigration economically beneficial smith2025demographic (27, 29, 4).
-Migration has also profoundly shaped Germany’s political development, from the resettlement of millions of expellees after WWII, to the labor migration of the 1950s–1970s, to refugee movements in the past decade kaya2017inclusion (52, 36).
-This raises important questions: how does German politics approach migration, and to what extent does it express solidarity – or anti-solidarity – toward migrants?
-Answering such questions requires systematic analysis of political discourse across long time spans, traditionally relying on manual textual analysis and qualitative interpretation, which remain powerful tools randour2020twenty (78).
-However, such methods are labor-intensive bhattacherjee2012social (11) and thus often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized babalola2021current (7).
-Computational social science (CSS) emerged to address the scale and complexity of modern data, yet it often struggles with theoretical alignment and interpretive depth lazer2024future (62).
-Large language models (LLMs) promise to bridge this gap by enabling automated, in-depth analysis of large-scale text.
-While LLMs have shown strong performance on tasks like ideological stance detection ziems2024can (109), it remains an open question whether they can handle more complex, theory-driven sociopolitical annotation tasks.
-We provide a partial answer: If accompanied by substantial validation and careful interpretation, LLMs can approximate the performance of trained human annotators and generate annotations that help quantify large-scale trends of complex constructs.
-In particular, we apply LLMs to annotate solidarity and anti-solidarity towards women and migrants in German parliamentary speeches, compare LLM annotations to thousands of human reference annotations, and interpret the (anti-)solidarity trends between 1949-1969 and 2009-2025 in light of German political history.
-Figure 1: Overview of the annotation scheme adapted from thijssen2012mechanical (97).
+Migration is a defining feature of the 21st century(Wilkinson et al., 2022).
+Economic imbalances, demographic divergence, and climate change make migration are increasing pressures for migration, especially from the Global South to the Global North (World Bank, 2023; Almulhim et al., 2024; Marois et al., 2020).
+If governed responsibly, it can contribute to prosperity and support sustainable development (Gavonel et al., 2021; Adger et al., 2024).
+Within Europe, Germany is a major migration destination (European Commission, Eurostat, 2024), and migration has profoundly shaped its development, from post-Second World War expellee resettlement to labor migration in the 1950s–1970s and refugee movements in the 2010s-2020s (Kaya, 2017; Fröhlich, 2025).
+As the EU’s largest economy, Germany remains an attractive destination, while demographic decline and skilled-labor shortages make continued immigration economically beneficial (Bertoli et al., 2016; Donal Smith and Huang, 2025; Eissel et al., 2021; Angenendt et al., 2023).
+At the same time, anti-migration stances have become more visible in political debate, often in relation to perceived risks such as pressure on public services, labor-market and resource competition, and the rise of right-wing populism (Frey, 2020; Gessler and Hunger, 2023; Messing and Ságvári, 2021; Esguerra et al., 2023).
+This raises a central question: how does German politics discuss migration, and to what extent does it express solidarity or anti-solidarity toward migrants?
+Answering such questions requires systematic analysis of political discourse across long time spans to assess continuity and change in migration discourse.
+Such analysis has traditionally relied on manual textual analysis and qualitative interpretation (Randour et al., 2020).
+However, such methods are labor-intensive (Bhattacherjee, 2012) and, hence, often restricted to selective subsets of data, which risks sampling bias and leaves large corpora underutilized (Babalola and Nwanzu, 2021).
+Computational social science (CSS) emerged partly in response to the scale and complexity of modern data, which exceed the analytic reach of traditional approaches (Lazer et al., 2009).
+More recently, large language models (LLMs) have been introduced into CSS as a means of scaling annotation and automating parts of the research workflow (Pavlovic and Poesio, 2024).
+At the same time, their use raises several methodological challenges.
+(i) Although LLMs show strong performance on tasks such as ideological stance detection (Ziems et al., 2024), it remains unclear whether they can perform more complex, theory-driven sociopolitical annotation tasks.
+(ii) Even when LLMs appear to perform well, recent work questions whether their outputs can support valid downstream inference or serve as ground-truth labels, given their sensitivity to model choice, prompt design, and configuration, as well as their potential to introduce systematic bias (Baumann et al., 2025; Gligoric et al., 2025; Krsteski et al., 2025; Egami et al., 2023).
+(iii) These concerns connect to ongoing debate in CSS emphasizing that computational methods and digital data must be tightly integrated with social scientific theory, concepts, and theoretically meaningful research questions (Lazer and Ognyanova, 2024; Jungherr and Theocharis, 2017; Theocharis and Jungherr, 2021; Soemer et al., 2025).
+In a prior conference paper, we addressed challenge (i) by showing that LLMs can approximate human performance on a complex annotation task focused on (anti-)solidarity in parliamentary text and grounded in a theoretically informed, fine-grained taxonomy (Kostikova et al., 2024).
+In this paper, we extend that analysis in three ways: the first two are methodological, and the third is substantive.
+First, to address (i) more rigorously, we evaluate a broader range of LLMs, compare their predictions against a larger set of human reference annotations, and systematically assess the effects of prompts, model size, and historical versus contemporary data.
+Second, to address (ii), we examine whether strong annotation performance is sufficient for valid downstream inference by analyzing systematic model errors, testing ensemble predictions, and applying statistical bias correction (Egami et al., 2023).
+Third, to address (iii), we apply this validated pipeline to the analysis of solidarity and anti-solidarity toward migrants in German parliamentary speeches and interpret the resulting trends between 1949 and 1957 and between 2009 and 2025 in light of German political history.
+Figure 1: Overview of the annotation scheme adapted from Thijssen (2012).
 Statements are first classified at a high level as expressing solidarity, anti-solidarity, mixed, and none.
@@
 The example sentence in the figure illustrates exchange-based anti-solidarity.
-We focus on solidarity as a particularly demanding test case for such an approach.
-While frequently invoked in political discourse, solidarity has often been treated as a symbolic notion rather than an operational concept kozlowska2024solidarity (58), which has limited large-scale empirical research.
-To address this gap, we approach solidarity as a socially meaningful category that can be systematically studied in political text.
-For this purpose, we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97), which provides a structured and annotatable framework.
-It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure 1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification).
-Germany offers a particularly suitable context for such a study: solidarity is deeply rooted in its political vocabulary fiegle2003solidarite (34, 94, 84), and its multi-party system enables us to trace how meanings of (anti-)solidarity vary across parties and over time.
-In a prior conference paper, we already performed a first analysis of LLM annotations of (anti-)solidarity towards women and migrants on this data set kostikova-fine (57).
-In this paper, we 1) evaluate a much wider range of LLMs, 2) systematically evaluate the effect of prompts, model size, and historic (versus contemporary) data, 3) include a much larger data set of human annotations, 4) validate robustness by comparing the time series of model-predicted (anti-)solidarity shares from the selected model (Llama-3.3-70B) and an independent model (GPT-4), and 5) most importantly, we interpret the (anti-)solidarity trends towards migrants over time to derive novel insights for political science.
-In particular, we find a) that different groups of migrants are associated with different types of (anti-)solidarity in German parliamentary debates during the post-World War II period; and b) that anti-solidarity towards migrants has dramatically increased in the German parliament over the past 10 years, indicating a shift to the right (especially in the liberal and conservative parties) on this issue that motivates further research.
+We focus on solidarity because it is a particularly demanding test case for operationalization and annotation, both for human annotators and for LLMs.
+Solidarity is an essentially contested concept: it is difficult to define, value-laden, ambiguous, multi-faceted, context-dependent, and open to change over time (Kozłowska, 2024).
+This makes annotation demanding, as coders must map concrete and historically variable political statements onto abstract categories of (anti-)solidarity.
+To address this challenge, we approach solidarity as a theoretically informed category and apply solidarity typology of Thijssen (2012), which offers a structured, annotatable framework.
+It distinguishes four types of (anti-)solidarity – group-based, compassionate, exchange-based, and empathic (see the annotation scheme in Figure˜1) – based on the interplay of structural conditions (e.g., group membership) and intersubjective recognition (e.g., emotional identification).
+Germany is a particularly suitable context for such a study since solidarity is deeply rooted in its political vocabulary (Fiegle, 2003; Stjernø_2005; Reisz, 2006), and its multi-party system enables to trace how meanings of (anti-)solidarity vary across parties and over time.
+Overall, we find that in our task, the strongest models approach human agreement, and in some settings even surpass it, with GPT-5 performing best overall and gpt-oss-120B being the strongest open model.
+However, we conclude that raw model predictions are not sufficient for reliable downstream inference: although strong models achieve near-human agreement at the instance level, their errors remain structured and correlated, limiting the benefits of simple ensemble methods.
+At the same time, human agreement itself remains limited, especially at the fine-grained level, indicating substantial label variation.
+In response, combining soft-label model outputs with statistical correction, such as Design-based Supervised Learning (DSL) (Egami et al., 2023), achieves more accurate recovery of human-derived temporal trends.
+Substantively, this validated pipeline reveals high levels of migrant-directed solidarity in the postwar period, especially forms centered on shared belonging and support for vulnerable groups (group-based and compassionate), thereby corroborating existing qualitative research on early postwar inclusion politics.
+For the contemporary period, it reveals a marked shift since 2015 toward anti-solidarity in German parliamentary discourse, framed through exclusion, undeservingness, and burdens on collective resources (group-based, compassionate and exchange-based).
+To our knowledge, this is the first large-scale quantitative evidence of that shift.
 1.1 Related Work
-Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; and (ii) the use of LLMs as annotation tools in political and social science contexts.
-While prior studies have engaged with each of these dimensions individually, few have combined them to enable large-scale, longitudinal analysis of solidarity in political speech.
-We briefly review these literatures below.
+Our work builds on and connects the following directions of existing research: (i) sociological and political theories of solidarity and their operationalization in NLP; (ii) the use of LLMs as annotation tools in political and social science contexts; and (iii) valid downstream inference from imperfect machine-generated labels.
 Solidarity in Political Discourse and NLP
-Sociological theories of solidarity often begin with Durkheim’s classical distinction between mechanical and organic forms, which describe different bases for social cohesion.
-While influential, this binary model has been criticized for its structuralism and linear view of social change.
-thijssen2012mechanical (97) expands on Durkheim by integrating recognition-based theories such as those developed by honneth1996struggle (45, 46) and proposing a fourfold typology with group-based, compassionate, exchange-based, and empathic subtypes, which can be applied to both historical and contemporary political discourse.
-thijssen2022s (98) apply this typology to Belgian party manifestos, showing how different parties emphasize different solidarity frames in line with broader ideological divides.
+The theoretical foundation for our analysis is the solidarity framework of Thijssen (2012), which extends Durkheim’s distinction between mechanical and organic solidarity by incorporating recognition-based theories of Honneth (1996, 2014).
+Thijssen distinguishes four subtypes of solidarity—group-based, compassionate, exchange-based, and empathic—applicable to both historical and contemporary discourse.
+Thijssen and Verheyen (2022) apply this typology to Belgian party manifestos, showing how parties emphasize different frames in line with ideological divides.
+Survey and interview studies also examine solidarity longitudinally (Schönweitz et al., 2024; Sengupta et al., 2023), but these approaches are not suited for large-scale text analysis or fine-grained classification.
 We build on this by applying the typology to parliamentary debates, extending the temporal scope and scaling the analysis with LLMs.
-While some studies have examined solidarity longitudinally using interviews or survey data schonweitz2024solidarity (89, 92), these approaches are not designed for large-scale text analysis or fine-grained typological classification.
-At the same time, solidarity has received relatively little attention in computational text analysis, where more application-driven concerns such as hate speech and misinformation detection have dominated the field.
-Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements weber2024social (103, 28), the COVID-19 pandemic kalaloi2021mediated (51), and natural disasters santhanam2019stand (88).
-These works typically use supervised models to detect solidarity or anti-solidarity in large-scale Twitter data and track how such expressions evolve over short timeframes.
-While they demonstrate the feasibility of automated solidarity detection, they focus on specific, time-bound crises and treat solidarity as a binary sentiment-like category rather than a differentiated moral frame.
-In contrast, our work applies a theory-driven typology to political speech, enabling fine-grained analysis of solidarity in parliamentary discourse.
-LLMs as Annotation Tools in Political and Social Science.
+Solidarity has received relatively little attention in computational text analysis.
+Recent NLP studies have examined expressions of solidarity in social media during crisis events, such as refugee movements (Weber et al., 2024b; Eger et al., 2022), the COVID-19 pandemic (Kalaloi et al., 2021), and natural disasters (Santhanam et al., 2019).
+These studies focus on specific, time-bound crises and typically treat (anti-)solidarity as a binary classification task, whereas our work applies a differentiated, theory-based typology to parliamentary discourse.
+LLMs as Annotation Tools in Political and Social Science
 LLMs are increasingly used in political and social science research.
-Most applications focus on standard classification tasks such as political ideology detection (heseltine2024large, 43, e.g.); sentiment analysis lashitew2025corporate (61, 37); stance detection (cao2024can, 19, e.g.); and misinformation detection (huang2025unmasking, 47, e.g.).
-For broader surveys of LLM applications, benefits, and challenges in social and political science, see thapa2025large (95, 66, 109).
-However, conclusions about LLM performance in these fields vary.
-Some studies report that LLMs show near-human performance rytting2023towards (87) or even outperform humans, as well as traditional models, on short texts and tasks like ideology classification le2025positioning (63, 43).
-In contrast, several studies show that smaller fine-tuned models like BERT still outperform LLMs, especially when labeled data is available wang2024selecting (102, 18) or when tasks require expert-defined taxonomies and contextual reasoning ziems2024can (109).
+Prior work has considered annotation of political ideology (e.g.
+Heseltine and Clemm von Hohenberg, 2024), sentiment (Lashitew and Mu, 2025; Fu et al., 2024), stance (e.g.
+Cao et al., 2024), and misinformation (e.g.
+Huang et al., 2025).
+For broader surveys of LLM applications in social and political science, see Thapa et al.
+(2025); Li et al.
+(2024); Ziems et al.
+(2024).
+Conclusions about LLM performance vary: some find near-human (Rytting et al., 2023) or superior results on short texts (Le Mens and Gallego, 2025; Heseltine and Clemm von Hohenberg, 2024), while others show smaller fine-tuned models perform better with labeled data or expert taxonomies (Wang et al., 2024; Burnham et al., 2024; Ziems et al., 2024).
 Our work provides a systemic evaluation of LLMs for such a complex, theory-informed task, namely (anti-)solidarity detection.
 Beyond standard classification tasks, a subset of research has applied language models to more complex, theory-driven categories.
-For example, erhard2025popbert (30) develop PopBERT, a fine-tuned transformer trained to detect populist rhetoric and ideological dimensions in German Bundestag speeches using a custom annotation scheme.
-card2022computational (20) analyze 140 years of U.S.
-political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, identify rhetorical frames, and detect dehumanizing metaphors, demonstrating the utility of fine-tuned LMs for large-scale, theory-aligned analysis.
-Related work has applied fine-tuned models to pragmatic and rhetorical categories in German parliamentary debates, such as speaker attribution, speech acts, and opinion role labelling reinig2024politics (83, 81, 80).
-In contrast to fine-tuned approaches, ranjit2024oath (79) use general-purpose LLMs (e.g., GPT-4, Flan-T5) to apply a custom typology of attitudes toward homelessness, developed from framing theory and grounded theory.
-Their evaluation focuses on aggregate trends and correlations with external structural indicators, using LLMs to annotate a large corpus of tweets at scale and over time.
-Our work differs in several important ways.
-Unlike erhard2025popbert (30) and card2022computational (20), we use state-of-the-art LLMs and avoid the need for supervised data sets and fine-tuning.
-Unlike ranjit2024oath (79), who also use prompted LLMs, we adopt a typology grounded in established sociological theory rather than one developed inductively for a specific domain.
-We apply this typology to more formal political speech rather than short, informal social media texts, and evaluate model performance through systematic comparison with human annotations.
-Finally, our analysis spans over 150 years of parliamentary debate, encompassing two case studies, historical and contemporary, and enabling fine-grained, longitudinal insight into political discourse.
+Erhard et al.
+(2025) develop PopBERT, a fine-tuned transformer for detecting populist rhetoric in German Bundestag speeches, while Card et al.
+(2022) analyze 140 years of U.S.
+political speech with fine-tuned RoBERTa and BERT models to classify immigration tone, frames, and dehumanizing metaphors.
+Other work fine-tune models for pragmatic and rhetorical categories in German debates, such as speaker attribution, speech acts, and opinion roles (Reinig et al., 2024; Rehbein et al., 2024; Rehbein and Ponzetto, 2024).
+In contrast to fine-tuned approaches, Ranjit et al.
+(2024) use general-purpose LLMs (e.g., GPT-4, Flan-T5) with a custom typology of attitudes toward homelessness to annotate large-scale Twitter data and study aggregate trends.
+Valid Inference from Imperfect LLM Labels
+Recent work moves beyond raw annotation accuracy to ask whether machine-generated labels actually support valid downstream inference.
+As Egami et al.
+(2023) stress, downstream uses of surrogate labels often ignore measurement error, that is, the unknown, heterogeneous mismatch between gold-standard and surrogate labels, which can distort estimates and invalidate inference even when surrogate labels appear highly accurate (Egami et al., 2023; Knox et al., 2022).
+In CSS, such error is especially difficult to avoid because annotation tasks are often intrinsically challenging (Ziems et al., 2024), and prior work shows that LLM annotations often diverge from expert coding decisions, especially on more complex tasks (Lin and Zhang, 2025; de Pieuchon et al., 2024).
+In response, bias-corrected estimators combine machine-generated labels with smaller sets of gold-standard annotations to support valid downstream inference.
+These include Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023, 2024), Design-Based Supervised Learning (DSL) (Egami et al., 2023), and Confidence-Driven Inference (CDI) (Gligorić et al., 2025).
+We build most directly on DSL because it was developed with CSS applications in mind and fits our design, which combines a fixed decade-stratified human-coded sample with multiple historical trend analyses.
+Recent benchmarking also suggests that DSL often yields lower bias and more precise estimates than PPI in text classification tasks (de Pieuchon et al., 2025).
+DSL has been evaluated in a limited number of benchmark and applied settings.
+Egami et al.
+(2023) assess DSL in a custom regression task and in class prevalence estimation across CSS text classification datasets from Ziems et al.
+(2024).
+Rister Portinari Maranca et al.
+(2025) apply DSL to AI-assisted image labeling, and Krsteski et al.
+(2025) to LLM-generated survey responses.
+In settings closer to ours, Stuhler et al.
+(2026) study U.S.
+media representations of the timing and urgency of climate change using a hybrid pipeline of rule-based extraction and RoBERTa-based classifiers, and apply DSL to debias trend and regression estimates.
+Liu (2025) examines Chinese-language social media discourse around two major 2024 U.S.
+political events using LLM-based coding of misinformation, partisan stance, affective polarization, and attack language.
+There, LLM annotation is the main measurement framework, while DSL is presented as a robustness check and is mainly implemented in the user-engagement analysis rather than across all analyses.
+Across work using both raw and bias-corrected model predictions, our study differs in several respects.
+Unlike Erhard et al.
+(2025); Card et al.
+(2022); Stuhler et al.
+(2026), we use state-of-the-art prompted LLMs and avoid the need for task-specific fine-tuning, and apply DSL directly to their outputs.
+Unlike Ranjit et al.
+(2024) and Liu (2025), which also use prompted LLMs, we apply a theory-driven typology to more than 150+ years of parliamentary speech rather than social media text, and we make LLM validation and DSL-based bias correction central to inference across all analyses rather than treating them as robustness checks, thus enabling fine-grained, longitudinal insight into political discourse.
 2 Methodology
 2.1 Data
-We build on the DeuParl dataset introduced in DeuParl (101), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022.
-We extend this corpus to include protocols up to June 2025.
-We follow the preprocessing steps from DeuParl (101) and include metadata such as date, session number, and legislative period.
-We reuse the keyword sets (see a full list of keywords in A) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in kostikova-fine (57), consisting of 32 and 18 terms respectively.
-Based on these keywords, we extract 63k migrant-related and 138k women-related sentences (instances) from DeuParl, additionally including three sentences before and after each instance for context.
-We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) provided by poradaautomatische (77).
-See B for details.
-Fig.
-2 shows the number of instances over time, as well as their share relative to all sentences in the DeuParl dataset.
-Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015, coinciding with refugee crisis debates.
-Notably, migrant mentions also surpass women-related discourse in the immediate post-war years.
+We build on the DeuParl dataset introduced in Walter et al.
+(2021), which consists of plenary protocols from the German Reichstag (en.: imperial diet) (sourced from the Reichstagsprotokolle) and the Bundestag (en.: federal parliament) (via the Open Data platform), spanning 1867–2022111Protocols from the Volkskammer (the GDR parliament) are not included due to lack of availability.
+As a result, GDR-specific narratives and solidarity frames are absent from the corpus..
+We extend this corpus to June 2025, following the preprocessing steps from Walter et al.
+(2021) and including metadata such as date, session number, and legislative period.
+We reuse the keyword sets (see Appendix˜A in the Appendix) for the categories Migrant (en.: migrant) and Frau (en.: woman) introduced in Kostikova et al.
+(2024), comprising 32 and 18 terms, respectively.
+Based on these, we extract 63k migrant-related and 138k women-related sentences (hereafter, instances) from DeuParl, including the three preceding and three following sentences as context.
+We enrich each instance with speaker and party information, using Open Data metadata for recent sessions (20th–21st periods) and improved XML markup for earlier ones (1st–19th) (Porada, 2019); see Appendix˜B in the Appendix for details.
+We include women alongside migrants in the evaluation to test model generalizability across distinct marginalized groups, though our substantive analysis focuses on migrants.
+Figure˜2 shows the number of instances over time and their share relative to all DeuParl sentences.
+Mentions of both migrants and women increase substantially after 1990, with migrant-related speech peaking and briefly overtaking that of women around 2015 during refugee crisis debates.
+Notably, mentions of migrants also surpass women-related discourse in the immediate post-war years.
 Figure 2: Absolute number of instances (left) and their share relative to all DeuParl sentences (right) for the Woman and Migrant categories.
 2.2 Data Annotation
-To validate our models and enable further qualitative analysis, we build on the human-annotated dataset introduced in kostikova-fine (57), which comprises 2,864 instances (1,437 for migrants, 1,427 for women) from the DeuParl corpus.
-[^1] Annotation was carried out by five trained student assistants through an iterative process, achieving Cohen’s $\kappa$ of 0.62 (high level) and 0.42 (fine-grained).
-Final labels in the dataset are assigned by majority vote.
-Each instance is labeled according to a fine-grained solidarity framing scheme that distinguishes four high-level stance categories – solidarity (willingness to share resources, following Lahusen.2018 (60) and TwitterDataset (50)), anti-solidarity (restriction or exclusion of groups), mixed (both supportive and opposing elements), and none (neutral or unrelated content).
-Instances labeled as solidarity or anti-solidarity are further assigned one of four frame types adapted from thijssen2022s (98): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance in contributions), and empathic (respect for or rejection of social diversity).
-We extend the evaluation to a newly annotated set of 756 instances for Migrant with more historical coverage (before 1950).
-For this set, annotations were conducted by 4 annotators following the same procedure as in kostikova-fine (57) Statistics on the human-annotated dataset are presented in Fig.
-3.
-Overall, the label distribution differs notably by target group: woman-related instances are predominantly labeled as solidarity (74.1%), while migrant-related instances show a more diverse pattern, with lower solidarity (51.6%) and a higher share of anti-solidarity (21.3%).
-Figure 3: Distribution of instances in the human annotated dataset across time and target groups.
-Total indicates number of all annotated samples for the given target group, including the None label.
-See Table 2 for the actual numbers of instances in the human annotated dataset.
+To validate our models and enable qualitative analysis, we use a human-annotated dataset created in two batches.
+The first, introduced in Kostikova et al.
+(2024), comprises 2,864 instances (1,428 migrant-related, 1,437 woman-related) from DeuParl, annotated by five trained student assistants.
+The annotation guidelines were refined over multiple rounds of discussion among the annotators and authors; details of the annotation process and the final guidelines are provided in Appendix˜F and Appendix˜G in the Appendix.
+The second adds 756 migrant-related instances with broader historical coverage (pre-1950), annotated by four of the same assistants following the same procedure.
+Each instance is labeled according to a fine-grained solidarity framing scheme distinguishing four high-level stance categories – solidarity, defined as willingness to share resources Lahusen and Grasso (2018); Ils et al.
+(2021), anti-solidarity (restriction or exclusion), mixed (both supportive and opposing elements), and none (neutral or unrelated content).
+Solidarity and anti-solidarity are further assigned one of four subtypes adapted from Thijssen and Verheyen (2022): group-based (shared identity or exclusion), compassionate (protection or dismissal of marginalized groups), exchange-based (valuing contributions or claiming imbalance), and empathic (respect for or rejection of diversity).
+We refer to the subtypes as fine-grained annotation.
+The final gold-standard label for each instance is determined by majority vote.For the first batch, inter-annotator agreement reached a Cohen’s $\kappa$ of 0.62 for the high-level labels and 0.42 for the fine-grained labels.
+For the second batch, agreement reached 0.47 for the high-level labels and 0.37 for the fine-grained labels.
+We attribute the lower agreement in the second batch to the inclusion of older historical data, which we discuss further in Section˜3.1.
+Statistics on the human annotated dataset are shown in Figure˜3.
+Label distributions differ by target group: woman-related instances are predominantly solidarity (74.1%), while migrant-related instances are more diverse, with less solidarity (51.6%) and more anti-solidarity (21.3%).
+Figure 3: Distribution of human annotated instances by time and target group. Total indicates the number of instances per group (including None); see Table˜6 in the Appendix for exact counts.
 2.3 Models and Experiments
-We evaluate a diverse suite of models, ranging from smaller supervised models (BERT, fine-tuned SBERT) to proprietary LLMs (GPT-4) and recent open-weight instruction-tuned models.
-The final list includes:
+We evaluate a diverse set of models to cover variation in model size, provider, and access setting (open vs. closed), while also including supervised classifier baselines. Our selection comprises:
   1. (i)
-    smaller pre-trained LMs like BERT and SBERT;
+    smaller pre-trained LMs such as BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details are provided in Appendix˜C in the Appendix;
   2. (ii)
-    closed-source LLMs like gpt-4-1106-preview;
+    closed-source LLMs such as gpt-4-1106-preview and gpt-5-chat;
   3. (iii)
-    open-weight LLMs such as Llama-3-8B-It, Llama-3.3-70B-It (and fine-tuned Llama-3-70B-It), Gemma-2-9b-It, Gemma-2-27b-It, and Mistral-2-Large.
+    open-weight LLMs, including:
+    • gpt-oss-20B and gpt-oss-120B;
+    • Mistral-2-Large (123b parameters);
+    • Llama-4-Scout-17B-16E-Instruct;
+    • Llama-3.3-70B-It, Llama-3-8B-It;
+    • Qwen2.5-72B-Instruct;
+    • Gemma-2-9B-It and Gemma-2-27B-It.
 Data and Metrics
-We adopt the 70/15/15 train/dev/test split strategy as in kostikova-fine (57), generating three random splits of approximately 2,000 training, 400 development, and 430 test instances for each target group.
-These are used to train and evaluate supervised models.
-For inference-only experiments, we use only the test sets: Test 1 (original) and Test 2 (newly added, 756 instances), the latter offering broader historical coverage (1860s–1950s).
-We report Macro F1 to account for class imbalance, along with per-class F1 scores for both high-level and fine-grained classification tasks.
-[^2] High-level classification refers to the four stance categories: solidarity, anti-solidarity, mixed, and none.
-Fine-grained classification includes ten labels: the four subtypes of solidarity, the four subtypes of anti-solidarity, plus mixed and none.
-Models
-We use manually designed prompts for each target group.
-Each prompt consists of a classification instruction concatenated with the parliamentary instance, i.e., the keyword sentence plus its three-sentence left/right context as defined in 2.1.
-Note that, although speaker and party metadata are available in the corpus, this information is removed from the model input.
-The classification instruction includes (i) chain-of-thought reasoning in zero- and few-shot settings (via think step by step); (ii) label definitions; and (iii) a two-step classification format: high-level first, followed by subtype for solidarity or anti-solidarity prediction.
-[^3] For GPT-4, both steps are combined in a single prompt (see Figure 21 for Migrant and Figure 22 in the Appendix for Woman).
-For open-weight models, we instead use two prompts: one for high-level classification and, conditional on that output, a second for subtype selection (see Figure 24 for Migrant and Figure 27 in the Appendix for Woman).
-In few-shot mode, prompts include one representative example per label.
-We apply these prompts to closed-source models like gpt-4-1106-preview, and to open-weight LLMs including Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it (full precision) in inference-only setting.
-Larger models (Llama-3.3-70B-Instruct and Mistral-Large-Instruct-2407) are run with $Q4_{0}$ quantization.
-Additionally, we fine-tune Llama-3-8B-It and Llama-3-70B-It using a three-model pipeline (high-level and two subtype classifiers), applying instruction-style fine-tuning with true labels paired with GPT-4-generated explanations as training targets.
-We use BERT and SBERT as supervised baselines. Full model configurations, fine-tuning setups, and inference details for all models used in our experiments are provided in C.
+We adopt a 70/15/15 train/dev/test split of the first annotated batch (Kostikova et al., 2024), generating three random splits of about 2,000 training, 400 development, and 430 test (Test 1) instances in total, covering both target groups.
+Supervised models are trained and evaluated on these splits, while inference-only experiments use two test sets: Test 1 (original) and Test 2 (756 newly added migrant-related instances, 1860s–1950s).
+We report Macro F1 to account for class imbalance, along with per-class F1 for high-level (four stance categories) and fine-grained (ten labels) classification.
+[^2] High-level: solidarity, anti-solidarity, mixed, none.
+Fine-grained: four solidarity subtypes, four anti-solidarity subtypes, plus mixed and none.
+Prompting Setup
+For the inference-only experiments with LLMs, We use manually designed prompts for each target group.
+Each prompt consists of a classification instruction concatenated with the parliamentary instance (the keyword sentence plus three sentences of left/right context as defined in Section˜2.1).
+Speaker and party metadata are available in the corpus but removed from the model input.
+Instructions include (i) chain-of-thought cues in zero-/few-shot settings (think step by step); (ii) label definitions; and (iii) a two-step format: high-level classification followed by subtype prediction.
+[^3] For GPT-4, both steps are combined in a single prompt (Figure 25 and Figure 26 in the Appendix); for open-weight models, we use two prompts: one for high-level classification and, conditional on that, a second for subtype selection (Figure 27 and Figure 30 in the Appendix).
+In few-shot mode, one representative example per label is included.
+Model Ensemble
+We additionally evaluate a model ensemble to test whether combining strong open-weight LLMs improves performance relative to individual models.
+The ensemble setup also resembles the human annotation procedure in that final labels are derived from multiple judgments rather than a single annotator.
+The ensemble combines the best-performing open models, each in its strongest configuration (Table˜1), using majority vote.
+The label with the most votes is selected as the final prediction; in tie cases, votes are weighted by each model’s macro-F1 score as measured on the validation set.
+We report ensemble results in Section˜3.1.3.
+Design-based Supervised Learning (DSL)
+When employing LLMs as annotators, it may be that single labels are systematically preferred compared to human annotators, yielding biased estimates and incorrect conclusions.
+We use Design-based Supervised Learning (DSL; Egami et al.
+2023) to correct such systematic biases.
+DSL combines large-scale surrogate outputs with a subsample of human-coded data to construct bias-corrected pseudo-outcomes for inference.
+In our application, the surrogate outputs are soft label distributions derived from a committee of LLMs.
+For each parliamentary speech $i$, we observe surrogate labels $Q_{i}$ and covariates $X_{i}$.
+For a (small) subset, we also have human annotations $Y_{i}$.
+The surrogate labels $Q_{i}$ are soft-label distributions over annotation categories, computed as the proportion of committee models assigning each label, while the human labels $Y_{i}$ are distributions over categories constructed from multiple human annotators.
+Human annotations are observed only for a subset of instances under a decade-stratified annotation design.
+We assign each instance a decade-specific inclusion probability $\pi_{i}$, computed as the proportion of instances in that decade that received human annotation, and treat these as fixed design probabilities in the DSL correction.
+DSL corrects downstream estimands by constructing a bias-adjusted pseudo-outcome
+$\tilde{Y}_{i}=\hat{g}(Q_{i},X_{i})+\frac{R_{i}}{\pi_{i}}\big(Y_{i}-\hat{g}(Q_{i},X_{i})\big),$ (1)
+where $R_{i}$ indicates whether instance $i$ is human-annotated, and $\hat{g}(Q_{i},X_{i})$ estimates the conditional expectation of human labels given surrogate outputs and covariates.
+We model $\hat{g}$ using a multinomial regression trained on soft targets and implement $K$-fold cross-fitting.
+Because outcomes are multi-class, DSL is applied in a vector-valued form, giving a vector of DSL-adjusted class scores $\tilde{Y}_{i}$ for each instance. We use them directly in downstream analyses.
+We estimate separate DSL models for the high-level and fine-grained tasks to avoid carrying subtype-level error from fine-grained predictions into aggregated high-level estimates.
 3 Results
-In this section, we evaluate models against a human-annotated dataset to determine reliability for social scientific analysis of solidarity frames in parliamentary texts.
-Our aim is twofold: (i) to identify a model whose predictions approximate human-level annotation quality across both high-level and fine-grained stance categories, and (ii) to ensure the selected model is practically feasible to deploy at scale.
-Specifically, we ask whether recent open-weight models can match the performance of GPT-4, which was previously found to be highly reliable (though costly) kostikova-fine (57), through techniques such as few-shot prompting or lightweight fine-tuning.
-3.1 Model Performance Overview
-Table 1 reports macro F1 scores for each model across both inference-based and fine-tuned approaches.
-We benchmark these results against a human upper bound, estimated using a leave-one-out (LOO) procedure, [^4] For each annotator, we exclude their labels, compute a consensus label from the remaining annotators, and calculate the macro F1 score between the excluded annotator and the consensus.
-We then average these scores across all annotators.
-and report performance for both high-level and fine-grained classification tasks, with the latter shown in parentheses.
-Our observations are structured around two dimensions: (i) overall model performance, (ii) task and label-related challenges.
+In this section, we evaluate the models in two steps.
+We first assess instance-level annotation performance against human labels, including overall performance, task-specific challenges, and systematic error patterns in Section˜3.1.
+We then examine whether these model outputs support reliable downstream inference by comparing hard-label, soft-label, and DSL-corrected trend estimates against human-derived reference trends in Section˜3.2.
+3.1 Annotation Performance at the Instance Level
+Table˜1 reports macro F1 scores for each model under inference and fine-tuned settings.
+We benchmark against the leave-one-out F1 score of human annotators, [^4] For each annotator, we exclude their labels, compute a consensus from the others, and calculate the macro F1 score between the excluded annotator and the consensus; scores are then averaged across annotators.
+and show results for both high-level and fine-grained tasks (the latter in parentheses).
+We evaluate instance-level annotation performance along four dimensions: (i) overall performance of individual models; (ii) task- and label-related challenges; (iii) model ensemble performance; and (iv) error analysis.
 Model Method Woman Migrant (T1) Migrant (T2)
+Closed-weight models
 GPT-4 Zero-shot 0.37 (0.60) 0.42 (0.73) 0.29 (0.62)
-Few-shot 0.37 (0.54) 0.43 (0.63) 0.29 (0.51)
+Few-shot 0.37 (0.54) 0.43 (0.63) 0.26 (0.55)
+GPT-5 Zero-shot 0.36 (0.53) 0.42 (0.67) 0.30 (0.60)
+[…]
+Large open-weight models
+gpt-oss-20B Zero-shot 0.25 (0.43) 0.27 (0.55) 0.15 (0.49)
+Few-shot 0.28 (0.47) 0.40 (0.56) 0.22 (0.49)
+gpt-oss-120B Zero-shot 0.28 (0.55) 0.38 (0.66) 0.25 (0.55)
+[…]
 Mistral-2-Large Zero-shot 0.20 (0.50) 0.27 (0.56) 0.24 (0.49)
 Few-shot 0.26 (0.51) 0.38 (0.57) 0.26 (0.46)
+Llama-4-Scout Zero-shot 0.25 (0.46) 0.28 (0.52) 0.20 (0.50)
+Few-shot 0.25 (0.46) 0.41 (0.63) 0.25 (0.50)
+Qwen-2.5 Zero-shot 0.21 (0.40) 0.37 (0.62) 0.23 (0.52)
+Few-shot 0.22 (0.46) 0.43 (0.59) 0.27 (0.51)
 Llama-3.3-70B Zero-shot 0.23 (0.50) 0.32 (0.54) 0.25 (0.47)
-Few-shot 0.30 (0.50) 0.37 (0.65) 0.28 (0.53)
-Llama-3-70B Fine-tuned 0.26 (0.49) 0.27 (0.65) 0.21 (0.47)
-Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.17 (0.40)
-Few-shot 0.14 (0.35) 0.19 (0.34) 0.15 (0.40)
-Fine-tuned 0.16 (0.38) 0.20 (0.40) 0.13 (0.27)
-Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.25 (0.45)
-Few-shot 0.25 (0.55) 0.33 (0.48) 0.26 (0.46)
-Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.22 (0.46)
-Few-shot 0.23 (0.44) 0.31 (0.54) 0.24 (0.46)
-BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.23 (0.45)
-SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.18 (0.36)
-Human upper bound 0.34 (0.62) 0.48 (0.75) 0.35 (0.57)
+[…]
+Mid- and small-scale open-weight models
+Llama-3-8B Zero-shot 0.12 (0.35) 0.19 (0.41) 0.15 (0.40)
+Few-shot 0.14 (0.35) 0.19 (0.34) 0.12 (0.42)
+Gemma-2-27B Zero-shot 0.20 (0.48) 0.22 (0.48) 0.21 (0.46)
+Few-shot 0.25 (0.55) 0.33 (0.48) 0.21 (0.47)
+Gemma-2-9B Zero-shot 0.22 (0.45) 0.24 (0.44) 0.21 (0.47)
+Few-shot 0.23 (0.44) 0.31 (0.54) 0.23 (0.47)
+Classical baselines
+BERT (2000 train instances) 0.13 (0.26) 0.24 (0.46) 0.22 (0.48)
+SBERT (2000 train instances) 0.15 (0.39) 0.23 (0.43) 0.17 (0.37)
+Human reference 0.32 (0.67) 0.48 (0.78) 0.30 (0.58)
 Table 1: Comparative performance (macro F1) of models on high-level (in parentheses) and fine-grained tasks.
 ?Migrant (T1)?
-uses the original benchmark; ?Migrant (T2)?
+is the original benchmark; ?Migrant (T2)?
 is newly introduced in this paper.
-Green cells indicate the best-performing score for each target group, based on a model’s best setting (e.g., zero-shot or few-shot); yellow cells indicate the best-performing score from a different model (i.e., the second-best model overall).
-  1. (i)
-    Overall model performance.
-    • GPT-4 achieves the highest scores, but Llama-3.3-70B is a strong alternative.
-Across all target groups (Woman and both Migrant test sets), GPT-4 (zero-shot) consistently shows the top performance, coming closest to the human upper bound in both high-level and fine-grained classification.
-Llama-3.3-70B is the best-performing open model, showing meaningful gains from few-shot prompting (in particular in the high-level classification).
+Green cells mark the best score per target group (model’s best setting); yellow cells mark the best score among open models (second best overall).
+Gray text indicates suboptimal settings for a given model.
+3.1.1 Overall model performance
+  • GPT-5 leads, but some open models offer a strong alternative.
+Across Woman and both Migrant test sets, GPT-5 (few-shot) matches human annotators on the high-level task and surpasses them on the fine-grained task.
+Among open models, gpt-oss-120B matches the human reference value in fine-grained performance and comes close to GPT-4 and GPT-5 at the high level.
+    Among the remaining open-weight models, Llama-4-Scout, Qwen-2.5, and Llama-3.3-70B show the strongest performance.
+These models achieve comparable results across targets and evaluation levels (e.g., on Migrant T1, fine-grained F1 scores range from 0.38 to 0.43 and high-level scores from 0.59 to 0.65).
+While their relative strengths vary by setting, where Qwen-2.5 and Llama-4-Scout achieve the strongest fine-grained scores, and Llama-3.3-70B the highest high-level scores, their overall performance remains closely clustered.
   • Model size correlates with performance, but not perfectly.
-Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure 4 for the relationship between model size and performance).
-Mid-sized models follow, and smaller models perform worst.
-However, this relationship is not strictly linear: for example, Llama3-8B-it underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B-it.
+Larger models generally outperform smaller ones on both high-level and fine-grained tasks (see Figure˜4).
+However, this relationship is non-monotonic: moving from small to mid-sized models brings clear benefits, with notable exceptions such as Llama3-8B, which underperforms relative to smaller baselines like BERT and the similarly sized Gemma-9B.
+Among large models ($\approx$70B–120B+ parameters), performance gains diminish.
+Moreover, mid-sized models can match or outperform larger systems; for example, gpt-oss-20B clusters with Qwen-72B and Mistral-Large-2, while gpt-oss-120B approaches GPT-4/5 on the fine-grained benchmark.
+These results suggest that factors beyond parameter count influence model performance.
     Figure 4: Model size vs.
-performance on high-level and fine-grained classification tasks in Migrant (T1).
+performance on high-level and fine-grained classification in Migrant (T1).
 Each model is plotted with its best-performing configuration (zero- or few-shot).
 The x-axis is logarithmic, so spatial distances do not represent linear differences in parameter count.
-    • Examples help mid-sized models more than small or very large ones.
-Few-shot prompting improves performance primarily for mid-sized models such as Gemma-9B-it, Gemma-27B-it, Mistral-2-Large, and Llama-3.3-70B-it.
-In contrast, smaller models like Llama-3-8B-it show no benefit, and GPT-4 sees minimal gains—likely because its zero-shot performance is already near-optimal.
-Fine-tuned variants (e.g., Llama-3.3-70B-
-      it
-      , Llama-3-8B-it) often show no improvement over inference-only models.
-  2. (ii)
-    Task and label-related challenges.
-    • All models perform worse on the fine-grained task, underscoring its complexity: unlike the more intuitive high-level categories (e.g., solidarity, anti-solidarity), it requires alignment with annotation guidelines grounded in social theory, where even human annotators show limited agreement (reaching macro F1 score 0.34 for Woman and 0.48 for Migrant).
-    • At the high level, solidarity is the easiest to detect (see Figure 5; full scores in Table 3 in the Appendix), especially in the Woman category, where even smaller models like BERT (0.77) and Gemma-9B (0.78) perform strongly.
-Anti-solidarity and mixed stance are the most challenging categories, with GPT-4 achieving the highest high-level anti-solidarity scores (0.64 for Migrant, 0.65 for Woman), but mid-size models (e.g., Mistral, Llama-3.3-70B, Gemma-27B, Gemma-9B) show moderate performance (with scores ranging from 0.52 to 0.54).
-Only GPT-4 and Llama-3.3-70B reliably handle mixed stances (e.g., 0.38–0.43 for Migrant, 0.30–0.37 for Woman).
-    • On the fine-grained level, GPT-4, Llama-3.3-70B, and Gemma models offer the most balanced coverage across subtypes.
-GPT-4 stands out on exchange-based solidarity (0.56), group- and exchange-based anti-solidarity for Migrant (0.47 and 0.46), and on empathic and compassionate anti-solidarity for Woman (0.67 and 0.50).
-Among open models, Llama-3.3-70B performs particularly well on compassionate solidarity (0.51), and shows strong results on group- and exchange-based anti-solidarity (0.42 and 0.35).
-Interestingly, the Gemma models are able to distinguish the rare compassionate anti-solidarity subtype for Migrant with some success.
-Most other models struggle with fine-grained anti-solidarity, with scores close to zero.
-      Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For the Migrant group, the two test sets (Test 1 and Test 2) are combined.
+  • Examples help mainly mid-sized models.
+Most models benefit from few-shot prompting, with the most consistent improvements observed for mid-sized models such as Gemma models, Mistral-2-Large, Llama-3.3-70B and Llama-4-Scout.
+Smaller models like Llama-3-8B show only marginal gains, while GPT-4 demonstrates minimal improvement, likely due to its strong zero-shot performance.
+Overall, the effect of few-shot prompting is widespread but not uniform across model scales.
+3.1.2 Task and label-related challenges
+  • All models (and human annotators) perform worse on the fine-grained task, which indicates its complexity.
+  • Solidarity is easier to detect than anti-solidarity or mixed stance (Figure˜5).
+Even smaller models perform well on solidarity, especially in the Woman category (e.g., BERT 0.77, Gemma-9B 0.70).
+In contrast, anti-solidarity and mixed stances remain challenging, with anti-solidarity being most difficult for Woman.
+This pattern is likely due in part to class imbalance (Table˜6 in the Appendix), especially because anti-solidarity is very rare for Woman and mixed is less common in both target groups.
+That said, GPT models (both closed-source models and gpt-oss-120B) achieve the strongest anti-solidarity performance (0.66–0.70 for Migrant; 0.53–0.65 for Woman), while mid-sized models reach 0.56–0.59.
+Only GPT-4/5, gpt-oss-120B and Llama-3.3-70B consistently capture mixed stances, with F1 scores between 0.30 and 0.50.
+Llama-4-Scout shows some asymmetry for Migrant: strong performance on solidarity (0.78) and anti-solidarity (0.63) but near failure on mixed stance (0.08).
+  • On the fine-grained level, GPT-5, gpt-oss-120B, Llama-3.3-70B and Llama-4-Scout offer the most balanced performance on the fine-grained level (see Table˜7 in the Appendix).
+Best fine-grained macro F1 is achieved by GPT-5.
+Among non-GPT-4/5 models, Llama-3.3-70B and Llama-4-Scout follow.
+Nevertheless, all models perform poorly on specific subtypes, including empathic solidarity (e.g., 0.11 for Migrant with Llama-3.3-70B) and compassionate anti-solidarity (e.g., 0.06 for Migrant with gpt-oss-120B).
+Empathic anti-solidarity fails across all models.
+For Woman, all models consistently struggle with fine-grained anti-solidarity.
+    Figure 5: F1 scores for high-level labels based on the best-performing configuration for each model. For Migrant, the two test sets (Test 1 and Test 2) are combined.
   • Performance drops on historical data.
-All models perform worse on the historical Migrant (T2) set, including the strongest ones.
-This reflects a broader pattern: model performance tends to decline on older data, as shown by macro F1 scores over time (see Figure 16)—so does agreement among human annotators (Cohen’s Kappa).
-However, while human trends are smooth and monotonic, model trajectories are less stable.
-Stronger models such as GPT-4, Mistral-2-Large, Llama-3.3-70B, and BERT show a general upward trend but exhibit clear dips in the 1930s and 1960s.
-In contrast, smaller models (e.g., Llama-3-8B, SBERT, Gemma-9B) show little to no temporal coherence.
-In summary, GPT-4 remains the strongest model overall.
-While performance generally scales with model size, mid-sized models can approximate GPT-4 when given examples.
-Among open-weight alternatives, Llama-3.3-70B emerges as the most viable option, particularly for the Migrant category, delivering competitive results in both high-level and fine-grained classification, while offering a more cost-effective solution.
-We therefore adopt it for the subsequent social scientific analyses.
-However, since it still lags behind GPT-4 in capturing certain fine-grained subtypes, we further validate its reliability by comparing the temporal patterns produced by both models in 4.1.
-3.2 Error Analysis
-In this subsection, we examine the errors made by the two strongest models in our experiments, GPT-4 and Llama-3.3-70B, as determined by macro F1 scores.
-To identify where these models struggle, we compare their confusion matrices on the full Migrant test set (Figure 6 for high-level; see Figure 17 in the Appendix for confusion matrices for all models included in the study and Figure 18 in the Appendix for fine-grained label confusion matrices) with a human reference, constructed by averaging confusion matrices from four leave-one-out (LOO) comparisons among annotators.
-Figure 6: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined.
-Human Annotation scores are computed by averaging the confusion matrices from four LOO comparisons, where each annotator is compared to the consensus of the others.
-Both models exhibit notable differences from human annotators.
-In particular, both models frequently annotate true Solidarity instances as None (108 instances for GPT-4, 128 for Llama) and true Anti-solidarity instances as None (59 for GPT-4, 60 for Llama), especially for common subtypes such as group-based and compassionate.
-These errors are far less frequent in human annotations (32 and 12, respectively), suggesting that models may more often underestimate both solidarity and anti-solidarity.
-Llama additionally shows a tendency to misclassify true Solidarity as Anti-solidarity (30 cases vs.
-11 for GPT-4 and 13 for human annotators) and, to a lesser extent, true Anti-solidarity as Solidarity (18 vs.
-4 for GPT-4 and 10 for humans).
-Some of these misclassifications arise when a speaker cites an anti-solidarity statement by another party, which Llama occasionally misinterprets as anti-solidarity (see Examples 9 and 11 in Table 14 in the Appendix; these are discussed in more detail in Section 4.3.
-However, manual inspection indicates that such cases are rare.
+All models, including the strongest, perform worse on the historical Migrant (T2) set.
+This reflects a broader trend: performance declines on older data (see Figure˜18 in the Appendix), as does human agreement (Cohen’s Kappa).
+We examine errors on the historical data in more detail in Section˜3.1.4.
+3.1.3 Model Ensemble Performance
+We construct a majority-vote ensemble from the best-performing open models in Table˜1: Llama-3.3-70B, Qwen-2.5, and gpt-oss-120B.
+We exclude GPT-5 due to its higher cost and because the subsequent social-scientific analysis focuses on open models.
+We find that the ensemble does not outperform the strongest individual model (gpt-oss-120B), indicating that simple majority voting does not consistently improve performance in this setting (see Table˜2).
+This is in line with prior work showing that ensemble methods are most effective when individual models achieve comparable performance while making at least partly uncorrelated errors (Trad and Chehab, 2024).
+Model Woman Migrant (T1) Migrant (T2)
+gpt-oss-120B (Few-shot) 0.38 (0.55) 0.48 (0.69) 0.28 (0.57)
+Llama-3.3-70B (Few-shot) 0.30 (0.50) 0.38 (0.65) 0.29 (0.58)
+Qwen-2.5 (Few-shot) 0.22 (0.46) 0.43 (0.59) 0.27 (0.51)
+Ensemble 0.34 (0.60) 0.47 (0.70) 0.29 (0.60)
+Table 2: Performance (macro F1) of models included in the ensemble and the resulting majority-weighted ensemble on fine-grained and high-level tasks (in parentheses).
+In tie cases, votes are weighted by validation-set macro-F1.
+Tie-breaking weights are based on validation-set performance, ranking gpt-oss-120B >> Llama-3.3-70B >> Qwen-2.5 for the migrant setting and Llama-3.3-70B >> gpt-oss-120B >> Qwen-2.5 for the woman setting.
+In summary, GPT-5 achieves the highest overall performance, surpassing human annotators on fine-grained level and approaching them on the high-level task.
+However, its closed-weight nature and high cost make it less suitable for our large-scale social scientific analysis.
+Among open models, gpt-oss-120B performs comparably to GPT-5 and approaches human-level performance (even matching it on the fine-grained level), making it the strongest open-weight alternative.
+That said, the fact that these models reach or exceed human reference values does not imply that they can be used as direct substitutes of human annotators.
+First, the observed macro-F1 scores remain low enough that substantial and systematic model errors can still occur, including errors that human annotators are less likely to make; moreover, recent work shows that even high agreement with human consensus can coexist with systematic misclassification patterns (Bavaresco et al., 2025).
+We therefore examine these patterns through error analysis (Section˜3.1.4).
+Second, in our setup, the human reference is derived from leave-one-out agreement and therefore inherits the uncertainty present in the human annotations.
+As a result, perfect F1 may itself be unattainable because some disagreement among human annotators likely reflects genuine label variation.
+In a subjective task like ours, this motivates moving beyond hard labels to consider label distributions instead (Section˜3.2.1).
+We address both issues next.
+3.1.4 Error Analysis
+We analyze errors of the strongest models through their confusion matrices on the full Migrant test set, comparing them with a human reference aggregated from leave-one-out (LOO) annotator comparisons.
+[^5] Because the human panel compares each annotator to a LOO consensus, whereas the model panels compare predictions to gold labels, we use row-normalized confusion matrices so that error patterns for each reference label are more directly comparable.
+Figure˜6 shows the top models; full matrices appear in Figure˜19 and Figure˜20 in the Appendix.
+We consider four aspects of model error: (i) high-level confusion; (ii) confusion within (anti-)solidarity subtypes; (iii) errors in historical instances (1867-1900); and (iv) ensemble failure analysis, with particular attention to error correlation across models.
+Since our substantive analysis focuses on migrant-related discourse, we restrict the error analysis to the Migrant category.
+(a) High-level confusion matrices.
+(b) Fine-grained confusion matrices.
+Figure 6: Row-normalized confusion matrices (%) for the best-performing configuration of each of the top models by macro-F1, as reported in Table˜1, on the two Migrant test sets combined (Test 1 and Test 2).
+Each row sums to 100, showing how items with a given reference label are distributed across predicted labels.
+Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators.
+The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
+(i) High-level confusion
+Most models recover the broad high-level structure of the task.
+GPT-5 and gpt-oss-120B show the clearest separation between solidarity, anti-solidarity, and none, but even these stronger models do not fully match human error patterns.
+Relative to the human LOO panel, models systematically overassign cases to none, suggesting that they often fail to detect low-intensity or indirect stance expressions, despite the prompt instruction to consider even slight expressions of solidarity regardless of the main topic of the text Example 11 in Table˜14).
+The models also preserve mixed much less often: for reference-mixed cases, the share predicted as mixed is 42% for humans, compared with 27% for GPT-5, 22% for gpt-oss-120B and Qwen, and 33% for the ensemble; Llama-3.3-70B is the only exception, reaching 47%.
+Instead, models resolve these cases into one-sided solidarity or anti-solidarity labels, suggesting that models overlook weaker cues of the opposing stance, as in Example 11.
+A smaller but qualitatively important error type arises when speakers from party A quote or paraphrase anti-migrant rhetoric from party B to criticize party B and its stance.
+In such cases, models sometimes misclassify the speaker’s stance as anti-solidarity (e.g.
+Llama and Qwen in Examples 11 and 12).
+(ii) Confusion within (anti-)solidarity subtypes
+Within solidarity, and focusing on the better-represented labels (Table˜6), human disagreement is greatest between group-based and compassionate solidarity: in the human LOO matrix, 18% of group-based solidarity cases are labeled as compassionate solidarity by the held-out annotator.
+Models show less confusion in this direction (5-7%), but more often predict group-based solidarity for human-labeled compassionate solidarity instead (e.g.
+32% for gpt-oss-120B).
+Humans also identify compassionate solidarity more consistently overall (66% vs.
+roughly 40% for models).
+The two subtypes can co-occur in migration debates.
+Human annotators reflect this ambiguity, whereas models tend to shift such cases toward group-based solidarity, apparently relying more heavily on cues such as integration, rights, and legal inclusion (Examples 12, 12, and 13).
+However, rights language alone does not imply group-based solidarity.
+When rights claims are framed in terms of vulnerability, deprivation, or need, compassionate solidarity is the stronger reading, as it aims to place the other ?in an equal position as I am?.
+When such claims are framed in terms of shared membership or common rights and duties, group-based solidarity is more appropriate (Thijssen, 2012; Thijssen and Verheyen, 2022).
+A related pattern is observed for other subtypes: confusion occurs when broader integration language associated with group-based solidarity overlaps with characteristics of other frames, as in Example 13.
+Since Thijssen and Verheyen (2022) describe the categories as conceptually close and the coding as interpretive, we treat such variation as meaningful uncertainty rather than noise.
+Within anti-solidarity, the sharpest subtype confusions appear in the rarer categories, especially empathic and compassionate anti-solidarity.
+These cases are often collapsed into group-based anti-solidarity, suggesting that models over-rely on broad cues of protecting the interests of the native in-group (see Example 14).
+Nevertheless, some models also show confusion among the better-represented anti-solidarity categories such as exchange-based anti-solidarity and group-based anti-solidarity (22% cases for Llama-3.3-70B, 21% for humans, see Example 13).
+(a) High-level confusion matrices.
+(b) Fine-grained confusion matrices.
+Figure 7: Row-normalized high-level and fine-grained confusion matrices (%) for the best-performing configurations of the top models on the combined Migrant test sets, restricted to 1867-1900.
+As in Figure˜6, rows sum to 100, Human Annotators (LOO) denotes aggregated leave-one-out annotator comparisons, and the Ensemble aggregates Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
+(iii) Confusion for historical instances
+The historical sample (restricted to 1867-1900, Figure˜7) shows an even stronger overprediction of none.
+Across nearly all models, errors for many true classes collapse into none, more than in the human LOO panel; this is particularly visible for Qwen-2.5-72B, Llama-3.3-70B, and the ensemble.
+Human disagreement also increases on the historical subset, butannotators more often distribute errors across neighboring stance categories.
+Examples 14 and 14 indicate why this matters for historical inference.
+Here, solidarity is expressed through legal and constitutional reasoning rather than overt emotional language, and models fail to recognize such low-intensity stance at all.
+Overall, strong models reproduce the broad label structure of the task, but their errors remain systematic rather than random.
+As a result, trends in (anti-)solidarity and mixed stance are likely to be underestimated, especially in older instances.
+Within subtypes, group-based solidarity is likely to absorb other solidarity claims, while rarer categories are often collapsed or disappear entirely.
+The ensemble mitigates these biases only modestly, if at all (e.g.
+it still overpredicts none and underpredicts mixed relative to humans).
+This highlights the need for correction techniques such as DSL, which we analyze in Section˜3.2.2.
+(iv) Ensemble failure analysis
+We begin by asking how often ensemble errors contain any useful signal at all.
+On the combined Test 1 and Test 2 data, the ensemble makes 1,112 errors.
+Of these, 747 (67.2%) occur on instances where all models are wrong, whereas only 365 (32.8%) occur when at least one model predicts the correct label.
+Thus, most ensemble failures occur on examples where none of the component models predicts the correct label.
+The remaining 365 cases suggest that complementary signal exists in principle.
+To quantify how much performance could be recovered from this signal, we construct an oracle that selects a correct model prediction whenever at least one model is correct.
+The oracle reaches markedly higher macro-F1 than the ensemble–0.66 (0.87) on Test 1 and 0.45 (0.73) on Test 2, compared with 0.47 (0.70) and 0.27 (0.56), respectively.
+This indicates that although useful complementary signal is present in the model pool, simple ensembling fails to recover much of it in practice, which explains the ensemble’s substantially lower empirical F1.
+A closer look at model overlap helps explain why this signal is difficult to exploit.
+Across the pooled Test 1 and Test 2 data, pairwise error Jaccard scores are consistently high (0.57-0.64; see Appendix˜D in the Appendix and Table˜4), indicating that models often fail on the same instances.
+Moreover, conditional on two models both being wrong, they predict the same wrong label in roughly half of cases (45-49%), which means that shared errors are often reinforced rather than corrected.
+Taken together, these results suggest that the main limitation is limited error diversity across models on difficult instances.
+Section˜3.2 examines how the biases discussed in this section affect downstream trend estimates.
+3.2 Improving Downstream Reliability
+The preceding analyses identify two challenges for using LLM annotations in downstream inference.
+First, disagreement between human annotators does not necessarily reflect error, but may indicate intrinsic ambiguity and multiple defensible interpretations.
+Following work that discusses such cases under the umbrella term of label variation (Plank, 2022), we therefore move beyond hard labels and compare label distributions.
+Second, the error analysis shows that model predictions are systematically biased, with some labels over- or under-predicted relative to human annotations.
+Such bias can distort downstream analyses when predictions are aggregated into temporal or group-level trends.
+We address both issues in this section: we first assess the similarity of model and human label distributions (Section˜3.2.1), and then test whether debiasing with DSL improves recovery of human-derived trends (Section˜3.2.2).
+3.2.1 Label distribution similarity
+For each instance with at least three human annotations, we derive a human label distribution from the annotators’ votes.
+For the model ensemble, we analogously construct a label distribution from the votes of the three models (Llama-3.3-70B, Qwen-2.5, and gpt-oss-120B).
+For individual models, we use sharp distributions that assign 100% probability mass on the predicted label.
+We compare these distributions using Jensen–Shannon divergence (JSD) and total variation distance (TV), where lower values indicate higher similarity (see Appendix˜E in the Appendix for definitions).
+The analysis covers both the high- and fine-grained levels.
+We compare model–human distances against three baselines: (i) human-human distance, averaged over ten random annotator splits; (ii) a randomized model baseline, with predictions shuffled across instances; and (iii) single-model baselines comparing each model to the human distributions.
+Table˜5 in the Appendix shows that the ensemble (TV 0.49; JSD 0.39) is closer to human label distributions (TV 0.45; JSD 0.35) than any individual model (TV 0.57–0.59; JSD 0.48–0.50) and far closer than the randomized baseline (TV 0.75; JSD 0.69), with the same pattern observed at the high level.
+These results indicate that strong LLMs reproduce human-like patterns of disagreement.
+At the same time, the ensemble–human distance does not surpass the human–human baseline.
+These results have the following implications: First, since model disagreement parallels human uncertainty, this variation should be preserved through soft labels rather than reduced to single hard predictions.
+Second, because models do not surpass human–human similarity and their errors remain correlated, these soft labels still contain structured bias.
+As a result, model predictions cannot be used directly without correcting for bias.
+We address this issue in the next section.
+3.2.2 Bias Correction with DSL
+We validate all trend estimators considered below using the complete set of 2,184 Migrant-related instances that were annotated by humans (hereafter the evaluation subset; see Table˜6 in the Appendix).
+Trends computed from these annotations serve as the reference.
+Because the evaluation subset is fully human-annotated, we simulate a partial-label setting for DSL evaluation.
+Specifically, we randomly retain human labels for only a subset of instances within each decade and treat the remaining labels as unobserved.
+We then apply DSL using the corresponding decade-specific sampling probabilities and compare the resulting trends with those computed from the full human annotations.
+We compare the human-derived reference trends against several alternative estimators.
+We first evaluate uncorrected estimators based on raw model outputs: (i) single-model hard-label predictions from the strongest models (Llama-3.3-70B, Llama-4-Scout, Qwen-2.5, gpt-oss-120B, GPT-5), (ii) a majority-vote ensemble of the best open models (Llama-3.3-70B, Qwen-2.5, gpt-oss-120B), and (iii) soft-label distributions obtained from these three open models’ outputs.
+We then evaluate debiased estimators obtained by applying DSL to (iv) a single strong model (Llama-3.3-70B) and (v) the soft-label ensemble.
+For each estimator, we compute temporal trends by grouping instances by decade and computing the share of instances assigned to each label in each decade.
+For DSL-based approaches, we use the bias-corrected pseudo-outcomes $\tilde{Y}_{i}$ directly in these aggregations.
+We then quantify agreement with the reference model by computing Pearson and Spearman correlations, as well as root mean squared error (RMSE), between model-based and human-based time series.
+We compute correlations separately for each label and report the average across labels.
+Table˜3 shows that the soft-label ensemble corrected with DSL recovers human-derived temporal trends more accurately than other approaches, including a single-model DSL estimator, at both the high-level and fine-grained levels.
+Model / Estimator Pearson Spearman RMSE
+High-level trends
+Llama-3.3-70B 0.87 0.86 0.08
+Llama-4-Scout 0.74 0.76 0.08
+Qwen-2.5 0.72 0.71 0.11
+gpt-oss-120B 0.84 0.80 0.07
+GPT-5 0.82 0.82 0.07
+Majority vote ensemble 0.88 0.87 0.08
+DSL (Llama-3.3-70B) 0.85 0.83 0.06
+DSL (soft-label ensemble) 0.89 0.91 0.06
+Fine-grained trends
+Llama-3.3-70B 0.64 0.58 0.09
+Llama-4-Scout 0.58 0.58 0.07
+Qwen-2.5 0.36 0.38 0.11
+gpt-oss-120B 0.44 0.52 0.10
+GPT-5 0.56 0.50 0.08
+Majority vote ensemble 0.64 0.67 0.10
+DSL (Llama-3.3-70B) 0.50 0.49 0.07
+DSL (soft-label ensemble) 0.68 0.67 0.04
+Table 3: Comparison with human-based decade-level trends across estimators. The table reports average Pearson and Spearman correlations and RMSE (reference: human majority vote).
+At the high level, the strongest single models and ensemble variants already achieve high agreement with human trends, with Pearson and Spearman correlations above 0.80 and low RMSE.
+In this setting, DSL shows moderate but consistent improvements: DSL applied to the soft-label ensemble achieves the highest correlations (Pearson 0.89; Spearman 0.91) and the lowest RMSE (0.06).
+(a) Human majority vote
+(b) gpt-oss-120B
+(c) DSL (soft labels)
+Figure 8: High-level (anti-)solidarity trends for the Migrant category under different labeling and correction strategies based on the 2k evaluation subset.
+Trends are computed as label shares within each decade (see definitions (a) and (c) in Section˜3.2.2).
+The grey shaded area marks the period 1930–1949, excluded due to limited data availability during and immediately after the NS dictatorship.
+Figure˜23 in the Appendix shows the same trends for the remaining configurations.
+In contrast, fine-grained trends are substantially harder to approximate from raw model predictions alone.
+Single-model and ensemble estimators show lower correlations and higher error: notably, even models that achieve the highest macro F1 scores in Section˜3.1.1 show only moderate agreement with human-derived fine-grained trends (for example, gpt-oss-120B reaches a Pearson correlation of 0.44 with an RMSE of 0.10, and the majority-vote ensemble reaches 0.51 with an RMSE of 0.08).
+This indicates that strong instance-level performance does not guarantee accurate recovery of fine-grained temporal trends.
+Here, DSL has a clear effect, particularly in reducing error: DSL applied to soft-label ensembles achieves the lowest RMSE (0.04) and higher correlations (Pearson 0.68) than all uncorrected approaches.
+(a) Human majority vote
+(b) gpt-oss-120B
+(c) DSL (soft labels)
+Figure 9: Fine-grained (anti)solidarity trends for the Migrant category under different labeling and correction strategies based on the 2k evaluation subset.
+Trends are computed as label shares within each decade (see definitions (a) and (c) in Section˜3.2.2).
+The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
+Figure˜24 in the Appendix shows the same trends for the remaining configurations.
+Figure˜8 and Figure˜9 illustrate these results.
+First, DSL reduces overuse of the None category and correspondingly increases estimated trends for solidarity and anti-solidarity, for example for Llama-3.3-70B.
+This pattern is also reflected in Table˜8 in the Appendix, which compares per-label correlations with human-annotated decade-level trends for Llama-3.3-70B and soft-label DSL.
+There, DSL reduces the RMSE for None from 0.13 to 0.07.
+Box 3.1: Label Shares Over Time Each speech belongs to a time period $t$ (e.g.
+a decade or a year) with $N_{t}$ speeches and receives a high-level or fine-grained label $\ell$.
+The estimated label share $\hat{p}_{t,\ell}$ in period $t$ is defined as follows: (a) Hard-label trends (e.g.
+human majority vote, gpt-oss-120B): $\hat{p}_{t,\ell}=\frac{N_{t,\ell}}{N_{t}}$, where $N_{t,\ell}$ denotes the number of speeches in period $t$ assigned label $\ell$.
+(b) Soft-label trends [^6] Reported in Figure˜24 in the Appendix.: $\hat{p}^{\text{soft}}_{t,\ell}=\frac{1}{N_{t}}\sum_{i\in t}Y_{i,\ell}$, where $Y_{i,\ell}$ denotes the probability of label $\ell$ for speech $i$, obtained from the distribution of annotator or model labels.
+(c) DSL (soft-label) trends: $\hat{p}^{DSL}_{t,\ell}=\frac{1}{N_{t}}\sum_{i\in t}\tilde{Y}_{i,\ell}$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
+At the fine-grained level, raw model outputs, including gpt-oss-120B, Llama-3.3-70B, Qwen-2.5, and their ensemble variants, show unstable trends, with dominance shifting between group-based, exchange-based, and compassionate categories across decades.
+DSL achieves more stable trends that follows with the human reference more closely, particularly within solidarity.
+Most notably, all raw model trends overemphasize group-based solidarity at the expense of compassionate solidarity.
+DSL corrects this imbalance and restores the dominance of compassionate solidarity observed in human annotations.
+Table˜8 in the Appendix shows that these corrections correspond to substantial error reductions, with RMSE for group-based solidarity decreasing from 0.20 to 0.06 and for compassionate solidarity from 0.20 to 0.08.
+A similar pattern is observed for fine-grained anti-solidarity.
+In raw model outputs, anti-solidarity is often limited to a single dominant subtype or disappears entirely; for example, empathic and compassionate anti-solidarity are absent from trends derived from Llama-3.3-70B predictions.
+DSL restores these subtypes and recovers differentiation between group-based and exchange-based anti-solidarity.
+In sum, these results suggest that combining soft-label model outputs with DSL is a viable strategy for studying long-run trends when full human annotation is not feasible.
 4 Temporal Trends in Solidarity Towards Migrants
-In this section, we apply the best-performing open-weight model identified in 3, Llama-3.3-70B, to the full Migrant dataset of 63k instances to analyze trends across time.
-We focus on the Migrant dataset in particular because migration has long been one of the most contested issues in German politics, consistently framed through ?crisis narratives?
-that have shaped national identity and migration governance across decades frohlich2025migration (36).
-This makes it a compelling case for evaluating whether model-based annotations align with historical and political science literature.
-It is also reflected in our human-annotated dataset, where migrant-related instances display a more diverse label distribution, including substantially more anti-solidarity, compared to the predominantly solidarity-oriented discourse on women.
-More specifically, 4.1 compares the long-term (anti-)solidarity trends across human annotators and LLMs and thereby further validates the results of our models.
-In doing so, we also note two periods with notable shifts in solidarity, namely the post-war period and the recent past, which we cover in more detail in Section 4.2 and Section 4.3.
-(a) GPT-4
-(b) Llama-3.3-70B
-Figure 7: Solidarity trends for Migrant predicted by GPT-4 and Llama-3.3-70B.
-Values are normalized within each decade; percentages may not total 100% because instances classified as ?None?
-are not shown (though included in the calculation).
-The period 1933–1949 (grey shaded area) is excluded due to limited data availability during and immediately after the Nazi dictatorship.
-4.1 Trends with GPT-4 vs with Llama-3.3-70B
-(a) GPT-4
-(b) Llama-3.3-70B
-Figure 8: (Anti-)solidarity frames trends for Migrant predicted by GPT-4 and Llama-3.3-70B.
-Percentages are normalized within each decade across all eight subtypes; thus, the curves in both panels together sum to 100% per decade.
-The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
-Figures 7 and 8 show the development of (anti-)solidarity for GPT-4 and Llama-3.3-70B on a high and fine-grained level across 155 years.
-The GPT-4 reference data is taken from kostikova-fine (57), based on a subset of 18.3k migrant-related instances (from an original 58k-instance dataset; see 2.1) spanning 1867–2022.
-While Llama-3.3-70B is applied to the full dataset (expanded to 63k instances through 2025) for the main analysis, we restrict it to the same 18.3k-instance subset for this comparison.
-To validate the annotations of the models, we first compare them to each other.
-Comparing the trends between GPT-4 and Llama-3.3-70B, we observe a high correlation of 0.96 on the high level and between 0.77 and 0.95 across the more frequent fine-grained subtypes (see Table 4 in the Appendix).
-This indicates that the models show a high degree of agreement in temporal trends.
-[^5] Additionally, Llama-3.3-70B shows near-identical trends between the 18.3k subset and the full 63k dataset (e.g., $\rho$ = 0.996 for solidarity and anti-solidarity, and $\sim$0.97–0.98 for subtypes).
-For reference, Figure 3 shows the high-level trends across time for human annotations ($N=2,084$).
-We observe similar relative distributions of (anti-)solidarity on a high level.
-Direct trend correlations between Llama-3.3-70B and human annotations are also substantial ($\rho$ = 0.84 for anti-solidarity, $\rho$ = 0.76 for solidarity, $\rho$ = 0.47 for mixed).
-In terms of Cohen’s $\kappa$ (computed on the 2,084 human annotated instances), GPT-4 slightly outperforms Llama-3.3-70B ($\kappa$ = 0.57 vs.
-0.50 at the high level; 0.38 vs.
-0.36 for subtypes), but differences are modest.
-Overall, GPT-4 and Llama-3.3-70B appear to produce similar results that agree to a moderate level with human annotations.
-Given that it is an open weight model that we can run on our own computing hardware at low further cost, we utilize Llama-3.3-70B for the subsequent analysis.
-4.1.1 Trend Stability Across Keywords.
-Before we move on to interpreting the observed trends derived from Llama-3.3-70B predictions, we verify that they are not artifacts of our keyword-based data collection (see 2.1).
-As Figures 19 and 20 in the Appendix show, individual keywords vary considerably in their temporal distribution and in the balance of solidarity vs.
-anti-solidarity.
-To rule out the possibility that our long-term trends merely reflect such keyword effects, we perform two robustness tests.
-First, we conduct a stability test by randomly sampling 200 subsets of at least 5 keywords each and comparing the resulting high-level trends using Pearson’s correlation coefficient.
-Across all 19,900 pairwise comparisons, the trends are highly consistent (average correlation of 0.92 for solidarity, 0.81 for anti-solidarity, and 0.94 for mixed; with 25% quantiles of 0.90, 0.70, and 0.92, respectively).
-[^6] For more detail on keywords and the stability test, see A.
-Second, we restrict the analysis to the two most frequent and temporally widespread keywords, Ausländer (en.: foreigners) and Flüchtlinge (en.: refugees), which occur in every decade (19.7k and 12k mentions, respectively).
-Together, the resulting plots (Figure 23 in the Appendix) and the stability test results confirm that overall trends in solidarity and anti-solidarity are preserved both under this restricted vocabulary and across different keyword selections.
-4.1.2 Notable Trends in 1867-2022.
-Inspecting the trends in Figure 7, we observe four periods with notable shifts in solidarity: Between 1890-1920, in the 1930s, in the post-war period, and in the recent past after 2010.
-Regarding the 1890-1920 period, we observe some disagreement between the models: GPT-4 shows an increase in solidarity between 1900 and 1920, while Llama-3.3-70B does not.
-We also note that the agreement between human annotators and between models and annotators is lower for historic data, such that we do not consider this trend to be sufficiently reliable for further analysis.
-The 1930s and 1940s data points are outliers, as the parliament was swiftly sidelined by the Nazi regime and the only 109 available samples from 1933 to 1949 mostly concern inter-state relations (mentioning, e.g., ?foreign creditors?) that fall into the ?None?
-category.
-Hence, we exclude these data points from further analysis, as well.
+In this section, we analyze temporal trends in the full Migrant dataset (63k instances) using DSL-corrected soft labels.
+We relate the resulting patterns to findings from political science, sociology, and history on migration to assess whether the temporal trends recovered by the DSL-corrected estimates are consistent with established interpretations.
+For these analyses, DSL uses the decade-stratified human annotation sample of 2,184 Migrant instances (3.5% of the full corpus) to estimate bias-corrected label distributions for the full corpus of 63k speeches.
+Since human annotations were obtained under a decade-stratified sampling design (see Section˜2.2), we estimate DSL with decade covariates and decade-specific labeling probabilities $\pi_{i}$, and reuse the resulting pseudo-outcomes $\tilde{Y}_{i}$ for all analyses that condition only on time, including the sub-period analyses introduced in Section˜4.1.2.
+For analyses that condition on additional variables, such as party or keyword, we re-estimate DSL on the relevant analysis sample, since the prediction function $\hat{g}(Q_{i},X_{i})$ changes when additional covariates are included.
+In these analyses, party or keyword indicators enter the DSL prediction step as additional covariates alongside decade indicators, while labeling probabilities $\pi_{i}$ remain defined at the decade level, consistent with the annotation design.
+We first present the DSL-corrected trends for the full dataset, summarizing the broad long-term pattern in parliamentary discourse about migrants and briefly assessing whether the aggregate trends are robust to variation in keyword composition.
+We then identify two periods with particularly notable shifts in solidarity–the post-war period and the recent past–which we analyze in more detail in Section˜4.2 and Section˜4.3.
+4.1 Overall Solidarity Trends with DSL, 1867-2025
+Figure 10: Solidarity trends for Migrant speeches computed using DSL-corrected soft labels (see definition (c) in Section˜3.2.2) on the full 63K dataset.
+Figure˜10(a) shows the high-level (anti-)solidarity trends, while Figure˜10(b) shows distribution of fine-grained (anti-)solidarity subtypes over time.
+The period 1930–1949 (grey area) is excluded due to limited data during and after the Nazi dictatorship.
+The overall trends from 1867 to 2025 are shown in Figure˜10.
+Across most of the observable period, and especially from the beginning of the 20th century through the 2010s, solidarity exceeds anti-solidarity in parliamentary discourse about migrants.
+This pattern is driven mainly by compassionate solidarity, with group-based solidarity playing a less prominent but persistent role.
+Solidarity peaks in the postwar decades, while anti-solidarity remains comparatively low until the late twentieth century.
+From the 1980s onward, anti-solidarity rises (especially its group-based, compassionate, and exchange-based subtypes), with the sharpest jump occurring between the 2010s and 2020s.
+Before turning to specific historical periods, we assess whether these aggregate trends are robust to changes in keyword composition over time.
+4.1.1 Robustness to Keyword Composition
+As Figure˜21 and Figure˜22 in the Appendix show, individual keywords vary widely in temporal distribution in the balance of solidarity vs.
+anti-solidarity associated with them.
+To ensure our long-term trends are not driven by such effects, we conduct two robustness tests.
+First, we conduct a stability test by randomly sampling 200 subsets of at least five keyword groups (grouping inflectional variants).
+For each subset, we re-estimate the high-level DSL trends (as defined in Section˜3.2.2) and compare the resulting time series to the full-sample trend using Pearson’s correlation coefficient.
+The trends remain broadly stable, with mean correlations of 0.93 for solidarity, 0.88 for anti-solidarity, and 0.82 for mixed; for more detail on keywords and the stability test, see Appendix˜A in the Appendix.
+Second, we restrict the corpus to the two most frequent and temporally widespread terms, Ausländer (en.: foreigners) (19.7k mentions) and Flüchtlinge (en.: refugees) (12k), and re-estimate DSL on this reduced sample.
+Both checks confirm that overall solidarity and anti-solidarity trends are preserved both under this restricted vocabulary (see Figure˜33) and across different keyword selections.
+4.1.2 Notable Trends in 1867-2025
+Inspecting Figure˜10, we identify three periods with notable shifts in solidarity: 1890-1900s, the post-war period, and the recent period after 2010.
+We note that the 1930s and 1940s are outliers, as the parliament was swiftly sidelined by the Nazi regime and the corpus contains only 109 available instances from 1933 to 1949.
+Hence, we exclude these data points from further analysis.
+The first of these, 1890-1900, is difficult to interpret with confidence.
+In this early part of the series, even the strongest model in our ensemble performs worse than in later decades in terms of macro-F1, and agreement between human annotators is also lower (see Figure˜18).
+Although DSL corrects the resulting estimates, we do not consider this early trend sufficiently reliable for detailed substantive analysis.
+This period also contains fewer instances than the post-war and recent periods that we analyze in detail below (see Figure˜2).
 For the two remaining periods, 1949–1957 and 2009–2025, we take a more detailed look.
-In the postwar years (Section 4.2), parliamentary discourse reveals high levels of solidarity and little anti-solidarity toward migrants, despite the severe economic and infrastructural strain of the time braun2014immigration (14).
-This raises our first guiding question: Why is there so much solidarity (and so little anti-solidarity), and what was it based on?
-In contrast, the recent period (Section 4.3) shows a clear rise in anti-solidarity and a decline in solidarity, beginning in the early 2010s.
-These shifts likely reflect the political impact of recent refugee movements and the growing influence of anti-immigration views, such as those promoted by the far-right party Alternative für Deutschland (AfD) (en.: Alternative for Germany), which has gained substantial electoral support in Germany since its founding in 2013.
-This leads to our second question: What are the drivers of anti-solidarity in the contemporary period?
 4.2 Germany after the Second World War, 1949-1957
-The Second World War set off the largest wave of forced migrations in European history, with tens of millions of people, both non-Germans and ethnic Germans, displaced or resettled across Central and Eastern Europe schulze2017commemoration (90).
-The populations displaced by the war were highly heterogeneous, including refugees, displaced persons (DPs), forced laborers, and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans.
-For those who settled in postwar Germany, integration posed significant political and societal challenges, shaped by regional disparities, economic hardship, and tensions with the local population chevalier2024forced (21).
-This subsection focuses on the early postwar period (1949–1957), when the Federal Republic of Germany institutionalized solidarity as part of nation-building.
-This period is analytically important because solidarity was no longer merely a moral gesture, but a tool of governance, embedded in legal frameworks like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy kittel2020stiefkinder (54).
-Crucially, such measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups from the benefits provided by these laws.
-However, since solidarity in this period was neither static nor uniform, shaped by diverse migrant groups, competing political ideologies, and contested notions of deservingness ther1996integration (96), we draw on Thijssen’s typology of solidarity thijssen2012mechanical (97) to guide our analysis.
-By distinguishing between identity-, need-, and contribution-based forms, this framework allows us to examine how solidarity was constructed and sustained amid institutional strain.
+The Second World War triggered the largest wave of forced migration in European history, with tens of millions of people displaced or resettled across Central and Eastern Europe (Wessel, 2017).
+The affected populations were heterogeneous—refugees, displaced persons (DPs), forced laborers—and—forming the largest group of expellees after 1945 in both Europe as a whole and Germany—ethnic Germans.
+For those who settled in postwar Germany, integration posed political and societal challenges shaped by regional disparities, economic hardship, and tensions with locals (Chevalier et al., 2024).
+This subsection examines the early postwar period (1949-1957), when the Federal Republic institutionalized solidarity as part of nation-building.
+This period is analytically important because solidarity became a tool of governance, embedded in laws like the Lastenausgleich (LAG) (en.: Equalization of Burdens), which redistributed resources between the war-damaged and the less affected to promote social cohesion and democratic legitimacy (Kittel, 2020).
+These measures applied specifically to Volkszugehörige (en.: ethnic Germans), excluding other migrant groups.
+Since solidarity was neither static nor uniform, shaped by diverse migrant groups, competing ideologies, and contested notions of deservingness (Ther, 1996), we employ Thijssen’s typology of solidarity (Thijssen, 2012) to guide our analysis.
+By distinguishing identity-, need-, and contribution-based forms, we analyze how solidarity was constructed and sustained amid institutional strain.
 4.2.1 Overall solidarity distribution
-Throughout the 1949–1957 period, solidarity dominated parliamentary discourse on postwar population movements.
-As shown in Figure 9, [^7] We use raw counts as they provide a more interpretable view of stance distribution during this period, given the overwhelming dominance of solidarity and the low frequency of categories like anti-solidarity.
-mentions of solidarity consistently outpaced both anti-solidarity and mixed statements, at times reaching ten times their frequency.
-Figure 9: Distribution of (Anti-)solidarity mentions in Bundestag debates, 1949–1957 (raw counts).
-LAG refers to the Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG to the Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG to the Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
-Mentions of solidarity closely track key legislative moments, particularly in the debates leading up to the passage of major laws.
+During 1949–1957, solidarity dominated parliamentary discourse on postwar population movements.
+As Figure˜11 shows, [^7] We report (DSL-adjusted) counts rather than proportions (see definition (d) in Section 4.2.1) because counts make spikes visible whenever migration-related laws are debated and enable comparison of debate volume across months.
+mentions of solidarity consistently exceeded anti-solidarity and mixed statements, at times by a factor of ten.
+Figure 11: Estimated counts of (anti-)solidarity mentions in Bundestag debates, 1949–1957 (DSL-adjusted, see definition (d) in Section˜4.2.1).
+LAG == Lastenausgleichsgesetz (en.: Equalisation of Burdens Law); BVFG == Bundesvertriebenengesetz (en.: Federal Expellees Act); HAuslG == Gesetz über die Rechtsstellung heimatloser Ausländer im Bundesgebiet (en.: Law on the Legal Status of Stateless Foreigners in the Federal Territory).
+Mentions of solidarity closely track key legislative moments.
 A major spike in March 1950 aligns with early discussions on the integration of expellees and Article 131, which addressed the rehabilitation of former public servants of the Nazi regime, many of them displaced.
-A second cluster from January to May 1952 reflects debates on the LAG, preceding its adoption in May (see an excerpt from such a debate in Example 1 in Table 10 in the Appendix).
-Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified rights and integration measures, see Example 2 for such a debate.
-Later peaks in September 1954 and June 1955 reflect continued debate over LAG implementation and budget issues.
+A second cluster in early 1952 reflects debates on the LAG, preceding its adoption in May (see Example 15 in Table˜19 in the Appendix).
+Another sharp rise in February–March 1953 corresponds to readings of the Bundesvertriebenengesetz (BVFG) (en.: Federal Expellee Law), which defined the legal status of expellees and refugees and codified integration measures (Example 15 ).
+Later peaks in September 1954 and June 1955 reflect debate over LAG implementation and budgets.
 Smaller but sustained increases in 1956 relate to new support measures for Sowjetzonenflüchtlinge (en.: Soviet-zone refugees) and Aussiedler (en.: resettlers).
-Figure 10: Distribution of (Anti-)solidarity frame subtypes from 1949 to 1957 (raw counts).
-An inset plot shows the evolving composition of solidarity subtypes, normalized within each 6-month period.
-Values represent the proportional share of each subtype among all solidarity mentions during that period.
-The prominence of solidarity in this period is unsurprising.
-Parliamentary debates used it not only to justify compensation, but also to navigate postwar fragmentation and competing claims to victimhood and belonging hughes1999shouldering (49).
-This multifaceted nature of solidarity is evident in the distribution of all four subtypes across key legislative moments, as shown in Figure 10.
-Group-based and compassionate frames were especially prominent, reflecting an emphasis on both shared national belonging and moral obligation toward displaced populations (see, e.g., Example 1 in Table 10, where solidarity is tied to the right to the homeland and framed as a legal entitlement).
-More striking is the shift in the balance between these two subtypes over time.
-As shown in the inset of Figure 10, group-based solidarity dominated in the early 1950s but declined after mid-1953, while compassionate steadily increased.
-This suggests a rhetorical transition: once legal belonging was formalized, most notably with the LAG in 1952 and BVFG in 1953, solidarity no longer needed to be framed primarily in terms of shared identity.
-The shift may also reflect demographic change within the displaced populations, as newer arrivals like Sowjetzonenflüchtlinge from the GDR and Aussiedler (en.: resettlers) from the Eastern Bloc countries were less easily framed as co-nationals boke1996fluchtlinge (13).
-This trend aligns with Thijssen’s reading of Durkheim: as group-based solidarity extends to more diverse groups, it gives way to need-based appeals thijssen2012mechanical (97) (see Example 3 in Table 10 in the Appendix).
-These patterns show that solidarity was used in parliamentary debates to manage the challenges of postwar reconstruction, as reflected in key legislative reforms.
-While the claimant groups of the broadest measure LAG were highly diverse (bombed out, evacuated, expelled, financially damaged), their eligibility for compensation under the law was based on Volkszugehörigkeit (en.: German ethnicity), excluding many from formal recognition kittel2020stiefkinder (54, 73).
-The concept of Volksdeutsche was a fluid and constructed Nazi category whose ambiguous definition nonetheless had concrete and deadly implications bergen2005tenuousness (9).
-We therefore examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in parliamentary discourse in that period.
+Box 4.1: Estimated Monthly Counts of (Anti-)Solidarity Speeches are grouped by month $t$, with $N_{t}$ speeches in month $t$.
+We define the estimated number of mentions of label $\ell$ in month $t$ as: (d) DSL-adjusted monthly counts: $\hat{C}^{DSL}_{t,\ell}=\sum_{i=1}^{N_{t}}\tilde{Y}_{i,\ell}$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
+Figure 12: Distribution of (anti-)solidarity subtypes, 1949–1957 (DSL-adjusted expected counts, see definition (d) in Section˜4.2.1).
+Inset shows the evolving composition of solidarity subtypes, normalized within each six-month period as shares of all solidarity framing (see definition (c) in Section˜3.2.2).
+The prominence of solidarity in this period can be explained by the fact that debates used it not only to justify compensation but also to manage postwar social and economic fragmentation and competing claims to belonging (Hughes, 1999).
+The distribution of all four subtypes across key legislative moments (Figure˜12) reflects this mix of solidarity rationales.
+Compassionate and group-based frames dominate throughout the period (see inset of Figure˜12).
+By emphasizing shared national belonging (group-based) and the suffering of affected populations (compassionate), these frames construct a victim-centered Schicksalsgemeinschaft (en.: community of fate), uniting Germans through shared loss (Gengler, 2020) (see Example 15in Table˜19 in the Appendix [^8] Example labels are based on the dominant DSL subtype, that is, the subtype with the highest DSL-adjusted score.
+This is useful for illustrating which subtype is most strongly reflected in a given speech within the DSL-based analysis; however, since DSL is designed for aggregate inference, these labels do not guarantee corrected true labels for individual instances.).
+This coexistence of the two subtypes is consistent with Thijssen (2012), who argues that compassionate and group-based solidarity form a dialectical pair: identification with a collective (?we?) generates obligations toward its members, while emphasizing their suffering reinforces that collective bond.
+Additionally, a small but persistent share of exchange-based solidarity appears throughout the period.
+It reflects debates over efficient allocation and the reintegration of displaced populations into labor and housing markets, with measures such as the LAG framed as tools to rebuild productive capacity (Examples 15 and 18).
+These patterns show that solidarity in parliamentary debates mediated the challenges of postwar reconstruction through key legislative reforms.
+However, since these reforms differentiated between migrant groups, we examine how (anti-)solidarity subtypes were expressed across key migrant-related terms in this period.
 4.2.2 Solidarity Frames Across Keywords in 1949-1957
-This analysis includes only migrant-related keywords that appear at least 200 times in combination with a solidarity, anti-solidarity subtype or mixed stance.
-Keywords dominated by none or with low total counts are excluded (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants), etc.); see Table 5 for full distributions.
-Figure 11 shows the distribution of (anti-)solidarity frames across migrant-related keywords in 1949–1957.
-Across all groups, solidarity was far more common than anti-solidarity, with group-based and compassionate frames dominating.
-However, the distribution of subtypes varies between groups.
-Vertriebene (en.: expellees) are most often associated with group-based solidarity, accounting for around half of all expressions with a stance, while also receiving a substantial share of compassionate solidarity (nearly 30%).
-This reflects their dual position: as key representatives of German victimhood (alongside prisoners of war) during and after the World War II moeller2001war (71), they were both integrated into West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice.
-Still, around 5–6% of stance expressions toward Heimatvertriebene fall into the anti-solidarity category, particularly group-based (solidarity limited to ?us?, the in-group).
-These expressed concerns that expellees strained local infrastructures and were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment kossert2008kalte (56).
-Perceptions of injustice were heightened by the uneven distribution of expellees across West German states braun2014employment (15), particularly in regions where their prominent presence in public institutions and politics sparked fears of Überfremdung (en.: overforeignization) and contributed to a sense of unequal burden-sharing levy2005memories (64) see Example 4 in Table 10.
-Figure 11: Relative distribution of (Anti-)Solidarity subtypes across selected migrant-related keywords, 1949–1957.
-Bars represent the proportion of each subtype within each keyword group, thus all subtypes for a given group sum to 100%.
-Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are primarily framed through compassionate solidarity (45–52%), followed by group-based solidarity (25–30%).
-The label Flüchtlinge initially functioned as a broad catch-all category in the late 1940s nachum2018semantics (74).
-With the 1953 BVFG, it was more clearly differentiated into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler.
-Yet their Germanness was not valued equally: while Vertriebene were framed as victims of expulsion and central to West German nation-building, the label Flüchtlinge retained post-war connotations of marginality and foreignness nachum2018semantics (74).
-This helps explain a higher share of compassionate solidarity, as well as the notable anti-solidarity (around 10%).
-Sowjetzonenflüchtlinge, in turn, were primarily framed as Cold War refugees fleeing communist oppression and thus not integrated into the nation building narrative like the Vertriebene, prompting more compassionate than group-based solidarity nachum2018semantics (74, 13) (see Example 5).
-Ausländer (en.: foreigners) were the main targets of anti-solidarity.
-Not a single coherent group, they were a discursively constructed ?other?, including DPs, Eastern European border crossers, wartime laborers, and early foreign workers smith1994germany (93).
-Often contrasted with German groups like Vertriebene or Einheimische, they were framed as non-belonging, economically burdensome, and sometimes criminalized holian1945missing (44) (see Example 6).
-However, despite this exclusionary framing, solidarity with Ausländer, especially group-based, exceeded anti-solidarity.
-This might be partly due to the category’s fluidity, which included ethnic Germans not yet formally recognized as such ruhkopf2023institutionalisierte (86) in the early postwar years.
-Politicians like Lukaschek also linked DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community ruhkopf2023institutionalisierte (86) (Example 7).
-The second-highest solidarity subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy within the Western alliance ruhkopf2023institutionalisierte (86) (Example 8).
-The 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure, granted DPs civic and social rights largely to meet conditions for sovereignty and improve the country’s international standing holian1945missing (44, 86).
-At the same time, instances of genuine solidarity toward non-ethnic German foreigners, particularly in CDU and SPD, statements, show that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 7 and 9).
-In sum, solidarity was expressed differently depending on the migrant group and was shaped by perceived ?deservingness?.
-Vertriebene were largely included as part of the national community, while others, especially Ausländer, were treated as outsiders and received comparatively less support.
-These differences reflected not only group characteristics but also the ideological strategies of political parties.
+The analysis includes migrant-related keywords for which the DSL-adjusted expected number of stanced instances (i.e., solidarity, anti-solidarity, or mixed rather than none) is at least 150.
+Keywords dominated by none or with low DSL-adjusted expected counts (e.g., Asylsuchende (en.: asylum seekers), Einwanderer (en.: immigrants), Emigranten (en.: emigrants)) are excluded; see Table˜9 in the Appendix for full distributions.
+Figure˜13 shows the distribution of (anti-)solidarity frames across these terms.
+Vertriebene (en.: expellees), a category reserved primarily for Volkszugehörige (en.: ethnic Germans), are most often associated with compassionate solidarity (about half of all expressions with a stance), and also with group-based solidarity (nearly 25%).
+This reflects their dual position: as key representatives of German victimhood during and after the World War II (Moeller, 2001), they were both central to West Germany’s nation-building narrative and portrayed as victims of Allied or foreign injustice.
+Box 4.2: Distribution of (Anti-)Solidarity Subtypes Across Groups Speeches are grouped by a categorical variable $g$ (a keyword or a political party).
+$I_{g}$ is the set of speeches belonging to group $g$.
+The share of subtype $\ell$ within group $g$ is defined as: (e) Subtype shares within groups: $\hat{s}_{g,\ell}=\frac{\hat{M}_{g,\ell}}{\hat{M}_{g}}$, where $\hat{M}_{g,\ell}=\sum_{i\in I_{g}}\tilde{Y}_{i,\ell}$ is the DSL-adjusted mass of subtype $\ell$ in group $g$, where $\tilde{Y}_{i,\ell}$ denotes the DSL-adjusted class score for label $\ell$ in speech $i$, as defined in Equation˜1 in Section˜2.
+$\hat{M}_{g}=\sum_{\ell\in L}\hat{M}_{g,\ell}$ is the total mass of group $g$ across all labels $L$ (four solidarity subtypes, four anti-solidarity subtypes, and mixed stance).
+Shares sum to one within each group $g$.
+Still, around 10% of stance expressions toward Heimatvertriebene were anti-solidarity, particularly group-based (solidarity limited to ?us?, the in-group), reflecting concerns that expellees strained local infrastructures or were unfairly privileged over equally suffering Einheimische (en.: locals), mirroring broader public resentment (Kossert, 2008).
+Perceptions of injustice grew with the uneven distribution of expellees across West German states (Braun and Mahmoud, 2014), especially where their prominence in institutions and politics sparked fears of Überfremdung (en.: overforeignization) and unequal burden-sharing (Levy and Sznaider, 2005) (see Example 15 in Table˜19).
+Figure 13: Relative distribution of (anti-)solidarity subtypes across selected migrant-related keywords, 1949–1957 (DSL-adjusted estimates; see definition (e) in Section˜4.2.2).
+Bars show each subtype’s share within a keyword group; shares sum to 100%.
+Full label distributions appear in Table˜9 in the Appendix.
+Flüchtlinge (en.: refugees), Sowjetzonenflüchtlinge (en.: Soviet-zone refugees), and Aussiedler (en.: resettlers) are also primarily framed through compassionate solidarity (50–72%), followed by group-based solidarity (15%).
+The label Flüchtlinge began as a broad catch-all in the late 1940s but, even after the 1953 BVFG differentiated it into Vertriebene, Sowjetzonenflüchtlinge, and Aussiedler, it retained connotations of marginality and foreignness (Nachum and Schaefer, 2018).
+This helps explain a more notable share of anti-solidarity (around 10%).
+Sowjetzonenflüchtlinge, framed as Cold War refugees from communist oppression, were less integrated into the nation-building narrative than Vertriebene, prompting more compassionate than group-based solidarity (Nachum and Schaefer, 2018; Böke, 1996) (see Example 16).
+Among the migrant categories, Ausländer (en.: foreigners) shows one of the highest relative shares of anti-solidarity.
+A discursively constructed ?other?, including DPs [^9] With the 1951 HAuslG, authority over the remaining DPs was transferred from Allied administration to the Federal Republic, and they were reclassified as heimatlose Ausländer (en.: stateless foreigners), a noncitizen status that granted them most civil rights without granting citizenship (Holian, 2011)., Eastern European border crossers, wartime laborers, and early foreign workers (Smith, 1994), they were contrasted with groups like Vertriebene or Einheimische and framed as non-belonging (evident in group-based anti-solidarity share), economically burdensome (exchange-based anti-solidarity), or even criminalized (Holian and Wilhelm, 2017) (see Example 16).
+Yet solidarity, especially group-based, still exceeded anti-solidarity, partly due to the category’s fluidity: it sometimes included ethnic Germans not yet formally recognized as such in the early postwar years (Ruhkopf, 2023).
+Speakers reinforced this dynamic by linking DPs and expellees through shared narratives of lost Heimat (en.: homeland), enabling symbolic inclusion into the national victim community (Ruhkopf, 2023) (Example 16).
+The second-highest subtype, exchange-based solidarity, reflects the perceived political utility of Ausländer in strengthening West Germany’s international legitimacy (Ruhkopf, 2023), for example through the 1951 Homeless Foreigners Law (HAuslG), passed under Allied pressure (Holian and Wilhelm, 2017) (Example 17).
+However, solidarity also extended to non-ethnic German foreigners, particularly in CDU and SPD statements, showing that solidarity with Ausländer was not limited to Germanness or foreign policy utility but also reflected broader humanitarian commitments (see Examples 16 and 17).
+In sum, compassionate framing dominated across most categories, reflecting a broader discourse of shared suffering in which displaced populations were positioned as victims of the war and its consequences.
+Nevertheless, solidarity was still shaped by perceived ?deservingness?
+and ethnic identity.
+While Vertriebene were more firmly incorporated into the national community as ethnic Germans, others, especially Ausländer, received comparatively less support and conditional inclusion (Nachum, 2013).
+These differences reflected both group characteristics and the ideological strategies of political parties.
 The next subsection examines how different parties in the early Bundestag framed (anti-)solidarity in their parliamentary speech.
 4.2.3 Distribution across parties
-Figure 12 shows the distribution of solidarity and anti-solidarity subtypes across parties for the period 1949–1957.
-While solidarity was the dominant stance overall, parties varied in how they framed it, driven by ideological and strategic choices.
-Some parties used solidarity to promote national integration, while others employed it to defend regional concerns.
-Broadly, we divide parties into three categories: mainstream governing parties, expellee advocacy parties, and smaller parties with ideological or regional agendas.
-The main parties – governing CDU/CSU and FDP, and opposition SPD – supported the integration of expellees and refugees through legal and institutional measures, primarily expressing group-based or compassionate solidarity (see Figure 12).
-The CDU/CSU played a leading role in proposing the LAG, which framed solidarity as a national obligation toward fellow Germans, and even included expellee representatives in the party ranks hughes1999shouldering (49).
-The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers connor2018refugees (22), resulting in occasional anti-solidarity statements.
-The FDP, while formally aligned with CDU/CSU and SPD on key refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation connor2018refugees (22) (see Example 10 in Table 10, where FDP speakers argued that expellees should build their own homes through self-help and cooperative effort, with state support limited to loans and low-interest credit).
-Expellee advocacy parties (GB/BHE and DP) [^8] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights) and Deutsche Partei (DP) (en.: German Party) explicitly positioned themselves as political representatives of displaced Germans.
-Following multiple successes in state elections, GB/BHE entered the federal government coalition in 1953, thereby exerting significant pressure on other parties to address the demands and concerns of expellees connor2018refugees (22).
-They framed solidarity primarily in terms of historical injustice and economic entitlement, calling for wealth redistribution, fair pensions, and affordable housing.
-Their rhetoric relied heavily on group-based solidarity, portraying expellees as full members of the national community who deserved compensation and full integration (see Example 11).
-Eventually, both major parties of the Federal Republic CDU and SPD managed to absorb the demands and concerns of the expellees, which made GB/BHE and DP less successful connor2018refugees (22, 54).
-Regionalist parties such as the WAV and BP [^9] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized group-based solidarity, but restricted it to ethnic Germans.
-The WAV, a short-lived nationalist party, appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories kranenpohl1998zwischen (59) (Example 12).
-The BP, responding to Bavaria’s disproportionate refugee burden, called for collective responsibility in crisis, fairer redistribution and showed preference for groups like the Sudeten Germans braun2020101330 (16) (see Example 13,where the BP argued that Sudeten and Southeast Germans should be concentrated in Bavaria because they were ?closest to the Bavarian cultural sphere?
-and thus easier to integrate).
-Both parties also expressed the highest levels of anti-solidarity, either by excluding non-German migrants (WAV) connor2018refugees (22) or opposing specific groups such as Silesians (BP).
-Figure 12: (Anti-)solidarity frames distribution across major parties in the Bundestag from 1949 to 1957, by election period.
-Percentages are normalized within each party, such that all subtype bars (solidarity and anti-solidarity) together sum to 100% per party.
-We include only those parties with a sufficient number of stanced instances $(N>100)$.
-The full label distribution across all parties is provided in Figure 6 in the Appendix.
-The KPD [^10] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) framed displaced populations through a lens of socialist integration, emphasizing group-based solidarity to support their incorporation into the new social order schwartz2008vertriebene (91).
-This solidarity, however, was assimilationist rather than inclusive: the party avoided the term Vertriebene, in line with Soviet repression of expellee identity heidemeyer2007deutsche (42), and this was reflected in its minimal use of empathic solidarity.
-The KPD referred almost exclusively to ?refugees?, advocating land reform, compensation, and integration as part of a broader effort to portray them as victims of fascism and capitalism and to gain political support connor2018refugees (22) (see Example 14, where the KPD argued that refugee employment could not be solved under capitalism).
-Some parties active in the early postwar years, such as the KPD, WAV, and BP, did not stay in parliament for long, but they played an important role in shaping political debates of the time.
-Rather than merely reflecting existing social tensions, as suggested by lipset1967cleavage (67), these parties actively contributed to defining what solidarity meant and who was considered deserving of support.
-As suggested by thijssen2022s (98), such framings were shaped by the political and social challenges of the period and reflected broader efforts to maintain or renegotiate social cohesion.
+Figure˜14 shows party distributions of solidarity and anti-solidarity in 1949–1957. Solidarity dominated overall, but parties framed it differently, shaped by ideology and strategy.
+The main parties – governing CDU/CSU and FDP, as well as opposition SPD – supported expellee and refugee integration through legal and institutional measures, primarily expressing compassionate or group-based solidarity (see Figure˜14).
+The CDU/CSU led in proposing the LAG, framing solidarity as a national duty toward fellow ethnic Germans, and even included expellee representatives in the party ranks (Hughes, 1999).
+The SPD also supported the law but voiced concerns about former Nazis and communist agents among the newcomers (Connor, 2018), resulting in occasional anti-solidarity statements.
+The FDP, while formally aligned with CDU/CSU and SPD on refugee issues, framed solidarity in more individualistic, exchange-based terms, opposing large-scale redistribution and favoring market-based solutions consistent with its liberal economic orientation (Connor, 2018) (see Example 17 in Table˜19).
+Figure 14: Distribution of (anti-)solidarity subtypes across Bundestag parties, 1949–1957 (DSL-adjusted estimates; see definition (e) in Section˜4.2.2).
+Bars show each subtype’s share within a party, summing to 100%.
+The analysis includes parties with at least 150 DSL-adjusted stanced statements.
+Full label distributions appear in Table˜10 in the Appendix.
+The expellee advocacy party GB/BHE [^10] Gesamtdeutscher Block/Bund der Heimatvertriebenen und Entrechteten (GB/BHE) (en.: All-German Bloc/League of Expellees and Deprived of Rights).
+The DP (Deutsche Partei (en.: German Party)) also addressed expellee demands, especially in Lower Saxony because of their electoral weight there, but it began as a Heimatpartei (en.: ?homeland party?) rather than a refugee party (Aschoff, 2014).
+This is visible in the party’s share of group-based anti-solidarity, reflected in its defense of local interests (Example 19).
+positioned itself as a political representative of displaced Germans.
+Following multiple successes in state elections, GB/BHE, founded by and for expellees, entered the federal coalition in 1953 and pressured other parties to address expellee demands (Connor, 2018).
+It expressed compassionate and group-based solidarity, portraying expellees as full members of the national community entitled to compensation and integration (see Example 17).
+Additionally, the GB/BHE shows the highest share of empathic solidarity (15%), consistent with its role as an expellee party defining expellees as a distinct community.
+Over time, CDU and SPD managed to absorb major expellee demands, which made GB/BHE less successful (Connor, 2018; Kittel, 2020).
+Regionalist parties such as WAV and BP [^11] Wirtschaftliche Aufbau-Vereinigung (en.: Economic Reconstruction Union) and Bayernpartei (en.: Bavarian Party) emphasized compassionate solidarity, but restricted it to ethnic Germans.
+The short-lived WAV appealed to expellees through anti-communist rhetoric and demands to reclaim eastern territories (Kranenpohl, 1998) (Example 18).
+The BP, responding to Bavaria’s disproportionate refugee burden, called for invoked collective responsibility, fairer redistribution and favored groups like the Sudeten Germans (Braun and Dwenger, 2020), which corresponds to its high share of group-based anti-solidarity (see Example 18).
+Both parties also expressed the highest levels of mixed stance, either by excluding non-German migrants (WAV) (Connor, 2018) or opposing specific groups such as Silesians (BP).
+The KPD [^12] Kommunistische Partei Deutschlands (en.: Communist Party of Germany) also addressed displaced populations, framing them through socialist integration and emphasizing compassionate and group-based solidarity, though in an assimilationist rather than inclusive sense.
+Avoiding the label Vertriebene in line with Soviet policy, the party instead spoke of ?refugees?, portraying them as victims of fascism and capitalism (Heidemeyer, 2007; Connor, 2018).
+Parties like KPD, WAV, and BP did not stay in parliament for long, but nonetheless shaped early postwar debatesand helped define solidarity and its boundaries of deservingness.
+As suggested by Thijssen and Verheyen (2022), such framings were conditioned by the period’s political and social challenges and reflected broader efforts to maintain or renegotiate societal cohesion.
 4.3 Recent German history 2009-2025
-Between 2009-2015, migration into Germany increased from just under 800,000 people per year to over 2 mio.
-people per year gessler2023politicization (41) in the context of the war in Syria.
-The peak in migration in 2015 has since been associated with the label ?migration crisis?
-gessler2023politicization (41, 36).
+Between 2009 and 2015, annual migration to Germany rose from just under 800,000 to more than 2 million people, peaking in 2015 in the context of the Syrian war (Gessler and Hunger, 2023).
+This period has since been associated with the label ?migration crisis?
+(Gessler and Hunger, 2023; Fröhlich, 2025).
 Since then, the right-wing AfD has made electoral gains and entered the German federal parliament in 2017.
-Another spike in migration occurred in 2022 due to the Russian invasion of Ukraine bamf2025 (33).
-Alongside these historical events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
+Another wave of migration followed the 2022 Russian invasion of Ukraine (Federal Office for Migration and Refugees, 2025).
+Alongside these events in the last four parliamentary periods (2009-2025), substantial shifts in solidarity occurred.
 4.3.1 Overall solidarity distribution
-Fig.
-13 shows the (anti-)solidarity trends according to Llama-3.3-70b annotations between 2009-2025.
-We observe a decrease in solidarity between 2014-2018, a somewhat stable period between 2018-2022, and another decrease in solidarity 2022-2025.
-The decrease in solidarity is associated with an almost corresponding increase in anti-solidarity In terms of (anti-)solidarity subtypes, we observe a peak in compassionate solidarity in 2014, followed by a sharp decrease.
-The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity.
-In terms of anti-solidarity, we observe a strong increase in group-based anti-solidarity, accounting for over 50% of all migrant-related statements that either expressed solidarity or anti-solidarity.
-Exchange-based anti-solidarity exhibited a small increase between 2015-2020 but dropped again with the beginning of the last parliamentary period in 2021.
-Overall, the 2020s mark the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements.
-This key finding is not only visible in the data provided by Llama-3.3-70B but also in human annotations (Figure 2).
-To provide more insight into this finding, we contextualize it with recent Germany history and a deeper look into party distribution.
 (a) Solidarity and anti-solidarity distribution
 (b) Solidarity and anti-solidarity frames distribution
-Figure 13: Fig.
-13(a) shows the fraction of solidarity, anti-solidarity, and mixed stance towards migrants from 2009 to 2025 (instances classified as ?None?
-are not shown, though included in the calculation).
-Fig.
-13(b) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes according to Llama-3.3-70B, where percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year).
-The spike in compassionate solidarity 2014 coincides with a time of heightened public solidarity with Syrian refugees: 15.4% of German survey respondents indicated solidarity as the most important personal value in 2014, up from 12.3% in 2013 (DeLeur2023, 23, p.
+Figure 15: (Anti-)solidarity trends for the Migrant category, 2009–2025.
+Figure˜15(a): Yearly DSL-adjusted shares of solidarity, anti-solidarity, and mixed stance (see definition (c) in Section˜3.2.2).
+The none category is included in the normalization but omitted from the visualization.
+Figure˜15(b): Yearly DSL-adjusted shares of solidarity (left) and anti-solidarity (right) subtypes (see definition (c) in Section˜3.2.2).
+Within each year, subtype shares are renormalized across the eight solidarity and anti-solidarity categories (excluding mixed and none), summing to 100%.
+Figure˜15 shows (anti-)solidarity trends from 2009–2025 (DSL-adjusted).
+Solidarity declined from 2014 to 2019, increased modestly between 2020 and 2022, and then saw a slight decline after 2022.
+This decrease in solidarity is associated with an almost corresponding increase in anti-solidarity.
+In terms of (anti-)solidarity subtypes, compassionate solidarity peaks around 2014–2016, drops sharply, briefly rebounds in 2021, [^13] This rebound coincides with the COVID-19 context, rising asylum applications, humanitarian responses to the Taliban takeover in Afghanistan, and the formation of the SPD–Greens–FDP coalition government (Brandt et al., 2024).
+and declines again.
+The decrease in solidarity since 2022 corresponds to a decrease across solidarity subtypes, especially in compassionate and exchange-based solidarity.
+In terms of anti-solidarity, group-based anti-solidarity rose strongly after 2017, accounting for over one third of all stanced migrant-related statements.
+Compassionate anti-solidarity spikes around 2020 and then declines after 2022.
+Overall, the period since 2019 marks the first time in post-war history that the fraction of anti-solidarity statements exceeds that of solidarity statements in the German parliament.
+This key finding is visible both in the DSL-adjusted estimates but also in human annotations (Figure˜2).
+We next contextualize it with recent Germany history and a deeper look into party distribution.
+The 2014 spike in compassionate solidarity coincided with heightened public support for Syrian refugees: 15.4% of German survey respondents named solidarity their most important value in 2014, up from 12.3% in 2013 (De Leur et al., 2023, p.
 38).
-Syrian refugees were commonly framed as needing help, motivating calls for a ?German welcome culture?
-Gebauer2023 (39) (see Example 15 in Table 14 in the Appendix).
-Over the course of 2015, a counter-narrative emerged, framing the influx of refugees as overburdening German capacities and as a ?refugee crisis?
-Gebauer2023 (39) (see Example 16).
-During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests and moved notably right, followed by substantial electoral gains in both state and national elections and the AfD entering the German parliament in 2017 Frey2020 (35, 39).
-Regarding the broader European context, DeLeur2023 (23) diagnose a ?solidarity fatigue?
-across Europe after 2015, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level.
-The counter-narratives mentioned by Gebauer2023 (39, 35, 23) match several notions of anti-solidarity according to thijssen2022s (98): refugees as an outgroup that stands in competition to the German population (group-based anti-solidarity), refugees as taking away resources and not contributing sufficiently (exchange-based anti-solidarity), and refugees as not deserving of compassion (compassionate anti-solidarity); see Example 17 where refugees are described as not contributing, straining resources, and threatening the welfare of the ?real?
-German population.
-We also note a general shift towards right-wing sentiments in the general German population Zick2023 (108), which may be associated with heightened anti-solidarity.
-More recently, the Russian invasion in Ukraine and the subsequent migration of Ukrainian refugees into Germany triggered another spike in public solidarity, albeit much less prominent Weber2024 (104).
-In our data, solidarity increases slightly in 2022, followed by a continued, rapid decline.
-Bojarskich2025 (12) identified several factors promoting anti-solidarity towards Ukraine, such as conspiracy mentality and populist voting, but also fears about economic and societal consequences, indicating exchange-based (insufficient contribution) or group-based (competition between groups) anti-solidarity (Example 18).
-Additionally, anti-Ukrainian framing in parliamentary debates may have been reinforced by the AfD’s long-standing pro-Russian alignment arzheimer2023russia (5), including opposition to sanctions and calls to end German military aid to Ukraine.
-In contrast to the postwar period, (anti-)solidarity shifts in the German parliament between 2009 and 2025 appear less attached to specific legislative initiatives.
-Because of Germany’s membership in the European Union, migration policy has become more complex and is negotiated between European institutions and member states geddes2020migration (40).
+Refugees were commonly framed as needing help, motivating calls for a ?German welcome culture?
+(Gebauer, 2023) (see Example 20 in Table˜23 in the Appendix).
+By early 2016, a counter-narrative emerged portraying refugees as overburdening German capacities and as contributing to a ?refugee crisis?
+(Gebauer, 2023) (see Example 20).
+During the same period, the party AfD—founded in 2013 with a focus on EU criticism—aligned itself with anti-migrant protests, shifted further towards the right of the political spectrum, and entered parliament in 2017 after major electoral gains (Frey, 2020; Gebauer, 2023).
+Regarding the broader European context, De Leur et al.
+(2023) diagnose a post-2016 ?solidarity fatigue?, driven by an inability to sustain prolonged compassion on the individual level and dwindling resources on the institutional level.
+These counter-narratives match several notions of anti-solidarity according to Thijssen and Verheyen (2022): refugees as an outgroup competing with German population (group-based, see Example 22), as undeserving of compassion (compassionate, see Example 23), and as straining resources without contributing (exchange-based, see Example 21).
+We also note a general shift towards right-wing sentiments in the general German population (Zick et al., 2023), which may be associated with heightened anti-solidarity.
+More recently, the Russian invasion of Ukraine and the resulting arrival of Ukrainian refugees in Germany, marking the largest net migration since 1950 (Atanisev et al., 2024), triggered another, though less prominent, spike in public solidarity (Weber et al., 2024a).
+In our data, solidarity increases slightly in 2022, followed by a continued decline.
+Bojarskich et al.
+(2025) identified several factors promoting anti-solidarity towards Ukraine among the German public, such as conspiracy mentality (i.e.
+receptivity to contrarian narratives and disinformation emphasizing the risks of supporting Ukraine) and support for populist parties, but also fears about economic and societal consequences, indicating group-based (competition between groups) or exchange-based (insufficient contribution) anti-solidarity (Example 21).
+Anti-Ukrainian framing may also have been reinforced by the AfD’s pro-Russian stance, including opposition to sanctions and military aid (Arzheimer, 2023).
+In contrast to the postwar period, (anti-)solidarity shifts between 2009 and 2025 appear less tied to specific legislation, likely because much of German migration policy is negotiated at the EU level (Geddes et al., 2020).
 Additionally, the AfD integrated anti-migrant talking points across a wide range of legislative topics.
-As such, we do not look deeper into specific laws passed during this time frame but rather focus on the trends for specific parties.
-Figure 14: Raw yearly counts of parliamentary statements by selected political parties, 2009–2025.
+As such, we focus less on legislation and more on party-specific trends.
+Figure 16: Raw yearly counts of parliamentary statements relating to migrants by selected political parties, 2009–2025.
 4.3.2 Distribution across parties
@@
 (f) Die Grünen
-Figure 15: Distribution of solidarity and anti-solidarity subtypes over time (2009–2025) for selected German parties.
-For each party, percentages are normalized within each year across all eight subtypes (thus, the curves in both panels together sum to 100% per year).
-Note that AfD was first represented in the Bundestag in 2017, and the FDP was not represented from 2013 to 2017.
-Therefore, data points are absent for these parties during those respective periods.
-Fig. 14 shows the raw numbers of migration-related statements per party and year, Fig. 15 shows the fine-grained solidarity trends.
+Figure 17: DSL-adjusted yearly shares of (anti-)solidarity subtypes across Bundestag parties, 2009–2025 (see definition (c) in Section˜3.2.2).
+Within each year and party, subtype shares are renormalized across the eight solidarity and anti-solidarity categories so that they sum to 100%.
+Note: the AfD entered the Bundestag only in 2017, while the FDP was absent from 2013–2017; accordingly, no observations appear for these periods.
+Figure˜16 shows the raw numbers of migration-related statements per party and year, Figure˜17 shows the fine-grained solidarity trends.
 The right-wing AfD
-utters almost no expression of solidarity and the anti-solidarity subtypes remain relatively stable over time, with group-based anti-solidarity being consistently dominant (ca.
-60% in 2018-2019 and ca.
-75% in 2021-2025), although a notable number of samples with exchange-based solidarity (ca.
-25%) occurs in the parliamentary period 2017-2021.
-Although the AfD is not the largest faction in parliament, they account for the largest share of anti-solidarity samples overall (72.5%).
-The sharp rise of anti-solidarity after 2017 (Figure 13(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure 14).
-This can be explained by the fact that many AfD speeches, across topics, contain references to migrants illegally crossing the border, criminality of migrants, or migrants imposing additional burden on the German society leaving less resources to address the actual topic of legislative debate (see Example 19 in Table 14 in the Appendix where an AfD member argues that refugees displace elderly people from care homes).
-The center-right CDU/CSU and the liberal FDP
+expresses almost no solidarity, with anti-solidarity stable over time and dominated by group-based frames, fluctuating around 50% throughout 2017-2025, although a notable number of instances with compassionate and exchange-based anti-solidarity (ca.
+25%) occurs, with the two frames alternating across years.
+Though one of the smallest factions in parliament, the AfD accounts for most anti-solidarity instances overall (53.3% over the whole period).
+The sharp rise of anti-solidarity after 2017 (Figure˜15(a)) aligns with the AfD’s entry into parliament and their rapid increase in migration-related utterances (Figure˜16).
+Many AfD speeches, across topics, link migrants to illegal border crossings, criminality, or additional burdens on the German society, without necessarily addressing the actual topic of legislative debate (see Example 21 in Table˜23 in the Appendix ).
+The center-right CDU/CSU and liberal FDP
 changed the most in our data.
-In the period 2009-2015, we observe a clear majority of solidarity samples, with a spike of compassionate solidarity for the CDU/CSU in 2014 (the FDP was not in parliament at that time).
-The main difference between both parties is more emphasis on exchange-based solidarity in the FDP, e.g.
-the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: skilled labor migration law) in 2023 (see Example 20, where the FDP highlights the law as enabling foreign workers to contribute to the German economy).
-However, for both parties, we observe a rise in group-based anti-solidarity, accounting for 80% of all samples in 2025.
-When inspecting the samples, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, which is classified as group-based anti-solidarity by Llama-3.3-70B (see Example 21 where a CDU/CSU speaker frames migration as a threat to political stability and an excessive burden on local infrastructure, with calls for deportations, faster procedures, and rejections at the border; and Example 22 where an FDP speaker portrays asylum as being misused for state-subsidized residence).
-The drastic shift toward group-based anti-solidarity over the last three years is insufficiently explained by recent literature.
-Partially, it may be that politicians from CDU/CSU and FDP take over populist rhetoric from AfD politicians who entered parliament since then Esguerra2023Accommodation (31).
-However,Lewandowsky2022 (65) and kesting2018discourse (53) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality.
-We also note that the first shift toward group-based anti-solidarity already occurred before the AfD enters parliament in 2017.
-Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population Zick2023 (108) by appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD.
-Overall, we cannot establish a full explanation but encourage further research into this striking and rapid trend.
+In 2009-2015, most statements expressed solidarity, with both parties focusing mostly on compassionate solidarity.
+However, the main difference between the two parties lies in their framing: the CDU/CSU emphasizes group-based solidarity, whereas the FDP foregrounds exchange-based solidarity, for instance by stressing the contributions of incoming skilled workers to the German economy, motivating a new Fachkräfteeinwanderungsgesetz (en.: Skilled Labor Migration Law) in 2023 (see Example 21).
+From the late 2010s onward, both parties shift toward anti-solidarity, most visibly in the CDU/CSU, where group-based anti-solidarity and compassionate anti-solidarity frames rise after the 2015 refugee crisis, decline during the Ukrainian refugee inflow in 2022, and resurge by 2025, reaching 33% and 24% of all frames, respectively.
+For the FDP, a comparable shift appears by 2025, where exchange-based, group-based, and compassionate anti-solidarity account for 16%, 12%, and 12%, respectively.
+When inspecting the instances, the framing of migration is mostly in legalistic terms, such as regulating and controlling migration and reducing numbers, marking migrants as an outgroup, for which group-based anti-solidarity has the highest DSL estimate among the anti-solidarity subtypes (see Example 22) and portraying migration as misuse of welfare benefits, corresponding to exchange-based anti-solidarity (see Example 22).
+The shift toward compassionate and group-based anti-solidarity over the last three years is insufficiently explained by recent literature.
+One possibility is that CDU/CSU and FDP adopted populist rhetoric from AfD after 2017 (Esguerra et al., 2023).
+However, Lewandowsky et al.
+(2022) and Kesting et al.
+(2018) did not observe effects of the AfD entering parliament on polarization, the use of populist language, or discourse quality.
+Notably, the first shift occurred even before the AfD entered parliament.
+Alternatively, it may be that CDU/CSU and FDP respond to an overall right-wing shift in the German population (Zick et al., 2023), appealing to voters with anti-migration rhetoric and thereby providing a more moderate (but still anti-solidarity) alternative to the AfD.
+Overall, our findings encourage further research into this striking and rapid trend.
 The center-left SPD
-also exhibits a spike in compassionate solidarity in 2014 and a slow decline in solidarity towards more group-based anti-solidarity (ca.
-20% in 2024).
-Part of this is explained by mis-classifications of Llama-3.3-70B because SPD speakers often cite anti-solidarity stances of other parties (especially AfD) and these citations are mis-classified as anti-solidarity by the speaker (see Example 23 where an SPD speaker cites AfD proposals putting foreigners under general suspicion in order to criticize them).
-However, there are also samples where SPD speakers refer positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 24).
+shows predominantly compassionate solidarity, with a spike in 2020, and is closely followed by group-based solidarity, which gains prominence post-2019, and almost taking over in 2025.
+This is consistent with the party’s humanitarian commitments and stronger emphasis on integration (Orhan, 2025).
+Additionally, exchange-based solidarity is present, at times reaching around 20%.
+In line with this perspective, SPD has repeatedly framed Germany as an Einwanderungsland (en.: country of immigration) and supported immigration policies aimed at attracting qualified workers to meet labor market demands (Sozialdemokratische Partei Deutschlands, 2021).
+However, there are also instances where SPD speakers express anti-solidarity by referring positively to their own policies having enabled deportations of declined asylum seekers or the difficulties associated with managing migration, similar to the frames of CDU/CSU and FDP speakers (see Example 22).
 The left Green and Linke parties
-only display very few instances classified as anti-solidarity and almost all of them are misclassifications where Green or Linke speakers cite anti-solidarity statements or parliamentary motions by other parties (see Example 25 ).
-These samples are nonetheless interesting because they express a polarizing influence of the AfD entering parliament, leading left-wing parties to positioning themselves in contrast Atzpodien2022 (6).
-Overall, Die Grünen initially emphasized compassionate solidarity, but later shifted to group-based appeals stressing integration and rights.
-In the 2024/25 campaign, their rhetoric focused strongly on exchange-based arguments about migrants’ economic and societal contributions Gruene2025 (17) (see Example 26).
-Similarly, Die Linke began with predominantly compassionate appeals, framing migration as a humanitarian duty, but later emphasized migration as integral to society and advanced a detailed pro-immigration agenda, thereby increasing group-based solidarity.
-In the 2024/25 campaign, they moved further toward empathic solidarity, criticizing stereotyping and positioning themselves against exclusionary discourses DieLinke2021 (25, 26) (see Example 27).
-A special case is the BSW, which split off from the Linke party in 2024, partially due to explicit anti-migration positions.
-However, our data set contains only very few samples from this party (38 statements), so we cannot provide a reliable quantitative analysis.
+only display very few instances classified as anti-solidarity.
+Overall, since 2021, Die Grünen moved from compassionate solidarity to group-based appeals, stressing integration and rights.
+In the 2024/25 campaign, and in the 2024/25 campaign strongly emphasized exchange-based arguments about migrants’ economic and social contributions (Bündnis 90/Die Grünen, 2025) (see Example 23).
+Die Linke likewise emphasized compassionate solidarity but since 2021 started increasingly framing migration as integral to society, advancing a detailed pro-immigration agenda and strengthening group-based solidarity.
+In 2024/25, the party continued to oppose stereotyping and position themselves against exclusionary discourses (Die Linke, 2021, 2025) (see Example 23).
+A special case is the BSW, which split from Linke party in 2024, partially due to explicit anti-migration positions.
+However, our data set contains only very few instances from this party (38 statements), so we cannot provide a reliable quantitative analysis.
 5 Discussion and Conclusion
-In this paper, we have evaluated LLMs on the complex task of annotating (anti-)solidarity in German parliamentary speeches and interpreted the annotated data to derive new insights into (anti-)solidarity towards migrants in the post-war period as well as recent past.
-In more detail, we provide both methodological and sociological insights:
-Methodological perspective.
-We find that instruction-tuned LLMs like Llama-3.3-70B and GPT-4 are effective tools for annotating solidarity frames in parliamentary debates, with open models offering a more cost-effective alternative to closed ones.
-In our experiments, Llama-3.3-70B performed competitively with GPT-4, especially in few-shot settings.
-Still, no model consistently outperformed GPT-4, and even GPT-4 did not fully attain the human upper bound.
-Existing research offers mixed evidence on the reliability of LLMs for annotation tasks: some highlight their limitations in handling subjective or expert-defined taxonomies ziems2024can (109), while others show they can substitute or outperform human coders heseltine2024large (43, 99), or suggest that smaller, fine-tuned models are more reliable in political science applications wang2024selecting (102).
-In our case, model performance generally improved with model size, but also varied by target group, historical period (with older data proving more difficult), and inference method.
-These findings highlight the absence of a one-size-fits-all solution: rather than relying on a single model, both model and method selection must be adjusted to the requirements of the specific task.
-Moreover, we emphasize the need for validation at every stage of the analysis pipeline.
-Even keyword search as a first filtering step can introduce biases in the data or can aggregate data that concerns very different sub-populations (such as different groups of migrants in the German post-war period).
-Regarding the LLM annotation itself, we validated model predictions against a large data set of human annotations (aggregating a year of labor involving multiple annotators), test multiple inference configurations (including zero- and few-shot prompting), and validate long-term trends produced by our two best-performing models (Llama-3.3-70B and GPT-4) against human annotations.
-We also conducted error analyses to identify where and why models fail to make sure that outputs can be reliably interpreted.
-This aligns with recent work advocating for human-in-the-loop workflows and partial human validation when using LLMs for annotation tasks, to ensure output quality and interpretability pangakis2024keeping (75, 100).
-Despite these efforts, it remains essential to acknowledge model-specific weaknesses when interpreting results.
-For example, we observed that Llama-3.3-70B occasionally misclassifies anti-solidarity, particularly in instances where criticism of anti-migrant rhetoric is sometimes mistaken for anti-solidarity framing itself.
-In such cases, models may overlook logical direction or target relevance and assign misleading labels despite otherwise coherent reasoning zhang2025mpvstance (107).
-One possible reason for this behavior could be the influence of complex political language or long prompt context, which may distract the model from the intended target huang2025breaking (48).
-Importantly, we could only reveal model errors because we checked the primary data before drawing conclusions.
-Social science perspective.
-Beyond the methodological contributions, our study also demonstrates how an LLM-based pipeline can support a theoretically informed quantitative analysis through two illustrative case studies: one historical (1949–1957) and one contemporary (2009–2025).
-For the early post-war period (1949–1957), we demonstrated how solidarity was shaped and institutionalized in the context of national reconstruction.
-While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insight for historical and social science research.
-First, it enables us to move beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity, and shifting rhetorical strategies across different migrant groups.
-Second, our approach helps integrate migration histories that are often treated separately, such as expellees, DPs, and refugees, into a shared analytical framework, allowing for direct comparison of how solidarity was articulated across groups.
+In this paper, we have evaluated LLMs on the complex, theory-driven task of annotating (anti-)solidarity in German parliamentary speeches and assessed whether their outputs support valid downstream social-scientific inference.
+We then use a validated and statistically corrected annotation pipeline to derive new insights into solidarity and anti-solidarity toward migrants in the postwar period as well as the recent past.
+In more detail, we provide both methodological and social scientific insights:
+Methodological perspective
+Our results show that strong instruction-tuned LLMs can reach human agreement in annotating solidarity frames in parliamentary debates, where GPT-5 performs best overall, while open models such as gpt-oss-120B and Llama-3.3-70B provide competitive and more cost-effective alternatives.
+At the same time, performance varies systematically by target group, historical period, prompting strategy, and model family, with older data and fine-grained labels proving substantially harder.
+Generally, existing research presents mixed evidence on the reliability of LLMs for annotation tasks: some studies mention limitations in handling subjective or expert-defined taxonomies (Ziems et al., 2024) or find that smaller fine-tuned models are more reliable in political science applications (Wang et al., 2024), while others show that LLMs can substitute for or outperform human coders (Heseltine and Clemm von Hohenberg, 2024; Törnberg, 2024).
+Our findings are in line with a more cautious interpretation: LLMs are useful annotation tools, but they should not be viewed as neutral coders, nor should their outputs be treated as ?ground truth?
+data; rather, they are ?measurement tools that require rigorous validation, calibration, and documentation?
+(Baumann et al., 2025).
+This idea is reinforced by our error analysis, which shows that even the strongest models in our study make systematic mistakes.
+For example, Llama-3.3-70B occasionally misclassifies anti-solidarity, mistaking criticism of anti-migrant rhetoric for anti-solidarity itself, and all models show a general tendency to overpredict the None label.
+More broadly, they underestimate Mixed cases and compress fine-grained subtype distinctions into broader categories, especially group-based subtypes of (anti-)solidarity.
+These errors become more severe on historical instances, where less explicit forms of solidarity are often missed.
+More importantly, these errors are not independent across models: they are often correlated, which limits the benefit of ensemble methods.
+This matters for downstream analysis because small but systematic misclassifications can accumulate into biased temporal trends.
+Recent work highlights this problem: high agreement with human annotators can coexist with systematic misclassifications, meaning that benchmark performance alone does not guarantee valid downstream inference (Bavaresco et al., 2025; Gligoric et al., 2025).
+Even highly accurate LLM-based surrogates can therefore result in misleading substantive conclusions when used directly in downstream analyses, because doing so ignores measurement error between gold-standard and surrogate labels (Egami et al., 2023).
+Consistent with this, our study finds that DSL-corrected soft-label outputs best recover human-derived trends, especially in fine-grained categories.
+This suggests that large-scale LLM annotation, combined with a limited set of expert annotations (around 2k in our case), is a viable strategy for downstream inference.
+The benefits are greatest when soft labels are aggregated across multiple strong models and corrected with DSL, rather than treated as hard predictions from a single model.
+Recent benchmarking further suggests that DSL performs especially well in low-sample settings, often outperforming alternative correction methods (de Pieuchon et al., 2025).
+From a resource-allocation perspective, this pipeline combines expensive expert labels with low-cost model labels in a way that improves inference.
+In our project, expert annotation cost over €29,000 for 3,500 instances, whereas model annotation scaled much more cheaply, whether via open-weight models such as gpt-oss-120B and Llama-3.3-70B on in-house NVIDIA A40 GPUs or via GPT-4, which annotated one third of the dataset for about €500.
+A smaller set of costly human labels provides the reference basis for validation and correction, while a much larger set of low-cost model labels extends coverage to corpus scale.
+Social science perspective
+Beyond methodological contributions, our study shows how an LLM-based pipeline can support theoretically informed quantitative analysis through two case studies: one historical (1949–1957) and one contemporary (2009–2025).
+For the early postwar period (1949–1957), we showed how solidarity was shaped and institutionalized in national reconstruction.
+While this era has been examined in previous historical research, our LLM-based, fine-grained analysis offers additional insights.
+First, it moves beyond the assumption that solidarity during this period was simply ?high?: while historical consensus holds that it largely benefited ethnic Germans, our data reveal coexisting forms of (anti-)solidarity and shifting rhetorical strategies across migrant groups.
+Second, our approach integrates migration histories that are often treated separately (expellees, displaced persons, and refugees) into a shared analytical framework, enabling direct comparison of how solidarity was articulated.
 This also sheds light on the emerging distinction between early DPs and the broader category of Ausländer, including those later identified as Gastarbeiter, showing that solidarity was not uniformly absent but selectively expressed.
 For the recent past (2009-2025), we see a striking shift from solidarity to anti-solidarity following the 2015 migration crisis.
-This may be partially explained by solidarity fatigue DeLeur2023 (23) or may be part of a larger right-wing shift Zick2023 (108) on migration issues.
-While stricter migration stances across Europe have been analyzed before, our study is the first to quantify the shift from solidarity to anti-solidarity in German parliament.
-Our data also revealed that the shift is most striking in the liberal and conservative parties, motivating further research.
-This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives aktas2024rise (2), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists bayerlein2021chasing (8).
-Comparative survey evidence further shows that anti-immigrant attitudes remain a central driver of right-wing populist support across the continent messing2021anti (70).
-This type of fine-grained, longitudinal analysis of solidarity would hardly be feasible at such scale using manual annotation alone which, in our project, cost over €29,000 for only 3,500 instances.
-By contrast, with access to computing resources (e.g., in our case, in-house hardware with NVIDIA A40 GPUs), open-weight models such as Llama-3.3-70B can annotate tens of thousands of instances at minimal additional cost.
-Even closed models like GPT-4 are far cheaper, with one third of our dataset annotated for about €500.
-While machine annotation is therefore not ?free?, it is vastly more scalable than human annotation and enables large-scale analyses that would otherwise be financially prohibitive.
-In this sense, our method bridges ?distant reading?
-approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded systematic interpretation of parliamentary speech while retaining full corpus coverage.
+This may reflect solidarity fatigue (De Leur et al., 2023) or a larger societal right-wing shift (Zick et al., 2023) on migration.
+While stricter migration stances across Europe have been studied, our study is the first to quantify this shift in German parliamentary debates.
+Our data also revealed that the shift is not only explained by the right-wing AfD party entering parliament, but it also visible in the conservative CDU/CSU party and the liberal FDP party, motivating further research.
+This development resonates with broader European trends, where populist radical right parties have increasingly mobilized around anti-immigration narratives (Aktas, 2024), and even mainstream conservative parties have shifted their positions on migration in response to the electoral pressure of right-wing populists (Bayerlein, 2021).
+Comparative surveys in Europe likewise show anti-immigrant attitudes remain a central driver of right-wing populist support (Messing and Ságvári, 2021).
+Taken together, our findings address a broader question in CSS: whether LLMs can support valid, theory-driven analysis of complex political concepts at scale.
+Our answer is yes, but only under clear conditions.
+LLMs are not neutral replacements for expert coders, and their outputs cannot be used uncritically as if they were ground-truth labels, particularly for theory-grounded annotation tasks.
+Yet when combined with human reference annotation, systematic validation, error analysis, and DSL-based correction, they can support serious social-scientific research by revealing patterns at a scale that would otherwise be accessible only through qualitative analysis.
+In this sense, our pipeline bridges ?distant reading?
+approaches common in large-scale text analysis with elements of ?close reading?, enabling theoretically grounded and systematic interpretation of parliamentary speech while retaining full corpus coverage.
 6 Limitations and Future Work
 While our analysis was performed carefully to avoid methodological pitfalls, we nonetheless acknowledge remaining limitations of our study.
-Human Label Variation:
-Even after months of training, discussion and improvements of the annotation manual, we still observe substantial disagreement between human annotators, especially on the fine-grained level.
-This indicates that (anti-)solidarity is intrinsically a challenging concept to annotate, especially on parliamentary speeches, and that reasonable disagreement may persist (e.g.
-whether anti-solidarity should be regarded as group-based or exchange-based when competition between groups around limited resources is highlighted).
-For this investigation, we used the majority label, but future work should consider methods that explicitly account for human label variation and embrace it plank-2022-problem (76).
-Political Bias:
-It may be that the LLMs compared in this paper may be subject to political bias as investigated by several prior works motoki2024more (72, 85).
-In particular, it may be that a left-wing political bias leads to more anti-solidarity being annotated.
-However, we do not observe a systematic bias toward anti-solidarity in the error analysis (Fig.
-6); instead, models often mislabel solidarity and anti-solidarity samples as ?None?.
-It may also be that the human annotators are biased towards certain labels, but this would need to affect all annotators equally because we do not observe a systematic disagreement towards one label, either.
-Finally, even if bias were present, this would not explain differences and trends in (anti-)solidarity (which are the main subject of our analysis) but rather the overall level of (anti-)solidarity.
-Analysis of (Anti-)solidarity towards women:
-In the same way we performed an analysis of (anti-)solidarity trends towards migrants, the (anti-)solidarity towards women could be analyzed.
-For example, recent work by Diabate2023 (24) suggests that there may be a shift to the left in Germany regarding gender, which would be an opposite trend compared to what we found regarding migration.
-Due to space and time constraints, such an analysis is not included in this work but appears highly promising for future work.
-Further model variations:
-We considered a range of models as well as different prompts, and fine-tuning (for Llama models and BERT models) but a vast range of possible model variations remains to be explored, such as ensembles of models, less quantized models, newer models, etc.
-It may well be that human annotation quality can be fully matched with more refined tools.
-Temporal Reliability:
-Both model performance (macro F1) and human agreement vary over time, with notable dips in some decades (e.g., the 1960s for Llama-3.3-70B; Fig.
-16).
-This reflects the difficulty of annotating historical language, but overall trend correlations between GPT-4, Llama-3.3-70B, and human annotations remain high.
-Since our analysis emphasizes broad temporal trends rather than isolated data points, these fluctuations do not undermine the validity of our conclusions.
+Human Coded Reference
+Our evaluation and DSL correction rely on human annotations as the best available expert-coded reference.
+At the same time, some degree of disagreement remains, especially at the fine-grained level, even after extensive training and refinement of the annotation manual.
+This reflects the intrinsic difficulty of interpreting theoretically informed categories in parliamentary speech and the possibility of reasonable disagreement (e.g.
+over whether competition over resources is group- or exchange-based anti-solidarity, or whether a text is better read as compassionate or group-based solidarity.).
+DSL improves correlation with human annotations, but it cannot remove uncertainty where the reference itself remains interpretive.
+Limits of Correction with DSL
+DSL improves recovery of human-derived trends, especially relative to raw hard-label predictions, but it does not remove all sources of error.
+Its performance depends on the quality of the surrogate model outputs, the coverage of the human-coded subset, and the sampling design.
+In settings with rare labels or sparse historical coverage, correction may remain unstable.
+Temporal Reliability
+Both model performance and human agreement fluctuate over time (e.g.
+the 1960s for several models; Figure˜18), reflecting the challenge of historical language.
+Although our correction framework mitigates part of this problem, historical variation remains a challenge for both annotation and downstream inference.
+Still, because our focus is on broad temporal trends rather than individual years, these fluctuations are less consequential for our main substantive conclusions.
+(Anti-)Solidarity towards Women
+Similar to migrants, one could analyze (anti-)solidarity trends toward women.
+Diabaté et al.
+(2023) notes a leftward shift in Germany on gender issues, potentially contrasting with our migration findings.
+Due to space and time constraints, we do not pursue this here, but consider it a promising avenue for future work.
+Political Bias
+LLMs compared in this paper may be subject to political bias as investigated by several prior works (Motoki et al., 2024; Rettenberger et al., 2025), with left-wing political bias potentially leading to more anti-solidarity being annotated.
+Yet our error analysis (Figure˜6) shows no systematic skew toward anti-solidarity; models more often mislabel instances as ?None?.
+Annotator bias towards certain labels is also possible but would need to be uniform across annotators, which we do not observe.
+Even if present, bias would affect only overall levels of (anti-)solidarity, not the differences and trends central to our analysis.
 Appendix
 Appendix A List of Keywords
-For Frau (en.: woman) we use Frauen, Frau, Mutter, Mädchen, Mütter, Ehefrau, Müttern, Hausfrauen, Hausfrau, Ehefrauen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Großmutter, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
+For Frau (en.: woman), we use Frauen (and Frau), Mütter (and Mutter, Müttern), Großmutter, Hausfrauen (and Hausfrau), Ehefrauen (and Ehefrau), Mädchen, Frauenförderung, Frauenquote, Dienstmädchen, Fräulein, Kriegerfrauen, Arbeiterfrauen, and Trümmerfrauen.
 Since Frau is the German word for woman but also for Mrs./Ms., we only include occurrences of the word Frau that are not followed by a capitalized word (which are probably surnames).
-For Migrant (en.: refugee) we use Flüchtlinge, Ausländer, Flüchtlingen, Zuwanderung, Vertriebenen, Ausländern, Asylbewerber, Migranten, Migration, Heimatvertriebenen, Aussiedler, Einwanderung, Ansiedler, Vertriebene, Zuwanderer, Asylbewerbern, Flüchtling, Heimatvertriebene, Sowjetzonenflüchtlinge, Aussiedlern, Einwanderer, Asylsuchenden, Asylsuchende, Bürgerkriegsflüchtlinge, Zuwanderern, Ansiedlern, Migrantinnen, Vertriebener, Emigranten, Kriegsflüchtlinge, Ausländerinnen, and Immigranten.
-When running stability tests (see Subsubsection 4.1.1), we ensure that each sample includes at least 5 keywords, covers at least 10% of the dataset, and spans at least 75% of the timeline, such that enough data is present to create the plots.
+For Migrant (en.: migrant) we use Flüchtlinge (and Flüchtling, Flüchtlingen), Ausländer (and Ausländern, Ausländerinnen), Asylsuchende (and Asylsuchende, Asylsuchenden), Asylbewerber (and Asylbewerbern), Heimatvertriebene (and Heimatvertriebenen), Vertriebene (and Vertriebenen, Vertriebener), Aussiedler (and Aussiedlern), Einwanderer (and Einwanderung), Ansiedler (and Ansiedlern), Zuwanderer (and Zuwanderern, Zuwanderung), Migranten (and Migrantinnen, Migration), Sowjetzonenflüchtlinge, Bürgerkriegsflüchtlinge, Kriegsflüchtlinge, Emigranten, and Immigranten.
+For robustness, we conducted a stability test by randomly sampling 200 subsets of at least 5 keyword groups (grouping inflectional variants) each, ensuring that each instance covered at least 10% of the dataset and spanned at least 75% of the timeline (such that enough data is present to create the plots).
+For each subset, we re-estimated the high-level DSL trends (as defined in Section˜3.2.2) and compared the resulting time series to the full-sample trend using Pearson’s correlation coefficient.
+The correlations remain high on average, with mean values of 0.93 for solidarity, 0.88 for anti-solidarity, and 0.82 for mixed; the 25% quantiles are 0.89, 0.85, and 0.75, respectively.
+This indicates that the main aggregate trends are broadly robust to keyword composition, although robustness is weaker for mixed.
 For the analysis of frequency of keywords over time, we calculate the percentages normalized for each keyword, i.e., a value of pp% in year $y$ implies that in year $y$ pp% of all sentences with this keyword occurred.
-The trends are shown in the Fig.
-19 in the Appendix.
+The trends are shown in the Figure˜21 in the Appendix.
 Appendix B Parties
 For speaker and party extraction, we parse the XML protocols of the Bundestag and Reichstag.
-For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by poradaautomatische (77).
+For modern sessions (20th–21st terms), we use the official Open Data XML, which already encodes speaker and party information; for earlier sessions (1st–19th terms), we rely on the improved markup provided by Porada (2019).
 Each new intervention is marked in the XML by a <redner> element containing both the speaker’s name and their party.
@@
 Hardware and Inference Settings.
-All inference runs were conducted on an NVIDIA A40 (80GB) with temperature set to 0.6 and top-p to 0.9. For Mistral-Large-Instruct-2407, we used an NVIDIA A100 to accommodate memory requirements.
+Inference was conducted on mixed CPU and GPU environments. Closed-source models (gpt-4-1106-preview, gpt-5-chat) were accessed via API.
+Open-weight models were run locally:
+  • gpt-oss-20B on two NVIDIA A40 (80 GB) GPUs in bfloat16 precision;
+  • Llama-3-8B-Instruct, Gemma-2-9b-it, and Gemma-2-27b-it in full precision on a single A40 GPU;
+  • Mistral-Large-Instruct-2407 on a rented NVIDIA A100 in 4-bit quantization to accommodate memory requirements;
+  • Llama-3.3-70B-Instruct in 4-bit quantization and Qwen2.5-72B-Instruct on two A40 GPUs in 8-bit quantization;
+  • and gpt-oss-120B and Llama-4-Scout-17B-16E-Instruct in 4-bit quantized form (Q4_K_M and Q4_K_XL, respectively) on CPU nodes (32 threads, 128 GB RAM) via llama.cpp [^14] https://github.com/ggml-org/llama.cpp.
+All generations used temperature =0.6=0.6 and top-p =0.9=0.9.
 Prompt Format.
-Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure 24 for Migrant and in Figure 27 in the Appendix for Woman).
+Each prompt consists of: (i) clear label definitions; (ii) chain-of-thought cues (e.g., think step by step); (iii) a two-step classification format (high-level → subtype) (see the prompt text in Figure˜27 for Migrant and in Figure˜30 in the Appendix for Woman).
 In few-shot settings, one representative example is included for each class label.
-GPT-4 receives all instructions in a single prompt (see the prompt text in Figure 21 for Migrant and Figure 22 in the Appendix for Woman).
+GPT-4 receives all instructions in a single prompt (see the prompt text in Figure˜25 for Migrant and Figure˜26 in the Appendix for Woman).
 Open-weight models receive separate prompts: one for high-level classification, and one for subtype prediction if the label is solidarity or anti-solidarity.
@@
 To counter class imbalance, minority classes are oversampled to parity with the majority class.
-We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer AdamW (68) and categorical crossentropy loss.
-Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 reimers2019sentence (82).
+We finetune for 20 epochs with a learning rate of 4e-4, a warmup ratio of 0.05, linear decay, AdamW optimizer (Loshchilov and Hutter, 2017) and categorical crossentropy loss.
+Additionally, we use the multilingual SentenceBERT model distiluse-base-multilingual-cased-v2 (Reimers and Gurevych, 2019).
 We modify the network by expanding the embeddings from 512 to 1024 dimensions, then reducing them back to 512 dimensions.
@@
 The finetuning is conducted over 20 epochs with a batch size of 8, starting with a learning rate of 5e-4, a warmup ratio of 0.03, and using AdamW for optimization with learning rate decay and categorical crossentropy.
-Fine-Tuned Llama Models
-We fine-tune Llama-3-8B-It on a single A40 GPU.
-Llama-3-70B-It is trained across two A40s using PyTorch FSDP and QLoRA [^11] https://www.philschmid.de/fsdp-qlora-llama3.
-Training spans 3 epochs with a constant learning rate of 2e-4, AdamW optimizer, and max gradient norm of 0.3.
+Appendix D Error Correlation
+For two models $m_{i}$ and $m_{j}$, let $\mathcal{E}_{i}$ and $\mathcal{E}_{j}$ denote the sets of instances misclassified by each model, and let $\mathcal{E}^{\text{same}}_{ij}$ be the subset where both models predict the same wrong label.
+We compute two measures of pairwise error overlap, reported in Table˜4.
+Error Jaccard
+$J_{ij}=\frac{|\mathcal{E}_{i}\cap\mathcal{E}_{j}|}{|\mathcal{E}_{i}\cup\mathcal{E}_{j}|}$
+This measures how often two models make errors on the same instances relative to their total error sets.
+Same-Wrong Rate
+$P_{\text{same-wrong}\mid\text{both-wrong}}=\frac{|\mathcal{E}^{\text{same}}_{ij}|}{|\mathcal{E}_{i}\cap\mathcal{E}_{j}|}$
+This measures the proportion of jointly misclassified instances on which two models predict the same incorrect label. Table˜4 reports the resulting values for all model pairs.
+Model 1 Model 2 Error Jaccard P(Same Wrong $\mid$ Both Wrong)
+gpt-oss-120B Qwen-2.5 0.637 0.494
+Llama-3.3-70B Qwen-2.5 0.610 0.448
+Llama-3.3-70B gpt-oss-120B 0.574 0.456
+Table 4: Pairwise overlap in model prediction errors.
+Error Jaccard measures the proportion of shared errors relative to the union of all errors between two models.
+P(Same Wrong $\mid$ Both Wrong) measures how often two models predict the same wrong label when both are incorrect.
+Appendix E Label Distribution Distance Metrics
+We compare model and human label distributions (Section˜3.2.1) using Jensen–Shannon Divergence and total variation distance.
+Jensen–Shannon Divergence.
+Let $p$ and $q$ denote discrete probability distributions over the same label set, and let $m=\frac{1}{2}(p+q)$. The Jensen–Shannon divergence is defined as
+$\mathrm{JSD}(p\,\|\,q)=\frac{1}{2}\mathrm{KL}(p\,\|\,m)+\frac{1}{2}\mathrm{KL}(q\,\|\,m),$
+where $\mathrm{KL}(\cdot\|\cdot)$ denotes the Kullback–Leibler divergence. Lower values indicate more similar label distributions.
+Total Variation Distance.
+Total variation distance measures absolute differences in probability mass:
+$\mathrm{TV}(p,q)=\frac{1}{2}\sum_{i}|p_{i}-q_{i}|.$
+It ranges from 0 (identical distributions) to 1 (disjoint support).
+Comparison Total Variation Jensen–Shannon Divergence
+Model ensemble vs humans 0.49 (0.32) 0.39 (0.22)
+Human vs human (avg 10 splits) 0.45 (0.19) 0.35 (0.12)
+Randomized model baseline 0.75 (0.59) 0.69 (0.47)
+gpt-oss-120B vs humans 0.58 (0.36) 0.48 (0.29)
+Qwen-2.5 vs humans 0.59 (0.42) 0.50 (0.34)
+Llama-3.3-70B vs humans 0.57 (0.40) 0.48 (0.32)
+GPT-5 vs humans 0.57 (0.33) 0.47 (0.25)
+Table 5: Distribution distances between human and model label distributions (lower = closer alignment). Values are shown as fine-grained (high-level).
 Label Woman Migrant Total per label
@@
 Evaluated on the full test set, including Frau and both test sets for Migrant, grouped by decade.
-Figure 17: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined.
-Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
-Figure 18: Fine-grained confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined.
-Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
-Table 3: F1 scores per model for fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman.
+Figure 19: High-level row-normalized confusion matrices (%) for the best-performing configuration of each model on the two migrant test sets (Test 1 and Test 2, combined).
+Each row sums to 100, showing how items with a given reference label are distributed across predicted labels.
+Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators.
+The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
+Figure 20: Fine-grained row-normalized confusion matrices (%) showing the best-performing configuration for each of the selected models on the two migrant test sets (Test 1 and Test 2), combined.
+Each row sums to 100, showing how items with a given reference label are distributed across predicted labels.
+Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators.
+The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
+Table 7: F1 scores for selected models on fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman.
 The top section contains fine-grained categories, and the bold rows show performance on high-level classification.
 Scores are chosen for the best performing configuration for the model in terms of Macro F1.
-Label
-GPT-4
-Mistral
-2-Large
-Llama
-3.3-70B
-Llama
-3-8B
-Gemma
-2-27B
-Gemma
-2-9B
-BERT SBERT
-Group-based Solidarity 0.33 0.30 0.29 0.25 0.25 0.34 0.32 0.16
-Exchange-based Solidarity 0.56 0.39 0.36 0.25 0.41 0.20 0.21 0.19
-Empathic Solidarity 0.04 0.27 0.10 0.08 0.15 0.14 0.00 0.03
-Compassionate Solidarity 0.48 0.42 0.51 0.22 0.45 0.43 0.51 0.41
-Solidarity (high-level) 0.78 0.52 0.69 0.60 0.58 0.63 0.66 0.51
-Group-based Anti-solidarity 0.47 0.25 0.42 0.32 0.42 0.41 0.28 0.20
-Exchange-based Anti-solidarity 0.46 0.29 0.35 0.12 0.28 0.29 0.09 0.22
-Empathic Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Compassionate Anti-solidarity 0.00 0.18 0.07 0.00 0.19 0.11 0.17 0.12
-Anti-solidarity (high-level) 0.64 0.53 0.52 0.44 0.56 0.54 0.35 0.35
-Mixed 0.38 0.26 0.43 0.10 0.17 0.19 0.24 0.15
-None 0.67 0.60 0.63 0.56 0.57 0.58 0.59 0.51
-Label
-GPT-4
-Mistral
-2-Large
-Llama
-3.3-70B
-Llama
-3-8B
-Gemma
-2-27B
-Gemma
-2-9B
-BERT SBERT
-Group-based Solidarity 0.28 0.24 0.25 0.00 0.20 0.20 0.24 0.08
-Exchange-based Solidarity 0.31 0.00 0.00 0.00 0.22 0.30 0.35 0.15
-Empathic Solidarity 0.27 0.36 0.41 0.24 0.33 0.41 0.22 0.14
-Compassionate Solidarity 0.43 0.32 0.51 0.00 0.37 0.21 0.62 0.52
-Solidarity (high-level) 0.85 0.67 0.79 0.51 0.75 0.70 0.77 0.78
-Group-based Anti-solidarity 0.31 0.36 0.22 0.00 0.27 0.26 0.00 0.00
-Exchange-based Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Empathic Anti-solidarity 0.67 0.00 0.25 0.07 0.00 0.17 0.00 0.00
-Compassionate Anti-solidarity 0.50 0.33 0.22 0.36 0.15 0.00 0.00 0.00
-Anti-solidarity (high-level) 0.65 0.44 0.39 0.31 0.48 0.35 0.12 0.10
-Mixed 0.30 0.40 0.37 0.16 0.32 0.30 0.15 0.26
-None 0.62 0.54 0.47 0.41 0.66 0.41 0.21 0.40
-Table 4: Agreement between GPT-4, Llama-3.3-70B, and human annotations on the Migrant dataset.
-Subtable LABEL:tab:model-human-agreement shows instance-level agreement; Subtable LABEL:tab:trend-agreement reports trend-level correlation across decades.
-Comparison Cohen’s $\kappa$ / Macro F1 N
-GPT-4 vs Llama-3.3-70B 0.47 (0.58) / — 17,696
-GPT-4 vs Human 0.38 (0.57) / 0.34 (0.63) 2,084
-Llama-3.3-70B vs Human 0.36 (0.50) / 0.35 (0.61) 2,084
-Label Pearson (p)
-Group-based Solidarity 0.560 ($p=2.41\times 10^{-2}$)
-Exchange-based Solidarity 0.777 ($p=4.00\times 10^{-4}$)
-Empathic Solidarity 0.108 ($p=0.691$)
-Compassionate Solidarity 0.915 ($p=7.10\times 10^{-7}$)
-Solidarity (high-level) 0.966 ($\boldsymbol{p=1.25\times 10^{-9}}$)
-Group-based Anti-solidarity 0.945 ($p=3.65\times 10^{-8}$)
-Exchange-based Anti-solidarity 0.898 ($p=2.29\times 10^{-6}$)
-Empathic Anti-solidarity 0.160-0.160 ($p=0.553$)
-Compassionate Anti-solidarity 0.841 ($p=4.44\times 10^{-5}$)
-Anti-solidarity (high-level) 0.976 ($\boldsymbol{p=1.15\times 10^{-10}}$)
-Mixed 0.685 ($p=0.0034$)
+Label GPT-5 GPT-4 gpt-oss 120b Qwen 2.5-72B Llama 3.3-70B Llama 4-Scout Mistral 2-Large Gemma 2-27B BERT
+Group-based S 0.42 0.33 0.37 0.39 0.29 0.35 0.31 0.25 0.31
+Exchange-based S 0.43 0.60 0.55 0.42 0.39 0.44 0.38 0.43 0.22
+Empathic S 0.48 0.05 0.26 0.26 0.11 0.17 0.30 0.11 0.00
+Compassionate S 0.60 0.50 0.48 0.50 0.57 0.59 0.42 0.46 0.53
+Solidarity (high-level) 0.81 0.80 0.78 0.69 0.71 0.78 0.54 0.58 0.67
+Group-based AS 0.47 0.51 0.46 0.39 0.47 0.42 0.31 0.44 0.32
+Exchange-based AS 0.43 0.48 0.47 0.40 0.42 0.48 0.30 0.30 0.10
+Empathic AS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Compassionate AS 0.20 0.00 0.06 0.16 0.08 0.19 0.16 0.21 0.18
+Anti-solidarity (high-level) 0.70 0.69 0.66 0.59 0.56 0.63 0.57 0.58 0.39
+Mixed 0.32 0.46 0.32 0.24 0.48 0.08 0.26 0.17 0.26
+None 0.66 0.71 0.68 0.65 0.66 0.68 0.64 0.59 0.62
+Label GPT-5 GPT-4 gpt-oss 120b Qwen 2.5-72B Llama 3.3-70B Llama 4-Scout Mistral 2-Large Gemma 2-27B BERT
+Group-based S 0.45 0.28 0.40 0.41 0.25 0.25 0.24 0.20 0.24
+Exchange-based S 0.50 0.31 0.40 0.00 0.00 0.00 0.00 0.22 0.35
+Empathic S 0.27 0.27 0.34 0.46 0.41 0.31 0.36 0.33 0.22
+Compassionate S 0.62 0.43 0.59 0.42 0.51 0.34 0.32 0.37 0.62
+Solidarity (high-level) 0.86 0.85 0.82 0.76 0.79 0.76 0.67 0.75 0.77
+Group-based AS 0.22 0.31 0.25 0.27 0.22 0.22 0.36 0.27 0.00
+Exchange-based AS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Empathic AS 0.50 0.67 0.47 0.00 0.25 0.36 0.00 0.00 0.00
+Compassionate AS 0.40 0.50 0.50 0.24 0.22 0.30 0.33 0.15 0.00
+Anti-solidarity (high-level) 0.59 0.65 0.53 0.30 0.38 0.39 0.44 0.48 0.12
+Mixed 0.50 0.30 0.38 0.33 0.37 0.43 0.40 0.32 0.15
+None 0.65 0.62 0.48 0.51 0.47 0.26 0.54 0.66 0.21
 Figure 21: Distribution of all Migrant keywords over the years, normalized per keyword. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
-Figure 20: Percentage of sentences showing solidarity/anti-solidarity per decade for all Migrant keywords from 1867 to 2025.
-The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
+Figure 22: (Anti-)solidarity trends for the Migrant category by keyword and decade, 1867-2025.
+Each panel shows DSL-adjusted decade-level shares of solidarity and anti-solidarity for one keyword (see definition (c) in Section˜3.2.2).
+The mixed and none categories are included in the four-class distribution but omitted from the visualization.
+Keywords are sorted by frequency, so estimates become less reliable toward the bottom right.
+(a) Human majority vote
+(b) Human soft labels
+(c) gpt-oss-120B
+(d) Model majority vote
+(e) Llama-3.3-70B
+(f) Qwen-2.5-72B
+(g) Model soft labels
+(h) DSL (soft labels)
+Figure 23: High-level solidarity trends for the Migrant category under selected labeling and correction strategies discussed in Section˜3.2.2 based on the 2k evaluation subset.
+Trends are computed as label shares within each decade (see definitions in Section˜3.2.2).
+The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
+(a) Human majority vote
+(b) Human soft labels
+(c) gpt-oss-120B
+(d) Model majority vote
+(e) Llama-3.3-70B
+(f) Qwen-2.5-72B
+(g) Model soft labels
+(h) DSL (soft labels)
+Figure 24: Fine-grained solidarity trends for the Migrant category under selected labeling and correction strategies discussed in Section˜3.2.2 based on the 2k evaluation subset.
+Trends are computed as label shares within each decade (see definitions in Section˜3.2.2).
+The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
+Label Model Pearson Spearman RMSE
+High-level trends (Human majority vote)
+Solidarity Llama-3.3-70B 0.907 0.894 0.101
+DSL (soft labels) 0.910 0.938 0.075
+Anti-solidarity Llama-3.3-70B 0.886 0.847 0.065
+DSL (soft labels) 0.928 0.921 0.041
+Mixed Llama-3.3-70B 0.810 0.830 0.029
+DSL (soft labels) 0.860 0.873 0.027
+None Llama-3.3-70B 0.910 0.906 0.134
+DSL (soft labels) 0.868 0.899 0.074
+Average Llama-3.3-70B 0.878 0.869 0.082
+DSL (soft labels) 0.892 0.908 0.054
+Fine-grained trends (Human majority vote)
+Group-based S Llama-3.3-70B 0.412 0.358 0.202
+DSL (soft labels) 0.679 0.678 0.062
+Exchange-based S Llama-3.3-70B 0.588 0.458 0.073
+DSL (soft labels) 0.461 0.508 0.055
+Empathic S Llama-3.3-70B 0.763 0.827 0.044
+DSL (soft labels) 0.559 0.519 0.022
+Compassionate S Llama-3.3-70B 0.823 0.772 0.204
+DSL (soft labels) 0.872 0.912 0.080
+Group-based AS Llama-3.3-70B 0.855 0.824 0.087
+DSL (soft labels) 0.930 0.908 0.058
+Exchange-based AS Llama-3.3-70B 0.706 0.642 0.024
+DSL (soft labels) 0.728 0.729 0.021
+Empathic AS Llama-3.3-70B 0.497 0.435 0.017
+DSL (soft labels) 0.365 0.442 0.018
+Compassionate AS Llama-3.3-70B 0.447 0.492 0.064
+DSL (soft labels) 0.842 0.660 0.036
+Average Llama-3.3-70B 0.636 0.601 0.089
+DSL (soft labels) 0.680 0.670 0.044
+Table 8: Per-label correlation between model predictions and human-annotated decade-level trends. Higher is better for Pearson and Spearman; lower is better for RMSE.
 Figure 25: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Migrant dataset.
 Figure 26: One-step zero-shot prompt set used for the evaluation of GPT-4 on the Frau dataset.
-Figure 23: (Anti-)solidarity trends in 1867-2022, restricted to the keywords Flüchtlinge and Ausländer.
-Figure 23(a) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes predicted by Llama-3.3-70B.
-Percentages are normalized within each decade across all eight subtypes, so the curves in both panels together sum to 100% per decade.
-Figure 23(b) shows the aggregated proportions of solidarity, anti-solidarity, and mixed labels per decade; totals may not reach 100% because instances classified as ?None?
-are omitted (though included in the calculation).
-The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
 (a) High-level prompt (Step 1).
 Figure 27: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B.
-The model is first given the high-level classification prompt in Figure 27(a).
-If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
+The model is first given the high-level classification prompt in Figure˜30(a).
+If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Figure˜31(a) or Figure˜32(a) is used.
 (a) Subtype prompt for Solidarity (Step 2a).
-Figure 25: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
+Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Figure˜27).
 (a) Subtype prompt for Anti-solidarity (Step 2b).
-Figure 26: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
+Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Figure˜27).
 (a) High-level prompt (Step 1).
 Figure 30: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset.
-The model is first given the high-level classification prompt in Figure 27(a).
-If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
+The model is first given the high-level classification prompt in Figure˜30(a).
+If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Figure˜31(a) or Figure˜32(a) is used.
 (a) Subtype prompt for Solidarity (Step 2a).
-Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
+Figure 31: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Figure˜30).
 (a) Subtype prompt for Anti-solidarity (Step 2b).
-Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
-Keyword Solidarity Anti-Solidarity Mixed None Total
-Asylsuchende 10 0 0 5 15
-Ausländer 168 52 9 525 754
-Aussiedler 99 1 5 61 166
-Einwanderer 23 1 0 21 45
-Emigranten 66 5 0 18 89
-Flüchtlinge 1828 23 84 1179 3114
-Heimatvertriebene 2471 20 68 1814 4373
-Sowjetzonenflüchtlinge 264 5 10 178 457
-Zuwanderer 32 0 2 46 80
-Zuwanderung 7 0 1 29 37
-Table 5: Distribution of Solidarity, Anti-Solidarity, Mixed, and None labels per keyword included in the post-war trends analysis (Section 4.2).
-BP CDU/CSU DP DRP DZP FDP GB/BHE KPD SPD WAV Z
-Group-based S. 95 707 56 13 23 221 118 50 667 93 5
-Exchange-based S. 5 163 5 4 1 46 14 8 73 13 0
-Compassionate S. 64 670 38 9 8 135 92 69 537 72 7
-Empathic S. 6 91 8 1 0 39 29 4 89 8 1
-Solidarity with no subtype 1 18 0 0 0 6 2 1 6 2 0
-Group-based AS 28 100 9 3 1 22 9 30 68 23 0
-Exchange-based AS 4 14 3 0 0 1 3 3 11 2 0
-Compassionate AS 1 3 1 0 0 0 0 3 3 0 0
-Empathic AS 0 1 0 0 0 0 0 0 0 0 0
-Mixed 15 63 5 1 2 23 8 4 33 3 0
-None 81 1564 78 10 11 389 94 54 754 47 4
-Total per party 300 3394 203 41 46 882 369 226 2241 263 17
-Table 6: Distribution of (anti-)solidarity subtypes by party included in the post-war trends analysis (Section 4.2). AS stands for Anti-Solidarity; S for Solidarity.
+Figure 32: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Figure˜30).
+Figure 33: Solidarity trends for Migrant speeches computed using DSL-corrected soft labels (see definition (c) in Section˜3.2.2) on the restricted corpus containing only the keyword groups Flüchtlinge and Ausländer.
+Figure˜33(b) shows the high-level (anti-)solidarity trends, while Figure˜33(a) shows the distribution of fine-grained (anti-)solidarity subtypes over time.
+In the fine-grained panel, percentages are normalized within each decade across the eight solidarity and anti-solidarity subtypes.
+The period 1930-1949 (grey area) is excluded due to limited data during and after the Nazi dictatorship.
+Keyword Solidarity Anti-Solidarity Mixed None N (statements)
+Asylsuchende 4.11 2.06 0.48 3.35 10
+Ausländer 291.43 118.86 21.43 326.28 758
+Aussiedler 118.43 12.48 15.22 25.87 172
+Einwanderer 2.84 3.21 0.20 19.75 26
+Emigranten 21.14 11.08 2.61 10.18 45
+Flüchtlinge 1917.08 321.69 174.29 859.94 3273
+Heimatvertriebene 3158.67 422.75 167.72 1006.85 4756
+Sowjetzonenflüchtlinge 308.99 27.13 2.63 138.25 477
+Zuwanderer 58.24 6.54 0.86 60.36 126
+Table 9: Distribution of DSL-adjusted expected probability mass for solidarity, anti-solidarity, mixed, and none per keyword included in the post-war trends analysis Section˜4.2.
+Values reflect summed DSL-adjusted expected counts.
+Totals correspond to the number of keyword-coded statements in the analyzed subset.
+Party Solidarity Anti-Solidarity Mixed None N (statements)
+BP 134.13 54.77 67.51 46.60 303
+CDU/CSU 2040.48 229.98 94.93 1050.62 3416
+DP 138.21 40.81 -0.54 26.52 205
+DRP 34.53 1.47 1.74 3.26 41
+FDP 539.79 38.02 56.78 251.41 886
+GB/BHE 318.75 27.61 9.03 17.62 373
+KPD 151.75 57.43 -1.14 18.96 227
+SPD 1595.06 153.56 37.13 478.25 2264
+WAV 191.81 19.72 37.12 15.35 264
+Zentrum 53.61 3.02 1.35 5.01 63
+Table 10: Distribution of DSL-adjusted expected probability mass for solidarity, anti-solidarity, mixed, and none per party included in the post-war trends analysis Section˜4.2.
+Values reflect summed DSL-corrected expected counts.
+Totals correspond to the number of party-coded statements in the analyzed subset.
 Original Text Translation
-(1) Predicted label: group-based solidarity towards expellees
+[…]
+?
+[…] das Interessante […] ist, daß nicht nur die östliche Landwirtschaft sondern auch die Industrie Ausländer braucht, daß die Zahlen der industrietätigen Ausländer im Lause der Jahre viel mehr gewachsen sind als die Zahlen der von der Landwirtschaft gebrauchten Arbeiter.
+[…] Das Maß des Bedürfnisses der Beschäftigung von Ausländern ist natürlich auch hier örtlich und nach Betrieben verschieden.
+Es gibt gewisse Betriebe, für die inländische Arbeiter kaum noch zu bekommen sind.
+[…] Es gibt aber auch industrielle Gebiete, die ohne diese Ausländer gar nicht weiter existieren können.
+?
+?
+[…] what is interesting […] is that not only eastern agriculture but also industry needs foreigners, and that the number of foreigners employed in industry has grown over the years much more than the number of workers needed in agriculture.
+[…] The extent of the need for the employment of foreigners naturally also varies here by locality and by sector.
+There are certain sectors in which domestic workers can scarcely be found anymore.
+[…] There are also industrial regions that simply could not continue to exist without these foreigners.
+?
+[…]
+?[…] Die Beschleunigung der Asylverfahren ist überfällig, und eine konsequente Rückführung […], und […] eine konsequente Abschiebung.
+Übrigens sind zügige Verfahren auch und gerade im Interesse der Menschen mit Bleibeperspektive, die wir dann umso besser integrieren können, weil wir ihnen eine Perspektive geben können.
+Zu guter Letzt müssen wir die Einwanderung nach Deutschland insgesamt neu ordnen, mit einem Einwanderungsgesetz, das diesen Namen verdient.
+Herr Kollege Castellucci, möglicherweise wird umgekehrt ein Schuh daraus: Wer Akzeptanz für legale Migration erhöhen will, der muss auch entschlossen illegale Migration bekämpfen.
+[…]?
+?Speeding up asylum procedures is overdue, as are consistent returns […] and […] consistent deportations.
+Swift procedures are also, and especially, in the interest of people who have a prospect of staying, because we can then integrate them all the better by giving them a perspective.
+Finally, we must reorganize immigration to Germany as a whole, with an immigration law worthy of the name.
+Mr Castellucci, perhaps the reverse is true: anyone who wants to increase acceptance of legal migration must also resolutely combat illegal migration.
+[…]?
+[…]
+?[…] Sie [die europäischen Staaten] sind aber nicht bereit, Flüchtlinge vor Menschenrechtsverletzungen zu schützen, sondern diffamieren sie als Schmarotzer und Bedrohung unserer Sicherheit, wie dies leider auch gerade durch den Kollegen der SPD wieder geschehen ist.
+Der Vertrag von Maastricht sieht die gemeinsame Einführung eines Visumzwangs vor, wenn auf Grund einer Notlage in einem Land eine größere Fluchtbewegung zu erwarten ist.
+So wird ganz bewußt verfolgten oder unter sonstigen Notlagen leidenden Menschen die Flucht erheblich erschwert, oft sogar unmöglich gemacht.
+Flüchtlinge werden damit geradezu in die Arme von Schlepperorganisationen getrieben, über die man sich hier so gerne beschwert.?
+?[…] But they [the European states] are not willing to protect refugees from human rights violations; instead, they defame them as parasites and a threat to our security, as unfortunately just happened again in the remarks of the SPD colleague.
+The Maastricht Treaty provides for the joint introduction of a visa requirement if, because of an emergency in a country, a larger refugee movement is to be expected.
+In this way, people who are persecuted or otherwise suffering acute hardship are deliberately making it much harder to flee, often even impossible.
+Refugees are thereby virtually driven into the arms of smuggling organizations, about which people here so readily complain.?
+Table 11: Error analysis examples.
+Bold text is the main sentence, the other sentences are for context.
+The reference label is the final human consensus label.
+Where relevant, we also report the distribution of individual human annotations.
+Original Text Translation
+Example (4): Predictions: group-based solidarity (gpt-oss, Qwen), group-based anti-solidarity (Llama). Human reference: group-based solidarity (10 Apr 2024)
+?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht.
+Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird.
+Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […].
+Sie ?übersehen?
+in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben.
+Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen.
+Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist.
+[…] ?
+?[…] Let us turn to the AfD’s draft and to what makes it simply atrocious.
+You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved.
+Not only does this place all foreigners under general suspicion […], but it also violates EU law […].
+You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency.
+But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is.
+[…]?
+[…]
+?
+[…] Diese Rückführung [der Evakuierten] muß beschleunigt werden, und wir müssen gesetzgeberische Maßnahmen ergreifen, damit die Evakuierten nicht immer wieder auf lange Zeit vertröstet werden, sondern in absehbarer Zeit zu ihren Rechten, zu einer menschenwürdigen Behandlung kommen.
+[…] Gerade das Vertrauen dieser Millionen von Menschen verpflichtet uns, sie auch in den nächsten Jahren nicht zu enttäuschen, […] diesen Eingliederungsprozeß auf allen Gebieten mit dem größten Nachdruck zu fördern […].
+In der politischen Diskussion hat zwar das Problem der sachlichen, wirtschaftlichen und sozialen Interessen der Vertriebenen und aller Kriegsgeschädigten kaum noch eine Rolle gespielt […].
+?
+?
+[…] This return [of evacuees] must be accelerated, and we must take legislative measures so that evacuees are not kept waiting indefinitely time and again, but instead obtain their rights and humane treatment within a foreseeable period of time.
+[…] It is precisely the trust of these millions of people that obliges us not to disappoint them in the years ahead either, […] and to promote this process of integration in every sphere with the greatest possible determination […].
+In political debate, the problem of the material, economic, and social interests of the expellees and all other victims of war has in fact hardly played any role any longer […].
+?
+[…]
+?Die Bundesrepublik hat […] einen Vorbehalt zur UN-Kinderrechtskonvention zurückgenommen, der ihr zusichern sollte, dass sie ausländische Kinder schlechter behandeln kann als inländische.
+[…] In den vorliegenden Anträgen […] ist dargelegt, was alles zu tun wäre: Heraufsetzung der Verfahrensmündigkeit von minderjährigen Flüchtlingen von 16 auf 18 Jahre, Ausbau und Verbesserung der Betreuung unbegleiteter minderjähriger Flüchtlinge durch die Jugendfürsorge, ihre Herausnahme aus dem Flughafenverfahren, keine Abschiebehaft für Minderjährige, keine Unterbringung von Kindern und Jugendlichen in Sammelunterkünften für Asylsuchende und Flüchtlinge.
+Unbegleitete Kinder und Jugendliche, deren Eltern nicht ausfindig gemacht werden können, sollen genauso behandelt werden wie elternlose deutsche Kinder.
+[…] ?
+?The Federal Republic has […] withdrawn a reservation to the UN Convention on the Rights of the Child that had ensured it could treat foreign children worse than German children.
+[…] The motions submitted […] outline what still needs to be done: raising the age of legal capacity in proceedings for minor refugees from 16 to 18; expanding and improving youth welfare support for unaccompanied minor refugees; removing them from the airport procedure; no detention pending deportation for minors; no accommodation of children and adolescents in collective facilities for asylum seekers and refugees.
+Unaccompanied children and adolescents whose parents cannot be located should be treated in the same way as German children without parents.
+[…] ?
+Table 12: Error analysis examples.
+Bold text is the main sentence, the other sentences are for context.
+The reference label is the final human consensus label.
+Where relevant, we also report the distribution of individual human annotations.
+Original Text Translation
+[…]
+?Das Hamburger Modell ist geradezu vorbildlich; denn hier werden nicht Menschen mit geringem Einkommen und Geflüchtete gegeneinandergestellt, sondern alle haben den gleichen Anspruch auf Mobilität, mit dem Unterschied, dass die Geflüchteten das von ihrem Taschengeld zwangsweise abgezogen bekommen.
+[…] ?Wir sind auf Migration angewiesen, aber die Flüchtlinge…?, könnte man sagen.
+[…] Ich bin so erzogen: Wenn jemand Hilfe braucht und du kannst helfen, dann hilf.
+Ich bin stolz und froh, dass unser Land helfen kann […].?
+?The Hamburg model is virtually exemplary, because here people with low incomes and refugees are not being pitted against one another; rather, everyone has the same entitlement to mobility, with the difference that refugees have it forcibly deducted from their allowance.
+[…] One might say, ?We depend on migration, but the refugees…?.
+[…] I was raised to believe this: if someone needs help and you can help, then help.
+I am proud and glad that our country is able to help […].?
+[…]
+?[…] Die große Aufgabe, die uns hier gesetzt war, war der Versuch, einen Kompromiß […] zwischen den Interessen der Ausländer in der Bundesrepublik Deutschland und den berechtigten Ansprüchen der deutschen Bevölkerung auf Aufrechterhaltung der öffentlichen Ordnung und Sicherheit zu finden.
+Das Gesetz ist ein Ausdruck unserer freiheitlichen demokratischen Grundordnung, indem es den Ausländern in der Bundesrepublik ein sehr hohes Maß von Entfaltungsmöglichkeiten gewährt.
+[…] Das Gesetz nimmt auch besonders Bedacht auf eine Verbesserung der Rechtslage der ausländischen Flüchtlinge […], die ihre Heimat unter dem Druck totalitärer Systeme verlassen mußten.
+[…]?
+?[…] The major task before us was to try to find a compromise […] between the interests of foreigners in the Federal Republic of Germany and the legitimate claims of the German population to the maintenance of public order and security.
+This law is an expression of our liberal democratic constitutional order in that it grants foreigners in the Federal Republic a very high degree of scope for self-development.
+[…] The law also gives special consideration to improving the legal position of foreign refugees […] who were forced to leave their homeland under pressure from totalitarian systems.
+[…]?
+[…]
+?[…] Mit diesem auf Angebotsorientierung im Migrationsrecht zielenden Gesetzentwurf wollen die Grünen im Kern nichts anderes, als die Einwanderung massiv vereinfachen und die Tore weit aufstoßen.
+Wir als Unionsfraktion stellen uns gegen die Beliebigkeit im Einwanderungsrecht und stehen daher auch klar für eine gezielte Auswahl bei der Migration ein.
+Der Staat und die Politik haben die Pflicht, bewusst darauf zu achten, dass die Migration für unser Land vor allem einen Mehrwert und keine Belastung darstellt, und genau darin liegt der Unterschied zum Asyl- und Flüchtlingsrecht.
+[…] ?
+?[…] With this bill, which is aimed at a supply-oriented approach to migration law, the Greens essentially want nothing other than to make immigration much easier and throw the gates wide open.
+As the CDU/CSU parliamentary group, we oppose arbitrariness in immigration law and therefore clearly support a targeted selection in migration.
+The state and policymakers have a duty to ensure consciously that migration represents above all an added value for our country and not a burden, and that is precisely the difference from asylum and refugee law.
+[…]?
+Table 13: Error analysis examples.
+Bold text is the main sentence, the other sentences are for context.
+The reference label is the final human consensus label.
+Where relevant, we also report the distribution of individual human annotations.
+Original Text Translation
+[…]
+?52 Prozent der britischen Muslime sagen nicht nur: ?Homosexualität ist moralisch falsch?, sondern: Homosexualität sollte verboten werden.
+[…] bei fortgesetzter Einwanderung [wird sich] der Anteil der Muslime in Deutschland in den kommenden Jahren auf 20 Prozent erhöhen.
+In vielen Städten und Gemeinden sind sie dann die Mehrheit.
+Der Islam kennt keine Minderheitenrechte.
+Das sind die dunklen Wolken, die über den Homosexuellen in Westeuropa aufziehen.?
+?52 percent of British Muslims say not only, ?Homosexuality is morally wrong?, but also: homosexuality should be banned.
+[…] if immigration continues, the share of Muslims in Germany will rise to 20 percent in the coming years.
+In many cities and municipalities, they will then be the majority.
+Islam recognizes no minority rights.
+These are the dark clouds gathering over homosexuals in Western Europe.?
+[…]
+?Ausländer bleibt unerachtet solchen Aufenthalts innerhalb der Mecklenburgischen Lande immer Ausländer, und einem solchen die Trauung innerhalb der diesseitigen Lande zu gestatten, ist allein das Großherzogliche Ministerium des Innern competent.
+[…] Vielmehr muß in solchen Fällen der Wiederverheirathung der Mecklenburger sein Recht zur Verheiratung durch einen gesetzlichen Domizilschein über das ihm für einen Mecklenburgischen Ort ertheilte Recht der Niederlassung und Verheiratung nachweisen und der Ausländer den Konsens des Großherzoglichen Ministeriums des Innern beibringen, bevor die Trauung geschehen darf.
+Ja, meine Herren, da sehen Sie nun, was in Mecklenburg das Indigenat aus dem Artikel 2 der Bundesverfassung für eine Bedeutung hat, und wie nothwendig es ist, daß irgend eine Ergänzung des Freizügigkeitsgesetzes […] erfolge, damit das Gesetz dort nicht blos eine Schale ohne Kern bleibe.?
+?A foreigner remains a foreigner within the lands of Mecklenburg despite such residence, and only the Grand Ducal Ministry of the Interior is competent to permit such a person to marry within these territories.
+[…] Rather, in such cases of remarriage, the Mecklenburger must prove his right to marry by means of a lawful certificate of domicile showing that he has been granted the right of residence and marriage in a Mecklenburg locality, and the foreigner must produce the consent of the Grand Ducal Ministry of the Interior before the marriage may take place.
+Gentlemen, there you can see what significance indigeneity under Article 2 of the federal constitution has in Mecklenburg, and how necessary it is that some supplementation of the Freedom of Movement Act […] should be made, so that the law there may not remain merely a shell without a kernel.?
+[…]
+?Wann und warum ein Ausländer ohne ein vorangegangenes gerichtliches Erkenntniß aus dem Reich oder einem Partikularstaat ausgewiesen werden darf, darüber existiren zwar keine gesetzlichen Normen und Bestimmungen […].
+Rechtfertigende Gründe der Ausweisung eines Ausländers nach dem allgemein anerkannten Völkerrechte sind folgende: erstens Gefährdung der öffentlichen Sicherheit durch ein feindliches Auftreten des Ausgewiesenen […]; zweitens […]; drittens[…], und viertens […].
+Damit sind alle Fälle erschöpft, in welchen nach dem modernen Völkerrecht Ausweisungen von einzelnen Personen stattfinden dürfen, und alle Lehrer des Völkerrechts sind um so vorsichtiger in der Empfehlung auch nur eines dieser Mittel, als sich das moderne Völkerrecht immer mehr der Theorie zuwendet, daß ein domizilirter Ausländer zu den wirklichen Staatsangehörigen gehört und als solcher zu behandeln ist.
+[…]?
+?Although there are no legal norms or provisions governing when and why a foreigner may be expelled from the Empire or from an individual state without prior judicial judgment […], the reasons justifying the expulsion of a foreigner under generally recognized international law are the following: first, endangerment of public security through hostile conduct on the part of the person expelled […]; second […]; third […]; and fourth […].
+These exhaust all the cases in which, under modern international law, expulsions of individual persons may take place, and all teachers of international law are all the more cautious in recommending even one of these measures because modern international law is increasingly turning toward the theory that a domiciled foreigner belongs among the true members of the state and is to be treated as such.
+[…]?
+Table 14: Error analysis examples.
+Bold text is the main sentence, the other sentences are for context.
+The reference label is the final human consensus label.
+Where relevant, we also report the distribution of individual human annotations.
+Original Text Translation
+(1) Dominant DSL subtype score: group-based solidarity towards expellees
 (May 14, 1952)
@@
 Im Gegenteil, […] daß die Rücksiedlung […] nur durchgeführt werden kann, wenn die körperliche, seelische und auch die wirtschaftliche Substanz der Vertriebenen erhalten geblieben ist, und das kann nur geschehen, wenn sie in die westdeutsche Wirtschaft eingegliedert werden.
-[…] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten; […] aber der Rechtsanspruch darf nicht so behandelt werden, als sei er nur etwas Zweitrangiges, angewiesen auf das, was übrigbleibt.?
+[…] Wir sind durchaus für die Ausgestaltung des Lastenausgleichs nach sozialen Gesichtspunkten […].
+?
 ?[…] the Equalization of Burdens Act and the integration of the expellees does not contradict our right to the homeland.
 On the contrary, resettlement […] can only be achieved if the physical, mental, and economic substance of the expellees has been preserved, and that can only happen if they are integrated into the West German economy.
-[…] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations; […] but the legal entitlement should not be treated as secondary, dependent only on what is left over.?
-(2) Predicted label: exchange-based solidarity towards expellees
+[…] We are fully in favor of shaping the Equalization of Burdens in accordance with social considerations […].
+secondary, dependent only on what is left over.
+?
+(2) Dominant DSL subtype score: group-based solidarity towards expellees
 (Feb 25, 1953)
-?Ich darf […] auf einige Gesichtspunkte hinweisen […] zum Bundesvertriebenengesetz.
-Der Ihnen vorliegende Entwurf geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen.
+?[…] Der Ihnen vorliegende Entwurf [zum Bundesvertriebenengesetz] geht […] von der Zielsetzung aus, die […] Leistungskraft der Vertriebenen und Flüchtlinge […] im Rahmen der Gesamtwirtschaft zur Geltung zu bringen.
 Die Eingliederungsbestimmungen sollen der Erfüllung berechtigter Ausgleichsansprüche hinsichtlich der verlorenen Existenz dienen […] mit dem Ziel, daß sich das gesamte Arbeitspotential, das der Einheimischen und der Vertriebenen, so betätigen kann, daß eine größtmögliche Wirkung erzielt wird.?
-?I would like to point out some aspects […] on the Federal Expellees Act.
-The draft before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy.
+?[…] The draft [on the Federal Expellees Act] before you starts from the objective […] of making […] productive capacities of the expellees and refugees […] effective within the overall economy.
 The integration provisions are intended to meet legitimate compensation claims with regard to lost livelihoods […] with the goal that the entire labor potential, that of natives and of expellees alike, can be deployed in such a way that the greatest possible effect is achieved.?
-(3) Predicted label: compassionate solidarity towards Soviet-zone refugees
+(3) Dominant DSL subtype score: compassionate solidarity towards Soviet-zone refugees
 (June 15, 1955)
@@
 But if we want to be sure that this reform really has a positive effect for the refugee, then […] the planned increase of the hardship fund […] must actually be implemented.?
-(4) Predicted label: group-based anti-solidarity towards expellees
+(4) Dominant DSL subtype score: group-based anti-solidarity towards expellees
 (Dec 13, 1950)
@@
 The burdens are unbearable.?
-Table 7: Examples for postwar period discussed in Section 4.2. Bold text is the main sentence, the other sentences are for context.
+Table 15: Examples for postwar period discussed in Section˜4.2.
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
 Original Text Translation
-(5) Predicted label: group-based solidarity towards Soviet-zone refugees
+(5) Dominant DSL subtype score: group-based/compass. solidarity towards Soviet-zone refugees
 (Sept 23, 1954)
-?[…] den Ländern [wurden] für die wohnungsmäßige Unterbringung der […] Sowjetzonenflüchtlinge […] 350 Millionen DM […] zur Verfügung gestellt […].
-Der Bundestag ist […] davon ausgegangen, daß die Sowjetzonenflüchtlinge im Lastenausgleich nicht berücksichtigt werden konnten und daß die Verweisung […] auf den Härtefonds nur als Interimslösung gedacht war.
-[…] Die Bundesregierung prüft […] noch […] Maßnahmen, um den bei den Sowjetzonenflüchtlingen auftretenden Notständen abzuhelfen.?
-?[…] the federal government provided 350 million DM for the housing of […] Soviet-zone refugees […].
-Parliament assumed that Soviet-zone refugees could not be included in the Equalization of Burdens and that referral […] to the hardship fund was intended only as an interim solution.
-[…] The federal government is still examining measures to address the emergencies arising among the Soviet-zone refugees.?
-(6) Predicted label: group-based anti-solidarity towards foreigners
+?[…] [Sowjetzonenflüchtlinge] meinen, es könnte ihnen sonst bei mangelnder Bereitwilligkeit der Bevölkerung zu neuen Opfern vielleicht nicht einmal mehr das zur Verfügung stehen, was ihnen heute aus dem Härtefonds zur Verfügung steht.
+[…] Wir halten es für unvertretbar, daß […] diese nach wie vor auf den genannten, im Rahmen des Gesamtlastenausgleichs verhältnismäßig geringen Betrag ohne Rechtsanspruch angewiesen sein sollen.
+[Die], deren Urstatus dem der Heimatvertriebenen in Westdeutschland doch gleich ist, im Lastenausgleich diesen Heimatvertriebenen nicht gleichgestellt sind.?
+?[…] [Soviet-zone refugees] fear that, if public willingness for further sacrifices declines, they may no longer receive even the support currently provided through the hardship fund […].
+[It is] indefensible that […] they should still be dependent on the aforementioned, comparatively small amount within the overall equalization scheme without any legal entitlement.
+[…] those whose original legal status is equivalent to that of expellees in West Germany are not treated equally under the Equalization of Burdens system.?
+(6) Dominant DSL subtype score: group-based anti-solidarity towards foreigners
 (June 21, 1956)
 ?[…] Wir müssen das Bundesgebiet nicht nur gegen politisch subversive Elemente, wir müssen uns auch gegen kriminelle Elemente abschirmen.
-[…] Für die Bundesrepublik kommt […] hinzu, daß sie […] Anlaß hat, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden.
+[…] die Bundesrepublik […] [hat] Anlaß, sich gegen Zehntausende von heimatlosen Ausländern abzuschirmen, die […] in der Absicht [kommen], bei uns wirtschaftliche Versorgung oder politisches Asyl zu finden.
 […] alle internationalen Verpflichtungen, die wir […] erfüllen, [gehören] selbstverständlich nicht unter das, was ich hier ausgeführt habe.?
 ?[…] We must shield the Federal territory not only against politically subversive elements, we but also against criminal elements.
-[…] For the Federal Republic there is the added circumstance that it has reason to shield itself against tens of thousands of stateless foreigners who […] come with the intention of seeking economic support or political asylum here.
+[…] the Federal Republic […] has reason to shield itself against […] stateless foreigners who […] come with the intention of seeking economic support or political asylum here.
 […] all international obligations which we […] fulfill, of course, do not fall under what I have set out here.?
-(7) Predicted label: group-based solidarity towards foreigners
+(7) Dominant DSL subtype score: group-based solidarity towards foreigners
+(Oct 18, 1950)
+?[…] [HAuslG] […] [wird] den Besatzungsmächten den Entschluß erleichtern, der Bundesrepublik die volle Verantwortung für die heimatlosen Ausländer zu übertragen.
+Wenn […] das deutsche Volk seine Bereitschaft bekundet, die Eingliederung der heimatlosen Ausländer in die deutsche Betreuung […] zu erleichtern, so gehen wir dabei von der Erwartung aus, daß auch die heimatlosen Ausländer bemüht sein werden, sich in die nun einmal gegebene Lage einzufügen.
+Sie werden sich dabei bewußt sein müssen, daß auch die Lage der deutschen Bevölkerung schwierig ist und daß viele Zehntausende insbesondere von Heimatvertriebenen noch immer arbeitslos sind und über keinen angemessenen Wohnraum verfügen.
+[…] der heimatlose Ausländer [wird] viele der Härten dieser Nachkriegszeit mit den Deutschen im Gebiete der Bundesrepublik teilen müssen.
+?
+?
+[…] [HAuslG] […] will make it easier for the occupying powers to decide to transfer full responsibility for stateless foreigners to the Federal Republic.
+If […] the German people declare their willingness to facilitate […] the integration of stateless foreigners into German administration and care, we proceed on the assumption that the stateless foreigners will also strive to adapt to the situation as it now stands.
+They must remain aware that the situation of the German population is also difficult and that many tens of thousands, especially expellees, remain unemployed and lack adequate housing.
+[…] stateless foreigners will have to share many of the hardships of this postwar period with the Germans in the territory of the Federal Republic.
+?
+(8) Dominant DSL subtype score: group-based solidarity towards foreigners
 (Oct 18, 1950)
 ?[…] Auch in der öffentlichen Fürsorge sollen sie Leistungen in gleicher Höhe erhalten wie deutsche Staatsangehörige.
-Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen wie Deutschen gewährt werden.
+Es soll ihnen Zutritt zu Schulen aller Art zu den gleichen Bedingungen… gewährt werden.
 […] die Bundesrepublik [gewährt] ihnen einen Rechtsschutz, dessen sie bedürfen, weil sie nicht mehr den Schutz ihres Heimatstaates genießen […].
@@
 ?[…] In public welfare too, they are to receive benefits equal to those of German citizens.
-They are to be granted access to schools of all kinds under the same conditions as Germans.
-[…] the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state […].
-The draft law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners.
+They are to be granted access to schools of all kinds under the same conditions […].
+the Federal Republic shall grant them legal protection, which they need because they no longer enjoy the protection of their home state.
+The law goes […] beyond the draft convention, whose principles set minimum standards for the treatment of stateless foreigners.
 […]?
-(8) Predicted label: group-based solidarity towards foreigners
+Table 16: Examples for postwar period discussed in Section˜4.2 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
+Original Text Translation
+(9) Dominant DSL subtype score: exchange-based solidarity towards foreigners
 (April 19, 1951)
@@
 […]?
-(9) Predicted label: empathic solidarity towards foreigners
+(10) Dominant DSL subtype score: group-based solidarity towards foreigners
 (June 15, 1955)
@@
 […]?
-Table 8: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
-Original Text Translation
-(10) Predicted label: exchange-based solidarity towards expellees,
+(11) Dominant DSL subtype score: exchange-based solidarity towards expellees,
 Party: FDP
@@
 If we only charged normal rents and let the capital bear no or low interest, then within at most fifteen years this house […] would become the property of the person concerned.?
-(11) Predicted label: group-based solidarity towards expellees,
+(12) Dominant DSL subtype score: group-based/
+compassionate solidarity towards expellees,
 Party: GB/BHE
@@
 once again.?
-(12) Predicted label: compassionate solidarity towards refugees,
+Table 17: Examples for postwar period discussed in Section˜4.2 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
+Original Text Translation
+(13) Dominant DSL subtype score: compassionate solidarity towards refugees,
 Party: WAV
@@
 […]?
-(13) Predicted label: group-based anti-solidarity towards expellees,
+(14) Dominant DSL subtype score: group-based anti-solidarity towards expellees,
 Party: BP
@@
 ?
-Table 9: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
-Original Text Translation
-(14) Predicted label: group-based solidarity towards refugees,
+(15) Dominant DSL subtype score: group-based solidarity towards refugees,
 Party: KPD
@@
 If the CDU speaker here has coined the fine, proud, and correct phrase ?Our house is German?, and if he concluded by saying ?It shall be the whole of Germany?, how then can one at the same time describe the fact that the expellees, refugees, and resettlers, as you call them, have politically organized themselves into parties as a danger of being overrun by foreigners??
-Table 10: Examples for postwar period discussed in Section 4.2 (continued). Bold text is the main sentence, the other sentences are for context.
+(16) Dominant DSL subtype score: exchange-based solidarity towards refugees
+Party: CDU/CSU
+(Jan 31, 1951)
+?[…] Es ist Voraussetzung, dass das deutsche Volk für das System des Lastenausgleichs einen Weg findet, damit das, was aus der deutschen Produktion und Wirtschaftskraft herausgenommen werden muss, […] dafür verwendet wird, die Kräfte der Kriegsgeschädigten, der Flüchtlinge und der Heimatvertriebenen wieder in die deutsche Produktion hineinzustellen und diese damit zu stärken.
+[…] Was wir an Eingliederungshilfe […], an Wohnraumhilfe […] und an Förderungshilfe aller Art […] verwenden, ist […] produktive Ausgabe.
+[…] das ist rund die Hälfte des gesamten Aufkommens.
+?
+?[…] It is a prerequisite that the German people find a way, within the system of burden sharing, to ensure that what must be taken from German production and economic strength is used […] to restore the productive capacity of war victims, refugees, and expellees and to reintegrate them into German production, thereby strengthening it.
+[…] What we allocate for integration assistance […], for housing assistance […], for all kinds of development assistance is […] productive expenditure.
+[…] that is roughly half of the total revenue.?
+Table 18: Examples for postwar period discussed in Section˜4.2 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
 Original Text Translation
-(15) Predicted label: compassionate solidarity towards refugees
+(17) Dominant DSL subtype score: group-based anti-solidarity towards expellees,
+Party: DP
+(March 18, 1953)
+?Also auch der dritte, vierte und fünfte Sohn eines ostdeutschen Bauern […] kann alle […] Vergünstigungen erhalten.
+[…] Welcher einheimische Bauer kann […] seinen drei oder vier Jungen eine solche Mitgift erarbeiten?
+Wir schaffen, wenn wir die alte Vorlage wiederherstellen, zweierlei Recht.
+Das dürfte so viel böses Blut zur Folge haben, dass das auch nicht im Interesse der Vertriebenen liegen kann.
+Herr Kollege Dr.
+Kather hat vorhin gesagt, dass 300 000 bäuerliche Familien aus dem Osten geflüchtet sind.
+Infolge der Formulierung des § 35 ist die Zahl der Bewerber um einen Betrieb weitaus größer geworden.
+Ich meine, wir sind unehrlich, wenn wir den Leuten Hoffnungen machen, die doch nicht zu erfüllen sind.
+?
+?Thus even the third, fourth, and fifth son of an East German farmer […] can receive all the benefits.
+[…] Which local farmer […] can earn such a dowry for three or four of his sons?
+If we restore the old version of the bill, we create two kinds of law.
+That would cause so much bad blood that it cannot possibly be in the interest of the expellees themselves.
+[…] Dr.
+Kather said that 300,000 farming families fled from the East.
+Because of the wording of § 35, the number of applicants for a farm has become far greater.
+I believe we are being dishonest if we give people hopes that cannot in fact be fulfilled.?
+Table 19: Examples for postwar period discussed in Section˜4.2 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
+Original Text Translation
+(18) Dominant DSL subtype score: compassionate solidarity towards refugees
 Party: Die Grünen
@@
 ?Ich bin froh, dass es heute bei uns in Deutschland eine andere Willkommenskultur gibt und dass sich viele Menschen mit hohem Engagement für das Wohl der Flüchtlinge, die im Augenblick zu uns kommen, einsetzen.
-[…] Ich will ganz klar sagen: Das Boot ist nicht voll; das Boot ist niemals voll.
+[…] Ich will ganz klar sagen: Das Boot ist nicht voll […].
 Heute geht es um Grundrechte von Menschen, denen die Grundrechte in ihren Heimatländern versagt wurden […].
@@
 ?I am glad that today in Germany we have a different welcoming culture, and that many people are working with great commitment for the well-being of the refugees who are currently coming to us.
-[…] I want to state very clearly: the boat is not full; the boat is never full.
+[…] I want to state very clearly: the boat is not full […].
 Today it is about the fundamental rights of people who have been denied those rights in their home countries […].?
-(16) Predicted label: group-based anti-solidarity towards refugees
+(19) Dominant DSL subtype score: group-based anti-solidarity towards refugees
 (Oct 1, 2015)
-?[…] Wenn die Zahlen weiterhin auf diesem hohen Niveau bleiben, dann wird Deutschland über kurz oder lang überfordert sein.
+?[…] Deutschland [wird] über kurz oder lang überfordert sein.
 Was die Registrierung, Antragstellung und Antragsbearbeitung anbelangt, ist dies eine riesige Herausforderung, und die Unterbringung sowie die Integration in unsere Gesellschaft und in den Arbeitsmarkt sind […] auf die Dauer nicht zu leisten […].
-allein im September sind mehrere zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen.
+allein im September sind zehntausend Flüchtlinge und Asylbewerber nach Deutschland gekommen.
 Sie wurden nicht registriert.
 Sie wurden nicht kontrolliert.
-[…] Damit besteht auch ein großes Sicherheitsrisiko.
+[…] Damit besteht auch ein Sicherheitsrisiko.
 ?
-?[…] If the numbers remain at this high level, Germany will sooner or later be overburdened.
+?[…] Germany will sooner or later be overburdened.
 Registration, application processing and decision-making are a huge challenge, and accommodation as well as integration into our society and the labor market will […] not be sustainable in the long run […].
@@
 […] This poses a major security risk.?
-(17) Predicted label: exchange-based anti-solidarity towards refugees
+Table 20: Examples for the 2009-2025 period discussed in Section˜4.3.
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
+Original Text Translation
+(20) Dominant DSL subtype score: exchange-based anti-solidarity towards refugees
 Party: AfD
 (May 16, 2018)
-?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug nach derzeitigem Stand um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind.
-In Deutschland, meine Damen und Herren, lebt bereits eine große soziale Unterschicht […].
+?[…] jährlich [verlassen] etwa 140 000 Hochkompetente […] dieses Land und im Gegenzug […] um die 200 000 sogenannte Flüchtlinge hereinkommen, von denen nicht nur kaum einer hochkompetent ist, sondern viele Analphabeten sind.
+In Deutschland lebt bereits eine große soziale Unterschicht […].
 Diese Zahl wächst seit 2015 durch die Zuwanderung von Gering- und Unqualifizierten rasant.
-Meine Damen und Herren, Deutschland ist kein reiches Land, sondern ein leistungsfähiges.
+[…] Deutschland ist kein reiches Land, sondern ein leistungsfähiges.
 ?
-?[…] each year about 140,000 highly qualified people leave this country, while in return, according to the current situation, about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate.
-In Germany, ladies and gentlemen, there already lives a large social underclass […].
+?[…] each year about 140,000 highly qualified people leave this country, while in return, […] about 200,000 so-called refugees enter, of whom hardly any are highly qualified, but many are illiterate.
+In Germany, there already lives a large social underclass […].
 Since 2015, this number has been growing rapidly through the immigration of low- and unqualified people.
-Ladies and gentlemen, Germany is not a rich country, but a capable one.?
-Table 11: Examples for the 2009-2025 period discussed in Section 4.3. Bold text is the main sentence, the other sentences are for context.
-Original Text Translation
-(18) Predicted label: group-based anti-solidarity towards refugees
+[…] Germany is not a rich country, but a capable one.?
+(21) Dominant DSL subtype score: group-based anti-solidarity towards refugees
 Party: AfD
@@
 Baerbock, feels primarily obligated to the Ukrainian population should run for office in Ukraine, but not in Germany.?
-(19) Predicted label: group-based anti-solidarity towards refugees
+(22) Dominant DSL subtype score: group-based anti-solidarity towards refugees
 Party: AfD
-(Dec 5, 2023)
+(12 May, 2023)
 ?[…] alte Leute sollen lieber aus ihren großen Wohnungen ausziehen, weil grün-linke Politik Millionen Gäste aus Afrika und dem Nahen Osten einlädt und gleichzeitig den Wohnungsbau abwürgt.
@@
 Es ist die Fortsetzung einer inländerfeindlichen… ?
-?[…] elderly people are supposed to move out of their large apartments because green-left politics invites millions of guests from Africa and the Middle East while simultaneously stifling housing construction.
-[…] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration [^12] Messereinwanderung: a propagandistic term coined by the AfD implying migrants bring knife crime.
-included.
+?[…] elderly people are expected to leave their large apartments because green-left politics invites millions of guests from Africa and the Middle East while choking off housing construction.
+[…] Rooms in supposedly Christian nursing homes are already being canceled because so-called refugees are more lucrative, knife migration included.
 […] This ?right to apartment swaps?
-proposed by the Left is a mixture of envy at the expense of elderly people and coercion at the expense of landlords.
-It is the continuation of an anti-native policy…?
-(20) Predicted label: exchange-based solidarity towards immigrants
+proposed by the Left is a mixture of envy at the expense of the elderly and coercion at the expense of landlords.
+It is the continuation of an anti-native [policy]…?
+(23) Dominant DSL subtype score: exchange-based solidarity towards immigrants
 Party: FDP
@@
 […]?
-(21) Predicted label: group-based anti-solidarity towards migrants
+Table 21: Examples for the 2009-2025 period discussed in Section˜4.3 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
+Original Text Translation
+(24) Dominant DSL subtype score: group-based anti-solidarity towards migrants
 Party: CDU/CSU
@@
 That is why it is good that we have agreed together in the coalition agreement on a migration turnaround: more deportations, faster procedures, and, yes, also rejections at the German borders, even in the case of asylum applications.?
-Table 12: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
-Original Text Translation
-(22) Predicted label: group-based anti-solidarity towards immigrants
+(25) Dominant DSL subtype score: exchange-based anti-solidarity towards immigrants
 Party: FDP
@@
 […] Ever more left-leaning policies […] will only enable the AfD to continue its diabolical work […].?
-(23) Predicted label: group-based anti-solidarity towards foreigners
-Misclassified; true: group-based solidarity
+(26) Dominant DSL subtype score: group-based anti-solidarity towards migrants
 Party: SPD
-(Apr 10, 2024)
-?Kommen wir zum Entwurf der AfD und zu dem, was ihn einfach grottenschlecht macht.
-Sie fordern, dass in allen Fällen, in denen ein Elternteil nichtdeutscher Herkunft ist, die Ausländerbehörde beteiligt wird.
-Nicht nur stellen Sie damit alle Ausländer/-innen unter Generalverdacht […], sondern Sie verstoßen damit auch gegen EU-Recht […].
-Sie ?übersehen?
-in Ihrem Entwurf, dass ein Großteil der Ausländer, die hier die Vaterschaftsanerkennung beantragen, längst eine Niederlassungserlaubnis haben.
-Aber gut, dass Sie gegen unsere ausländischen Mitbürgerinnen und Mitbürger hetzen.
-Ist ja wirklich Ihr politisches Tagesgeschäft, so bedauerlich das auch ist.
-[…] ?
-?Let us turn to the AfD’s draft and to what makes it simply atrocious.
-You demand that in all cases where one parent is of non-German origin, the immigration authority must be involved.
-Not only does this place all foreigners under general suspicion […], but it also violates EU law […].
-You overlook in your draft that a large proportion of the foreigners who apply here for recognition of paternity already have permanent residency.
-But of course, it fits your agenda to agitate against our foreign fellow citizens — that is really your daily political business, as regrettable as it is.
-[…]?
-(24) Predicted label: group-based anti-solidarity towards migrants
-Party: SPD
-(Dec 9, 2024)
+(12 Sept, 2024)
 ?Das Sicherheitspaket, das heute hier zur Debatte steht, ist der erste Schritt, um Konsequenzen aus der furchtbaren Katastrophe in Solingen zu ziehen.
@@
 We will take further steps to ensure that persons obliged to leave the country actually leave our country.?
-(25) Predicted: group-based anti-solidarity towards migrants
-Misclassified; true: empathic solidarity
-Party: Die Linke
-(June 5, 2025)
-?[…] Migration wird pauschal als sicherheitspolitisches Problem dargestellt.
-Die AfD führt die sogenannte gescheiterte Migrationspolitik in einer Liste von Krisen auf – neben dem Krieg in der Ukraine, der Coronapandemie und der hybriden Bedrohung.
-Damit betreibt sie eine gefährliche und spaltende Gleichsetzung.
-Sie erklärt das bloße Vorhandensein von Migration zur Bedrohung der inneren und äußeren Sicherheit und nennt sie in einem Atemzug mit Terrorismus, Krieg und Katastrophen.
-?
-?Migration is presented across the board as a security policy problem.
-The AfD lists the so-called failed migration policy alongside the war in Ukraine, the coronavirus pandemic, and hybrid threats.
-In doing so, it engages in a dangerous and divisive equation.
-It declares the mere existence of migration to be a threat to internal and external security, mentioning it in the same breath as terrorism, war, and disasters.?
-Table 13: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
+Table 22: Examples for the 2009-2025 period discussed in Section˜4.3 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
 Original Text Translation
-(26) Predicted label: exchange-based solidarity towards immigrants
+(27) Dominant DSL subtype score: exchange-based solidarity towards immigrants
 Party: Die Grünen
@@
 And because we need immigration, we need a culture of welcome.?
-(27) Predicted label: empathic solidarity towards migrants
+(28) Dominant DSL subtype score: compassionate solidarity towards migrants
 Party: Die Linke
@@
 Less migration makes the country safer — that is nonsense.?
-Table 14: Examples for the 2009-2025 period discussed in Section 4.3 (continued). Bold text is the main sentence, the other sentences are for context.
+(29) Dominant DSL subtype score: compassionate anti-solidarity towards asylum seekers
+Party: CDU/CSU
+(19 Feb 2016)
+?Das sind aber nicht nur Flüchtlinge, die Anspruch auf Schutz haben.
+[…] bei mehr als der Hälfte [konnte] keine Schutzbedürftigkeit festgestellt werden.
+Wenn wir bei Asylbewerbern, bei denen schon bei der Ankunft absehbar ist, dass sie keine Bleibeaussicht haben, nicht in kürzester Zeit entscheiden und sie in ihre Herkunftsländer zurückführen, werden wir noch mehr Anreize schaffen, ohne berechtigten Grund nach Deutschland zu kommen.
+[…] Das ist nicht nur ein klares Signal im Sinne von ?Wir nehmen nur wirklich Schutzbedürftige auf?, sondern entlastet auch deutlich unsere Kommunen bei der Unterbringung.
+Man kann dabei auch nicht von einer Einschränkung oder Aushöhlung des Asylrechts sprechen.
+?
+?These are not only refugees who are entitled to protection.
+[…] in more than half of the cases no need for protection could be established.
+If, in the case of asylum seekers for whom it is already apparent upon arrival that they have no prospect of remaining, we do not make a decision as quickly as possible and return them to their countries of origin, we will create even more incentives to come to Germany without justified reason.
+[…] This is not only a clear signal in the sense of ?We only take in those who genuinely need protection,?
+but it also significantly relieves our municipalities with regard to accommodation.
+Nor can one speak of a restriction or hollowing-out of the right to asylum in this context.?
+Table 23: Examples for the 2009-2025 period discussed in Section˜4.3 (continued).
+Bold text is the main sentence, the other sentences are for context.
+The label shown is the instance’s dominant DSL subtype, defined as the subtype with the highest DSL-adjusted score.
+Appendix F Annotation Process Details
+Annotators first assigned a high-level stance label and, where applicable, an (anti-)solidarity subtype.
+To document their decisions, they also identified the specific resource at issue (e.g.
+time, money, rights, or access to education), selected an indicator associated with the chosen category, highlighted supporting text spans, and provided a brief free-text explanation (1-2 sentences).
+Where relevant, annotators also marked expressions of the speaker’s own position and references to others’ positions.
+These additional annotations supported interpretation and documentation during the annotation process, but were not used in the analyses reported in this paper.
+The full annotation guidelines are provided in Appendix˜G in the Appendix, and Figure˜34 shows an example from the annotation file.
+(a) Columns for high-level and (anti-)solidarity categorizations.
+(b) Columns for providing explanations.
+Figure 34: Example of the annotation process from the annotation file.
+Figure˜34(a) illustrates the step where annotators choose a high-level label and an (anti-)solidarity subcategory, if applicable.
+Figure˜34(b) shows columns for detailed explanations, including the choice of a resource, an indicator, and providing free-text commentary
+Appendix G Annotation Guidelines Excerpt
+This appendix reproduces the annotation guidelines used by the annotators, with minor formatting edits.
+General Information
+Goal of the study and annotation
+The purpose of this study is to examine expressions of solidarity and anti-solidarity towards two specific vulnerable groups—women and migrants—within German parliamentary proceedings.
+The definitions of (anti-)solidarity are derived from the work Ils et al.
+(2021), and Eger et al.
+(2022).
+We distinguish four subtypes of solidarity and anti-solidarity based on Thijssen and Verheyen (2022): group-based (shared identity or exclusion), compassionate (protection or dismissal of vulnerable groups), exchange-based (contribution, reciprocity, or burden), and empathic (respect for or rejection of diversity).
+Details of the Annotation Process
+Annotation instructions
+In the next section (Annotation Categories), we provide general theoretical information about the categories used in the annotation.
+This material is intended to support a better understanding of the concepts underlying the task.
+It is critical that annotators provide a clear explanation for the chosen category or categories in the Explanation column, while keeping comments short.
+The Other comments column can be used for any additional observations.
+Contextual evaluation
+Pay close attention to the middle text, which is the sentence whose stance is to be assessed. The surrounding sentences (previous and next) are provided to support contextual interpretation.
+Highlighting sentences
+When labeling the instances, annotators highlight parts of or full sentences that guided the label choice.
+Green is used for solidarity categories and red for anti-solidarity categories.
+In cases of mixed stance, solidarity-expressing parts are highlighted in green and anti-solidarity-expressing parts in red.
+Pay attention to the target group
+If the target group in the Category column is women, but the instance is about migrants, the instance should be marked as None.
+For example, if the target group is women, then a sentence such as ?all refugees must now leave this country?
+should also be marked as None.
+Structure of the Annotation File
+The annotation file contains the following fields: ID: ID of the sentence.
+The first four numbers after the word debates indicate the year.
+Category: Indicates the target group (woman or migrant).
+Pay close attention to the target group.
+If the target group in the Category column is women, but the instance is about migrants, the instance should be marked as None.
+Sentences: The text to annotate, consisting of Previous, Middle, and Next.
+The surrounding sentences are provided for contextual interpretation.
+However, the central focus of the decision should be the middle text.
+Main category: One of Solidarity, Anti-solidarity, Mixed stance, or None.
+(Anti-)solidarity subcategory: One of the subcategories, selected only for instances labeled as Solidarity or Anti-solidarity.
+Explanation: Resource: The resource on the basis of which solidarity or anti-solidarity is expressed (e.g., access to Germany, rights and well-being).
+Indicators: An indicator motivating the selected anti-solidarity or solidarity type.
+Indicators are given in drop-down boxes for the respective (anti-)solidarity type and are also listed in the Annotation Categories section.
+Free text: A short explanation for the chosen label (1–2 sentences).
+Annotators should provide a clear explanation for the selected category.
+Other comments: Any additional observations.
+Annotation Categories
+Definition of solidarity (based on Ils et al.
+(2021)): solidarity is the readiness to share one’s own resources with others (be it time, financial means, rights, access to education etc.).
+1.
+Solidarity Group-based solidarity Definition.
+Group-based solidarity stems from a perceived similarity among members of the same group, leading to a shared identity.
+The group-based solidarity frames often refer to a certain (desired) commonality and a sense of togetherness due to common interests and goals, shared values and norms, or common rights and duties.
+Characteristics.
+To identify group-based solidarity, look for: 1.
+References to a shared identity, common interests, goals, values, or norms within a group (e.g.
+It’s our collective responsibility to…; we stand together in the fight…; We must unite to…).
+2.
+The existence of mutual obligations or shared rights and duties.
+3.
+Indication that the group must be integrated or even assimilated.
+4.
+Possibly, the identification of an outgroup that contrasts with the in-group.
+Emphasis on unity and shared goals/duties/rights among the whole group.
+Compassionate solidarity Definition.
+Compassionate solidarity is primarily about recognising and supporting individuals or subgroups who are marginalised, disadvantaged, or vulnerable.
+It extends assistance to these individuals or subgroups, emphasising their need for protection or help.
+Characteristics.
+To identify compassionate solidarity, look for: 1.
+Acknowledgement of the challenges, hardships, or vulnerabilities faced by individuals or subgroups (e.g.
+We must pay special attention to the needs…; We need to recognize the difficulties…).
+2.
+Emphasis on the need for their protection, assistance, or special consideration due to their marginalised or disadvantaged status (e.g.
+It is our duty to support…).
+Emphasis on extra care, support, and protection for those most vulnerable (both in-group and out-group).
+Exchange-based solidarity Definition.
+Exchange-based solidarity is linked to the principle of the role of exchange between partners with complementary qualities.
+Exchange-based solidarity is coded when a speaker refers to the usefulness of ‘exchange partners’ in terms of their actual or future contributions or willingness to contribute.
+These exchange partners can be rewarded or stimulated, and may also be required to contribute more to receive support.
+Characteristics.
+To identify exchange-based solidarity, look for references to a group: 1.
+That was actively doing, is doing and / or will be doing something useful.
+2.
+Should either be rewarded for past, present or future behavior or that it should be supported to continue that useful behavior.
+3.
+Does not receive enough and needs more support for its contribution (do not confuse it with compassionate solidarity).
+Empathic solidarity Definition.
+Empathic solidarity is a concept grounded in the intersubjective valuation of differences among individuals, recognising and respecting their unique characteristics.
+In contrast to compassionate solidarity, empathic solidarity is coded when a difference is not something that is bad or something that should disappear but is rather a strength or something positive.
+Empathic solidarity is coded when a speaker acknowledges diversity, difference, or unique characteristics as aspects that should be respected and taken into account.
+Characteristics.
+To identify empathic solidarity, look for: 1.
+Acknowledgement of a group’s diversity and that this diversity must be respected; 2.
+Recognition of the group’s unique and distinct characteristics; 3.
+Affirmation of the group’s right to self-expression, which can relate to their specific needs, traits, views, or belief choices.
+4.
+Advocacy for the group’s freedom to develop themselves the way they want.
+2.
+Anti-solidarity Group-based anti-solidarity Definition.
+The group-based anti-solidarity often points to an out-group characterised by pronounced differences (concerning interests and goals, values and norms, individual rights and responsibilities), and an expression of exclusion and lack of support for that group based on these differences.
+Characteristics.
+To identify group-based anti-solidarity, look for: 1.
+References to differing interests, goals, responsibilities, values, or norms between the groups and at the same time not willing to provide support based on the differences, sense of belonging to different groups.
+2.
+A stronger emphasis on one’s own group’s interests (e.g.
+It’s every person for themselves…; We must stand up for our own interests…; We must not sacrifice our personal goals for…).
+3.
+Resistance to assimilation or integration with other groups.
+Compassionate anti-solidarity Definition.
+Compassionate anti-solidarity is about disregarding and excluding individuals or subgroups either on the basis that they are perceived to be in a “good position” already, or are viewed as undeserving of help.
+It neglects these individuals or subgroups, downplaying their need for protection or help.
+Characteristics.
+To identify compassionate anti-solidarity, look for: 1.
+Denial of the challenges, hardships, or vulnerabilities faced by individuals or subgroups (e.g.
+We shouldn’t focus on the problems of…; We don’t need to pay attention to the struggles of…).
+2.
+Minimising the need for their protection, assistance, or special consideration (e.g.
+It is not our responsibility to aid…) as they are already in a ?good position?
+or are in a ?bad position?, but still do not deserve help.
+Exchange-based anti-solidarity Definition.
+Exchange-based anti-solidarity focuses on the idea that some groups take more than they give.
+It brings attention to groups that seem to receive a lot of resources or support, but do not contribute much or cause problems.
+This perspective supports punishing such groups, giving them less help, or making them contribute more.
+Characteristics.
+To identify exchange-based anti-solidarity, look for references to a group: 1.
+Talk about a group that has received a lot, but has not given much back or has caused harm.
+2.
+Suggestions that this group should be punished, or get less than they are currently getting.
+3.
+Calls for the group to give or do more in return for help.
+Empathic anti-solidarity Definition.
+Empathic anti-solidarity refers to the cases when a speaker refers to a group indicating that this group is different from everyone else, but this difference (needs, characteristics, beliefs, opinions) should not be respected or recognised.
+The speaker could argue that this difference should not be acknowledged.
+The core logic behind this is: ?You’re not like me, and that’s a shame.
+I don’t understand or respect you or your differences?.
+Characteristics.
+To identify empathic anti-solidarity, look for: 1.
+Disregard for a group’s diversity, arguing that this diversity should not be respected.
+2.
+Ignorance or dismissal of the group’s unique and distinct characteristics.
+3.
+Denial of the group’s right to self-expression, which can relate to their specific needs, traits, views, or belief choices.
+3.
+Mixed stance Definition.
+Mixed stance refers to instances where the speaker expresses both support and opposition towards the target group within the same discourse or text.
+It involves contradictions or conflicting statements regarding the speaker’s stance towards the target group.
+Characteristics.
+1.
+Annotate instances where the speaker presents a mixed stance by expressing both support and opposition towards the target group in the same discourse or text.
+2.
+Look for contradictions or inconsistencies in the level of support or opposition expressed towards the target group.
+3.
+Note that the middle sentence has more weight.
+4.
+None Definition.
+None refers to instances where the speaker neither expresses explicit support nor opposition towards the target group.
+The statement is unrelated to either solidarity or anti-solidarity or is unrelated to the group altogether.
+Characteristics.
+1.
+Annotate instances where the speaker maintains a neutral position and does not express clear support or opposition towards the target group.
+2.
+Look for statements that provide factual information, or refrain from taking a stance or expressing personal opinion.
+3.
+Also label the instance as None if you think that it is unrelated to the target group.